### v1

# Recovery of rational functions via Hankel pencil method and sensitivities of the poles

Abstract
In the paper, we develop a new method for the recovery of rational functions. Our idea is based on the property that Fourier coefficients of rational functions have the exponential structure and reconstruction of this exponential structure with the ESPRIT method in the frequency domain. Further we present sensitivity analysis for poles of rational functions reconstructed with our method in case of unstructured and structured perturbations. Finally, we consider several numerical experiments and, using sensitivities, explain the recovery errors for poles.
Keywords: Rational functions, sparse exponential sums, matrix pencil, ESPRIT, Hankel matrices, sensitivity analysis, structured and unstructured perturbations.
AMS classification: 41A20, 42A16, 42C15, 65D15, 94A12.
1 Introduction
For $M\in\mathbb{N}$, we define a rational function $r(z)$ of type $(M-1,M)$ by means of its partial fraction decomposition
$r(z):=\sum\limits_{j=1}^{M}\frac{\gamma_{j}}{z-z_{j}},$ (1.1)
where $\gamma_{j},\,z_{j}\in\mathbb{C}$, $\gamma_{j}\neq 0$ and poles $z_{j}$ are pairwise distinct.
We study reconstruction of rational functions (1.1) using as input data a finite set of their Fourier coefficients. We use the property that Fourier coefficients of rational functions have the exponential structure and we reconstruct this exponential structure in the frequency domain applying the ESPRIT method [17, 15]. The idea of recovery of rational functions by estimation of signal parameters in the frequency domain is not new. Similar technique for the recovery of or approximation by rational functions was employed in [6, 18, 19, 20]. We present modification of these ideas, by applying the ESPRIT method for the exponential recovery. Since ESPRIT is based on using the Hankel pencil, it makes the method suitable for sensitivity analysis.
The other methods for the recovery of rational functions (in time domain) such as the AAA method [13], the RKFIT (Rational Krylov Fitting) method [3], the Vector Fitting algorithm [12] and the Loewner framework method [1], were considered in [11]. Note that, in [6, 18, 19, 20], similar to [11], reconstruction from given noisy data is limited to numerical observations.
In [21], for the Loewner framework method the authors study sensitivities of Loewner pencil eigenvalues (which are poles of a transfer function defined by the corresponding Loewner matrices) with respect to unstructured and structured perturbations. Following approach in [21], we present the sensitivity analysis for the poles computed by our method. According to our method, poles located inside and outside the unit circle are reconstructed separately as eigenvalues of some Hankel matrix pencils. Further we consider two kinds of Hankel pencil perturbations. For the sensitivity $\rho$ with respect to unstructured perturbations, we present the connection with condition numbers of Vandermonde matrices that are used in factorizations of the corresponding Hankel matrices. The sensitivity $\eta$ with respect to structured perturbations of Hankel pencil is studied in [22]. We present modification of these results for our special reconstruction problem. Note that the sensitivity of eigenvalues of Hankel pencil perturbations was studied also in [10], [2]. In [10] and [2], the authors measure both structured and unstructured perturbations by the same quantity. Following approach in [21], we consider two different quantities for structured and unstructured sensitivities respectively. Further using sensitivity analysis, we explain reconstruction errors for poles from noisy Fourier coefficients. We refer also to the paper [4] for further application of the sensitivity analysis of Hankel pencil.
The method offered in this paper is also very much related to our ESPIRA method for the exponential recovery [7][9]. In ESPIRA, we use the property that the classical Fourier coefficients of exponential sums have a rational structure and to reconstruct this rational structure we apply the AAA method for rational approximation [13] in the frequency domain. Thus the idea of the method for rational recovery used in this paper can be considered as converse to the idea of our ESPIRA method.
Outline. In Section 2, we present the main ideas of the ESPRIT method for exponential recovery. In Section 3 we consider sensitivity analysis for Hankel pencil eigenvalues. Subsection 3.1 is dedicated to unstructured perturbations and in Subsection 3.2, we consider structured perturbations. In Section 4, we describe our main results. In Subsection 4.1, we present main ideas of our method for the recovery of rational functions (1.1). Subsection 4.2 is dedicated to the sensitivity analysis of poles reconstructed with our method. Finally, in Section 5, we present numerical experiments.
Notation. Let us agree that throughout the paper, we will use the matrix notation $\mathbf{X}_{L,K}:=(X_{ij})_{i,j=1}^{L,K}$ for complex matrices of size $L\times K$ and the submatrix notation $\mathbf{X}_{L,K}(m:n,k:\ell)$, to denote a submatrix of $\mathbf{X}_{L,K}$ with rows indexed $m$ to $n$ and columns indexed $k$ to $\ell$, where first row and first column has index 1 (even though the row and column indices for the definition of the matrix may start with 0). For square matrices, we often use the short notation $\mathbf{X}_{K}$ instead of $\mathbf{X}_{K,K}$. By $\left(\mathbf{X}_{L,K}\right)^{\dagger}$ we denote the Moore-Penrose inverse of $\mathbf{X}_{L,K}$. We consider $\|\mathbf{X}_{L,K}\|_{2}:=\sigma_{\max}(\mathbf{X}_{L,K})$ the 2-norm or spectral norm, and $\|\mathbf{X}_{L,K}\|_{\mathrm{F}}:=\left(\sum\limits_{i=1}^{L}\sum\limits_{j=1}^{K}|X_{ij}|^{2}\right)^{1/2}=\left(\sum\limits_{i=1}^{\min\{L,K\}}\sigma_{i}^{2}\right)^{1/2}$ the Frobenius norm of the matrix $\mathbf{X}_{L,K}$. The $\ell_{1}$- and $\ell_{2}$-norm of the vector $\boldsymbol{x}=(x_{1},\ldots,x_{K})$ are given as usually by $\|\boldsymbol{x}\|_{1}=\sum\limits_{j=1}^{K}|x_{j}|$ and $\|\boldsymbol{x}\|_{2}=\left(\sum\limits_{j=1}^{K}|x_{j}|^{2}\right)^{1/2}$ respectively.
2 ESPRIT method for the recovery of exponential sums
As was described in Introduction, we will apply the ESPRIT method [17, 15] to reconstruct the exponential structure of the Fourier coefficients of rational functions (1.1). In this section, we explain the main ideas of this method.
Let $M\in\mathbb{N}$ and $f$ be an exponential sum defined by
$f(t):=\sum\limits_{j=1}^{M}\gamma_{j}z_{j}^{t},\ \gamma_{j},z_{j}\in\mathbb{C},\ \gamma_{j}\neq 0,$ (2.1)
where nodes $z_{j}$ are pairwise distinct. Let $L,N\in\mathbb{N}$ be given such that $M\leq L\leq N$, $L$ is the upper bound for $M$. We want to reconstruct an exponential sum (2.1), i.e. to reconstruct parameters $M$, $\gamma_{j}$ and $z_{j}$ for $j=1,\ldots,M$, using as input data samples $\boldsymbol{f}:=\big{(}f(0),f(1),\ldots,f(2N-1)\big{)}^{T}$.
For $n,m\in\mathbb{N}$ and a vector $\boldsymbol{a}:=(a_{1},\ldots,a_{m})^{T}$, we define a Vandermonde matrix $\mathbf{V}_{n,m}:=\mathbf{V}_{n,m}(\boldsymbol{a})=(a_{j}^{k})_{k=0,j=1}^{n-1,m}$. Let now $\boldsymbol{z}:=(z_{1},\ldots,z_{M})^{T}$ be a vector of nodes of the exponential sum (2.1) and let $\mathbf{\Gamma}_{M}:=\operatorname{diag}(\gamma_{j})_{j=1}^{M}$ and $\mathbf{Z}_{M}:=\operatorname{diag}(z_{j})_{j=1}^{M}$. We create a Hankel matrix ${\mathbf{H}}_{2N-L,L+1}\coloneqq\left(f(\ell+m)\right)_{\ell,m=0}^{2N-L-1,L}$ and consider its factorization
$\mathbf{H}_{2N-L,L+1}=\mathbf{V}_{2N-L,M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{L+1,M}(\boldsymbol{z}).$ (2.2)
Since Vandermonde matrices in (2.2) have full column ranks, the matrix $\mathbf{H}_{2N-L,L+1}$ has rank $M$. Therefore in case of exact given data, the order $M$ of the exponential sum (2.1) is reconstructed as the numerical rank of the matrix $\mathbf{H}_{2N-L,L+1}$.
Consider further a Hankel matrix pencil problem
$z\mathbf{H}_{2N-L,L}(0)-\mathbf{H}_{2N-L,L}(1),$ (2.3)
where $\mathbf{H}_{2N-L,L}(s):=\mathbf{H}_{2N-L,L+1}(1:2N-L,1+s:L+s)=(f(\ell+k+s))_{k,\ell=0}^{2N-L-1,L-1}$ for $s=0,1$. Since the Hankel matrices $\mathbf{H}_{2N-L,L}(0)$ and $\mathbf{H}_{2N-L,L}(1)$ satisfy factorizations
$\mathbf{H}_{2N-L,L}(0)$ $=\mathbf{V}_{2N-L,M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{L,M}(\boldsymbol{z}),$ (2.4)
$\mathbf{H}_{2N-L,L}(1)$ $=\mathbf{V}_{2N-L,M}(\boldsymbol{z})\,\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{L,M}(\boldsymbol{z}),$ (2.5)
we obtain that
$z\,{\mathbf{H}}_{2N-L,L}(0)-{\mathbf{H}}_{2N-L,L}(1)=\mathbf{V}_{2N-L,M}(\boldsymbol{z})\,\mathrm{diag}\left(z-z_{j}\right)_{j=1}^{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{L,M}(\boldsymbol{z}).$
From the last property, we conclude that nodes $z_{j}$, $j=1,\ldots,M$ in (2.1) are eigenvalues of the matrix pencil (2.3). Let us denote by $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ the vector of eigenvalues of (2.3). In case when $M$ is known, we can choose $N=L=M$ and compute $z_{j}$ as eigenvalues of
$z\mathbf{H}_{M}(0)-\mathbf{H}_{M}(1).$ (2.6)
Now we discuss how we solve the matrix pencil (2.3). Consider the SVD of ${\mathbf{H}}_{2N-L,L+1}$,
${\mathbf{H}}_{2N-L,L+1}={\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,L+1}{\mathbf{W}}_{L+1},$ (2.7)
where ${\mathbf{U}}_{2N-L}$ and ${\mathbf{W}}_{L+1}$ are unitary square matrices and, in the case of unperturbed data,
${\mathbf{D}}_{2N-L,L+1}=\left(\begin{array}[]{cc}\mathrm{diag}\left((\sigma_{j})_{j=1}^{M}\right)&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{0}}\end{array}\right)\in{\mathbb{R}}^{(2N-L)\times(L+1)},$
where $\sigma_{1}\geq\ldots\geq\sigma_{M}$ denote the positive singular values of ${\mathbf{H}}_{2N-L,L+1}$. Thus (2.7) can be simplified to
${\mathbf{H}}_{2N-L,L+1}={\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1},$ (2.8)
where the (last) $L+1-M$ zero-columns in ${\mathbf{D}}_{2N-L,L+1}$ and the last $L+1-M$ rows in ${\mathbf{W}}_{L+1}$ are removed. Similarly as before, the SVDs of ${\mathbf{H}}_{2N-L,L}(0)$ and ${\mathbf{H}}_{2N-L,L}(1)$ can be directly derived from (2.8) and the matrix pencil in (2.3) now reads
$z\left({\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1}\right)(1:2N-L,1:L)$
$\hskip 142.26378pt-\left({\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1}\right)(1:2N-L,2:L+1),$
or equivalently
$z\left({\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1}\right)(1:2N-L,1:L)-\left({\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1}\right)(1:2N-L,2:L+1).$
Multiplication from the left with $\left(\mathrm{diag}(\sigma_{j}^{-1})_{j=1}^{M},{\mathbf{0}}\right)\in{\mathbb{R}}^{M\times(2N-L)}$ yields the matrix pencil
$z\,{\mathbf{W}}_{M,L+1}(1:M,1:L)-{\mathbf{W}}_{M,L+1}(1:M,2:L+1).$ (2.9)
Thus, eigenvalues of the matrix pencil (2.3) can be computed as eigenvalues of the matrix pencil (2.9). Then we define ${\mathbf{W}}_{M,L}(0)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,1\!:\!L)$, ${\mathbf{W}}_{M,L}(1)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,2\!:\!L+1)$, and determine the vector $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ as vector of eigenvalues of $\left({\mathbf{W}}_{M,L}(0)^{T}\right)^{\dagger}{\mathbf{W}}_{M,L}(1)^{T}$.
Having nodes $z_{j}$, we reconstruct coefficients $\gamma_{j}$, $j=1,\ldots,M$ as a least squares solution of the system of linear equations
$\sum\limits_{j=1}^{M}\gamma_{j}z_{j}^{k}=f(k),\quad k=0,\ldots,2N-1,$
which can be also written in a matrix form
[…]
3 Sensitivities of Hankel pencil eigenvalues
In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3). The concept of sensitivity arises from the perturbation of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [10] and further considered in [2]. In [21], the notion of sensitivities was defined for Loewner matrix pencil. In the paper [21], the authors define two sensitivities, which correspond to structured and unstructured perturbations. Using similar ideas, in [22], the author considered the structured sensitivity of Hankel matrix pencil.
3.1 Sensitivity of Hankel pencil eigenvalues with unstructured perturbations
Following approach offered in [21] for Loewner pencil, we will show that in case of unstructured perturbations of Hankel pencil, the sensitivity of eigenvalues depends on the numerical conditions of Vandermonde matrices from factorizations (2.4) and (2.5) (for $N=L=M$). We consider the generalized eigenvalue problem (GEP) for the Hankel matrix pencil (2.6),
$\boldsymbol{p}^{T}\mathbf{H}_{M}(1)=z\boldsymbol{p}^{T}\mathbf{H}_{M}(0),\quad\mathbf{H}_{M}(1)\boldsymbol{q}=z\mathbf{H}_{M}(0)\boldsymbol{q},$ (3.1)
where $\boldsymbol{p}=(p_{1},\ldots,p_{M})^{T}$ and $\boldsymbol{q}=(q_{1},\ldots,q_{M})^{T}$ are right and left eigenvectors of the GEP (3.1) respectively. Then under the perturbation
$\tilde{\mathbf{H}}_{M}(0):=\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)},\ \ \tilde{\mathbf{H}}_{M}(1):=\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)},$
the first order approximation of the eigenvalue perturbation $z^{(1)}$ is given by
$z^{(1)}=\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}.$ (3.2)
Formula (3.2) is obtained from the perturbed system
$(\boldsymbol{p}^{T}+(\boldsymbol{p}^{(1)})^{T}+\ldots)$ $(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)})=(z+z^{(1)}+\ldots)(\boldsymbol{p}^{T}+(\boldsymbol{p}^{(1)})^{T}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{{\mathbf{H}_{M}(0)}}),$
$(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)}))$ $(\boldsymbol{q}+\boldsymbol{q}^{(1)}+\ldots)=(z+z^{(1)}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)}))(\boldsymbol{q}+\boldsymbol{q}^{(1)}+\ldots),$
by retaining only first-order terms (see [10], [2], [21] and [22] for details).
Assume that the norm of perturbation matrices is bounded as follows
$\|\Delta_{\mathbf{H}_{M}(0)}\|_{2}\leq\epsilon\|\mathbf{H}_{M}(0)\|_{2},\ \ \|\Delta_{\mathbf{H}_{M}(1)}\|_{2}\leq\epsilon\|\mathbf{H}_{M}(1)\|_{2},$ (3.3)
where $\varepsilon>0$. Then from (3.2) taking into account (3.3), we get for the GEP (3.1)
$|z^{(1)}|$ $=\left|\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}_{L}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}\right|$
$\leq\frac{\|\boldsymbol{p}^{T}\|_{2}\,\|\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)}\|_{2}\,\|\boldsymbol{q}\|_{2}}{|\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}|}\leq\epsilon\rho,$
where $\rho$ is the sensitivity of the eigenvalue $z$ of the GEP (3.1) defined by
$\rho:=\frac{\|\boldsymbol{p}^{T}\|_{2}\,(|z|\,\|\mathbf{H}_{M}(0)\|_{2}+\|\mathbf{H}_{M}(1)\|_{2})\,\|\boldsymbol{q}\|_{2}}{|\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}|}.$ (3.4)
We derive further formulas for the exact computation and upper bounds of sensitivities $\rho_{j}$ of eigenvalues $z_{j}$, $j=1,\ldots,M$ of the GEP (3.1). Since Hankel matrices $\mathbf{H}_{M}(0)$ and $\mathbf{H}_{M}(1)$ satisfy factorizations (2.4) and (2.5) for $N=L=M$, left $\boldsymbol{p}^{(j)}$ and right $\boldsymbol{q}^{(j)}$ eigenvectors of the GEP (3.1) corresponding to the eigenvalue $z_{j}$ are given by formulas
$\boldsymbol{q}^{(j)}=\boldsymbol{p}^{(j)}=\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j},$ (3.5)
where $\boldsymbol{e}_{j}$ is the $j$th canonical vector in $\mathbb{R}^{M}$. Then using again the factorization (2.4) of the Hankel matrix $\mathbf{H}_{M}(0)$, we get
$(\boldsymbol{p}^{(j)})^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}^{(j)}$ $=\left(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\right)^{T}\,\mathbf{H}_{M}(0)\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}$
$=\boldsymbol{e}_{j}^{T}\,\mathbf{\Gamma}_{M}\,\boldsymbol{e}_{j}=\gamma_{j}.$ (3.6)
From (3.4) taking into account (3.6), we obtain that the sensitivity $\rho_{j}$ of the eigenvalue $z_{j}$ is given by
$\rho_{j}=\frac{1}{|\gamma_{j}|}\|(\boldsymbol{p}^{(j)})^{T}\|_{2}\,(|z_{j}|\,\|\mathbf{H}_{M}(0)\|_{2}+\|\mathbf{H}_{M}(1)\|_{2})\,\|\boldsymbol{q}^{(j)}\|_{2}.$ (3.7)
The ill-disposedness of the eigenvalue $z_{j}$ with respect to unstructured perturbations can be measured by checking whether $\rho_{j}$ is "large". Applying formulas (2.4), (2.5) and (3.5) to (3.7), we obtain for $j=1,\ldots,M$
$\rho_{j}$ $=\frac{1}{|\gamma_{j}|}\left\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\right\|_{2}^{2}\left(|z_{j}|\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}+\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}\right).$ (3.8)
The formula (3.8) is used to compute the exact values of sensitivities $\rho_{j}$.
We proceed with estimation of (3.7). Since from (2.4) and (2.5), we have
$\|\mathbf{H}_{M}(0)\|_{2}$ $\leq\|\mathbf{V}_{M}(\boldsymbol{z})\|_{2}\,\|\mathbf{\Gamma}_{M}\|_{2}\,\|\mathbf{V}^{T}_{M}(\boldsymbol{z})\|_{2},$
$\|\mathbf{H}_{M}(1)\|_{2}$ $\leq\|\mathbf{V}_{M}(\boldsymbol{z})\|_{2}\,\|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}\,\|\mathbf{V}^{T}_{M}(\boldsymbol{z})\|_{2},$ (3.9)
and taking into account the fact $\|\mathbf{X}\,\boldsymbol{e}_{j}\|_{2}\leq\|\mathbf{X}\|_{2}$, we obtain
$\rho_{j}\leq\frac{|z_{j}|\,\|\mathbf{\Gamma}_{M}\|_{2}+\|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}}{|\gamma_{j}|}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right),$
where $\kappa(\mathbf{X}):=\|\mathbf{X}\|_{2}\,\|\mathbf{X}^{\dagger}\|_{2}$ denotes the condition number of the matrix $\mathbf{X}$. For diagonal matrices $\mathbf{\Gamma}_{M}$ and $\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}$, it holds
$\|\mathbf{\Gamma}_{M}\|_{2}=\max\limits_{k=1,\ldots,M}|\gamma_{k}|,\ \ \ \|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}=\max\limits_{k=1,\ldots,M}|z_{k}\gamma_{k}|.$
Thus
$\rho_{j}\leq\zeta_{j}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right),$ (3.10)
where $\zeta_{j}$ is defined by
$\zeta_{j}:=\frac{1}{|\gamma_{j}|}\left(|z_{j}|\max\limits_{k=1,\ldots,M}|\gamma_{k}|+\max\limits_{k=1,\ldots,M}|\gamma_{k}z_{k}|\right).$ (3.11)
Define further a vector of unstructured sensitivities $\boldsymbol{\rho}:=(\rho_{1},\ldots,\rho_{M})^{T}\in\mathbb{R}^{M}$ and let $\boldsymbol{\zeta}:=(\zeta_{1},\ldots,\zeta_{M})^{T}\in\mathbb{R}^{M}$. Then from (3.10) for the $\ell_{2}$-norm of the vector $\boldsymbol{\rho}$ we get
$\|\boldsymbol{\rho}\|_{2}\leq\|\boldsymbol{\zeta}\|_{2}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right).$
From (3.8), taking into account (3.9) and (3.11), we obtain
$\rho_{j}\leq\zeta_{j}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}\,\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}.$
By summing up the last inequalities over $j=1,\ldots,M$, we get
$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}$
where $\zeta_{\max}:=\max_{j=1,\ldots,M}\zeta_{j}$. Finally using the property $\sum_{j=1}^{M}\|\mathbf{X}\boldsymbol{e}_{j}\|_{2}^{2}=\|\mathbf{X}\|_{\mathrm{F}}^{2}$, we obtain
$\|\boldsymbol{\rho}\|_{1}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\big{\|}\mathbf{V}^{\dagger}_{M}(\boldsymbol{z})\big{\|}_{\mathrm{F}}^{2}$
for the $\ell_{1}$-norm of the sensitivity vector $\boldsymbol{\rho}$.
3.2 Sensitivity Hankel pencil eigenvalues with structured perturbations
In this section we consider structured perturbations of the GEP (3.1) for the matrix pencil (2.6), i.e. the perturbation matrices have a particular structure, they are again Hankel matrices. We give a short description of the results obtained in [22]. Let $\boldsymbol{\delta}_{2M}:=(\delta_{0},\delta_{1},\ldots,\delta_{2M-1})^{T}$ be a vector of perturbations and we define
$\Delta_{\mathbf{H}_{M}(0)}:=(\delta_{k+\ell})_{k,\ell=0}^{M-1,M-1},\ \Delta_{\mathbf{H}_{M}(1)}:=(\delta_{k+\ell+1})_{k,\ell=0}^{M-1,M-1}.$
For the first order approximation of the eigenvalue perturbation $z_{j}^{(1)}$, $j=1,\ldots,M$ from (3.2) using (3.6), we obtain
$z^{(1)}_{j}=\frac{1}{\gamma_{j}}\,(\boldsymbol{p}^{(j)})^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z_{j}\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}^{(j)}.$ (3.12)
Define further matrices $\mathbf{Q}^{(j)}_{M,2M}(0):=\left(Q^{(j)}_{k,m}(0)\right)_{k,m=1}^{M,2M}$ and $\mathbf{Q}^{(j)}_{M,2M}(1):=\left(Q^{(j)}_{k,m}(1)\right)_{k,m=1}^{M,2M}$ by their entries
$Q^{(j)}_{k,m}(0):=\begin{cases}q^{(j)}_{m-k+1},&1\leq m-k+1\leq M,\\ 0,&\text{otherwise},\end{cases}$
and
$Q^{(j)}_{k,m}(1):=\begin{cases}q^{(j)}_{m-k},&1\leq m-k\leq M,\\ 0,&\text{otherwise}.\end{cases}$
Then
$\Delta_{\mathbf{H}_{M}(0)}\,\boldsymbol{q}^{(j)}=\mathbf{Q}_{M,2M}^{(j)}(0)\,\boldsymbol{\delta}_{2M},\ \Delta_{\mathbf{H}_{M}(1)}\,\boldsymbol{q}^{(j)}=\mathbf{Q}_{M,2M}^{(j)}(1)\,\boldsymbol{\delta}_{2M}.$
Since for $m=1,\ldots,2M$
$\left((\boldsymbol{p}^{(j)})^{T}\,\mathbf{Q}^{(j)}_{M,2M}(0)\right)_{m}$ $=\sum\limits_{k=1}^{2M}p^{(j)}_{k}\,Q^{(j)}_{k,m}(0)=\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}p^{(j)}_{k}\,q^{(j)}_{m-k+1},$
$\left((\boldsymbol{p}^{(j)})^{T}\,\mathbf{Q}_{M,2M}^{(j)}(1)\right)_{m}$ $=\sum\limits_{k=1}^{M}p^{(j)}_{k}\,Q_{k,m}^{(j)}(1)=\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}p^{(j)}_{k}\,q^{(j)}_{m-k},$
taking into account (3.12), we define vectors $\boldsymbol{\mathcal{S}}^{(j)}_{2M}:=\left(S^{(j)}_{1},\ldots,S^{(j)}_{2M}\right)^{T}\in\mathbb{C}^{2M}$, $j=1,\ldots,M$ by their components
$S^{(j)}_{m}:=\frac{1}{\gamma_{j}}\left(\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}p^{(j)}_{k}\,q^{(j)}_{m-k}-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}p^{(j)}_{k}\,q^{(j)}_{m-k+1}\right),$
$m=1,\ldots,2M$. Then the first order approximation of the eigenvalue perturbation $z_{j}^{(1)}$ in (3.12) can be given by
$z_{j}^{(1)}=(\boldsymbol{\mathcal{S}}^{(j)}_{2M})^{T}\,\boldsymbol{\delta}_{2M},\quad j=1,\ldots,M.$ (3.13)
The component $S^{(j)}_{m}$ shows the influence of the perturbation of the $m$th measurement (in this case it is the function value $f(m-1)$) on the $j$th eigenvalue $z_{j}$ of the Hankel matrix pencil (2.6). Using (3.5), we rewrite components $S^{(j)}_{m}$ in the following way
$S^{(j)}_{m}=$ $\frac{1}{\gamma_{j}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{m-k,j}$
$-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{m-k+1,j}\bigg{)}.$
Formula (3.13) can be also written in a matrix form
$\boldsymbol{z}^{(1)}=\mathbf{\mathcal{S}}_{M,2M}\,\boldsymbol{\delta}_{2M},$
where $\boldsymbol{z}^{(1)}:=(z_{1}^{(1)},\ldots,z_{M}^{(1)})^{T}$ and $\mathbf{\mathcal{S}}_{M,2M}:=(S_{j,m})_{j,m=1}^{M,2M}\in\mathbb{C}^{M\times 2M}$ be a matrix with the entries $S_{j,m}:=S^{(j)}_{m}$ that quantify the perturbation of the eigenvalues of matrix pencil (2.6).
Finally, we define a structured sensitivity of the eigenvalue $z_{j}$ of matrix pencil (2.6) with respect to the $m$th measurement by
$\eta_{j,m}:=|S_{j,m}|$
for $j=1,\ldots,M$ and $m=1,\ldots,2M$. Then the eigenvalue sensitivity $\eta_{j}$ for the eigenvalue $z_{j}$ with respect to structured perturbations of Hankel matrix pencil (2.6) is defined as the $\ell_{2}$-norm of the vector $\boldsymbol{\mathcal{S}}^{(j)}_{2M}\in\mathbb{C}^{2M}$. Thus we further introduce a vector $\boldsymbol{\eta}:=(\eta_{1},\ldots,\eta_{M})^{T}\in\mathbb{R}^{M}$ with its entries
$\eta_{j}:=\|\boldsymbol{\mathcal{S}}^{(j)}_{2M}\|_{2}=\sqrt{\eta_{j,1}^{2}+\ldots+\eta^{2}_{j,2M}},\quad j=1,\ldots,M.$
The ill-disposedness of the eigenvalue $z_{j}$ with respect to structured perturbations can be measured by checking whether $\eta_{j}$ is "large". The entry $\eta_{j}$ determines also the standard deviation of the eigenvalue perturbation $z_{j}^{(1)}$ (see [22] for details).
4 Recovery of rational functions via Hankel matrix pencil method and sensitivities of the poles
In this section, we describe the main ideas of our new method for the rational reconstruction and present analysis of the sensitivities of the poles.
4.1 Description of the method
Let $r(t)$ be a rational function defined by (1.1) and by $\mathbb{D}$ we denote the unit circle. Following approach in [19], for $k\in\mathbb{Z}$ we consider the integral
$\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t},$
which is closely related to the classical Fourier coefficient of a function $r(\mathrm{e}^{\mathrm{i}\varphi})$,
$\widehat{r}_{k}:=\frac{1}{2\pi}\int\limits_{0}^{2\pi}r(\mathrm{e}^{\mathrm{i}\varphi})\mathrm{e}^{-\mathrm{i}k\varphi}\,\mathrm{d}\varphi=\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t}.$
Then (see, for example, [19])
$\widehat{r}_{k}=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{-(k+1)},&k\leq-1,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{-(k+1)},&k\geq 1.\end{cases}$ (4.1)
From (4.1), we conclude that Fourier coefficients $\widehat{r}_{k}$ with negative indeces $k=-1,-2,\ldots$ contain information about poles inside the unite circle, and Fourier coefficients $\widehat{r}_{k}$ with positive indises $k=1,2,\ldots$ contain information about poles outside the unite circle.
Taking into account (4.1), we define an exponential sum $f(t)$ by
$f(t)=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{t},&t\geq 0,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{t},&t\leq-2,\end{cases}$ (4.2)
and from (4.1) we obtain the following interpolation property
$f(-(k+1))=\widehat{r}_{k},\quad k\in\mathbb{Z}\setminus\{0\}.$ (4.3)
Therefore the rational reconstruction problem (1.1) can be reformulated as the exponential interpolation problem (4.3). Our further goal is to reconstruct an exponential sum (4.2), using samples $f(m)$, $m\in\mathbb{Z}\setminus\{-1\}$ as input data. From (4.2), we get that values $f(m)$ with $m=0,1,2,\ldots$ contain information about nodes $z_{j}$ inside $\mathbb{D}$, and values $f(m)$ with $m=-2,-3,\ldots$ contain information about nodes $z_{j}$ outside $\mathbb{D}$.
Let $M_{1}$ and $M_{2}$ be numbers of nodes $z_{j}$ of the exponential sum (4.2) located inside $\mathbb{D}$ and outside $\mathbb{D}$ respectively. We assume that a number of poles $M$ as well as $M_{1}$ and $M_{2}$ are also unknown. Let $L,N\in\mathbb{N}$ and $M_{1},M_{2}\leq L\leq N$, where $L$ is the upper bound for $M$. We use values $f(k)$, $k=0,\ldots,2N-1$ ($\widehat{r}_{-k}$, $k=1,\ldots,2N$) to reconstruct nodes $z_{j}$ inside $\mathbb{D}$, and values $f(-k)$, $k=2,\ldots,2N+1$ ($\widehat{r}_{k}$, $k=1,\ldots,2N$) to reconstruct nodes $z_{j}$ outside $\mathbb{D}$. Further we rewrite an exponential sum (4.2) in the following way
$f(t)=\begin{cases}\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{j_{n}}(z^{(-)}_{j_{n}})^{t},&t\geq 0,\\ -\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{j_{n}}(z^{(+)}_{j_{n}})^{t}&t\leq-2,\end{cases}$ (4.4)
i.e. by $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ we denote poles located inside $\mathbb{D}$ and let $\gamma^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ be the corresponding coefficients and by $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$ we denote poles located outside $\mathbb{D}$ and let $\gamma^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$ be the corresponding coefficients.
To reconstruct exponential sums (4.4), we use the ESPRIT method described in Section 2. We present some modification of this method for our special parameter estimation problem (4.4). Note that the ESPRIT is not the only method to apply in the frequency domain. We can use the other methods for exponential recovery, for example, MPM (matrix-pencil method) [15] or approximate Prony method [16]. In [19], the author uses the AAK theory [14] in the frequency domain to reconstruct (1.1) and a matrix-valued version of (1.1).
Let $\boldsymbol{z}^{(-)}\in\mathbb{C}^{M_{1}}$ and $\boldsymbol{z}^{(+)}\in\mathbb{C}^{M_{2}}$ be two vectors defined by
$\boldsymbol{z}^{(-)}:=(z^{(-)}_{j_{1}},\ldots,z^{(-)}_{j_{M_{1}}}),\ \ \boldsymbol{z}^{(+)}:=\left(\frac{1}{z_{j_{1}}^{(+)}},\ldots,\frac{1}{z^{(+)}_{j_{M_{2}}}}\right).$
Let us define also diagonal matrices
$\mathbf{\Gamma}^{(-)}_{M_{1}}:=\mathrm{diag}(\gamma^{(-)}_{j_{n}})_{n=1}^{M_{1}},\ \mathbf{\Gamma}^{(+)}_{M_{2}}:=\mathrm{diag}(\gamma^{(+)}_{j_{n}})_{n=1}^{M_{2}},\ \mathbf{Z}_{M_{1}}^{(+)}:=\mathrm{diag}\left(z^{(-)}_{j_{n}}\right)_{n=1}^{M_{1}},\ \mathbf{Z}_{M_{2}}^{(+)}:=\mathrm{diag}\left(\frac{1}{z^{(+)}_{j_{n}}}\right)_{n=1}^{M_{2}}.$
According to the ESPRIT method described in Section 2, to reconstruct nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$, located inside the unit circle, we use the Hankel matrix
$\mathbf{H}^{(-)}_{2N-L,L+1}:=\left(f(\ell+k)\right)_{k,\ell=0}^{2N-L-1,L}=\left(\widehat{r}_{-(k+\ell)}\right)_{k=0,\ell=1}^{2N-L-1,L+1},$
and to reconstruct nodes $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$, located outside the unit circle, we use the Hankel matrix
$\mathbf{H}^{(+)}_{2N-L,L+1}:=(f(-(\ell+k+2)))_{k,\ell=0}^{2N-L-1,L}=\left(\widehat{r}_{k+\ell}\right)_{k=0,\ell=1}^{2N-L-1,L+1}.$
Consider factorizations of these matrices
$\mathbf{H}^{(-)}_{2N-L,L+1}$ $=\mathbf{V}_{2N-L,M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}^{(-)}_{M_{1}}\,\mathbf{V}^{T}_{L+1,M_{1}}(\boldsymbol{z}^{(-)}),$ (4.5)
$\mathbf{H}^{(+)}_{2N-L,L+1}$ $=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L+1,M_{2}}(\boldsymbol{z}^{(+)}).$ (4.6)
Since all Vandermonde matrices in (4.5) and (4.6) have full column ranks, the matrix $\mathbf{H}_{2N-L,L+1}^{(-)}$ has rank $M_{1}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ has rank $M_{2}$. Therefore in case of exact given data, the numbers $M_{1}$ and $M_{2}$ are reconstructed as numerical ranks of matrices $\mathbf{H}_{2N-L,L+1}^{(-)}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ respectively.
First we consider the reconstruction of nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$, located inside the $\mathbb{D}$. Since in this case, the Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ have structure similar to the structure of the Hankel matrices $\mathbf{H}_{2N-L,L+1}$ in the Section 2, following ideas of the Section 2, we get that nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ are reconstructed as eigenvalues of the Hankel matrix pencil
$z\mathbf{H}^{(-)}_{2N-L,L}(0)-\mathbf{H}^{(-)}_{2N-L,L}(1),$ (4.7)
where for $s=0,1$
$\mathbf{H}^{(-)}_{2N-L,L}(s):=\mathbf{H}^{(-)}_{2N-L,L+1}(1:2N-L,1+s:L+s)=(f(\ell+k+s))_{k,\ell=0}^{2N-L-1,L-1}.$
Consider now the case of nodes $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$, located outside the $\mathbb{D}$. For $s=0,1$, define matrices
$\mathbf{H}^{(+)}_{2N-L,L}(s):=\mathbf{H}^{(+)}_{2N-L,L+1}(1:2N-L,1+s:L+s)=(f(-(\ell+k+s+2)))_{k,\ell=0}^{2N-L-1,L-1},$
and consider their factorizations
$\mathbf{H}^{(+)}_{2N-L,L}(0)$ $=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}),$ (4.8)
$\mathbf{H}^{(+)}_{2N-L,L}(1)$ $=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{3}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}).$ (4.9)
Since according to (4.8) and (4.9) we have
$z\,{\mathbf{H}}^{(+)}_{2N-L,L}(0)-{\mathbf{H}}^{(+)}_{2N-L,L}(1)=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathrm{diag}\left(z-\frac{1}{z^{(+)}_{j_{n}}}\right)_{n=1}^{M_{2}}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}),$
then inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of nodes $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ are reconstructed as eigenvalues of the Hankel matrix pencil
$z\mathbf{H}^{(+)}_{2N-L,L}(0)-\mathbf{H}^{(+)}_{2N-L,L}(1).$ (4.10)
If we assume that $M_{1}$ and $M_{2}$ are known, we can reconstruct nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ as eigenvalues of the Hankel matrix pencil ($N=L=M_{1}$ in (4.7))
$z\mathbf{H}^{(-)}_{M_{1}}(0)-\mathbf{H}^{(-)}_{M_{1}}(1),$ (4.11)
and inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of nodes $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ are reconstructed as eigenvalues of the Hankel matrix pencil ($N=L=M_{2}$ in (4.10))
$z\mathbf{H}^{(+)}_{M_{2}}(0)-\mathbf{H}^{(+)}_{M_{2}}(1).$ (4.12)
Having nodes $z^{(-)}_{j_{n}}$ and $z^{(+)}_{j_{n}}$, coefficients $\gamma^{(-)}_{j_{n}}$ and $\gamma^{(+)}_{j_{n}}$ can be computed as least squares solutions of the overdetermined systems of linear equations
$\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{j_{n}}(z^{(-)}_{j_{n}})^{-(k+1)}=\widehat{r}_{-k},\quad k=1,\ldots,2N,$
and
$-\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{j_{n}}(z^{(+)}_{j_{n}})^{-(k+1)}=\widehat{r}_{k},\quad k=1,\ldots,2N,$
respectively.
Remark 4.1.
Note that coefficients $\widehat{r}_{k}$ can be also computed applying the quadrature formulas on some grid of points. In that case, we will use the function values instead of Fourier coefficients as input data for the algorithm. This question was discussed in details in [19], where the author consider the random access model and the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$. In this paper, we are interested in the presentation of a modification of the method from [19] and in the explanation of the results of this method regarding recovery of noisy data with the use of sensitivities.
4.2 Sensitivities of poles
In this subsection, we consider sensitivities of poles $z_{j}$ of the reconstruction problem (1.1) that is solved with the method described in Subsection 4.1. First we consider the case of unstructured perturbations. Poles $z_{j_{n}}^{(-)}$, $n=1,\ldots,M_{1}$ located inside the unit circle $\mathbb{D}$, are reconstructed as eigenvalues of the matrix pencil (4.11). Since Hankel matrices in the matrix pencil (4.11) have similar structure as Hankel matrices in the matrix pencil (2.6), we apply ideas from Subsection 3.1 without any changes and obtain for the sensitivities $\rho_{n}^{(-)}$, $n=1,\ldots,M_{1}$ of the poles $z^{(-)}_{j_{1}},\ldots,z^{(-)}_{j_{M_{1}}}$ in the $\mathbb{D}$ the formula for the exact computation
$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
×$\times\left(|z_{j_{n}}^{(-)}|\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}+\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{Z}_{M_{1}}\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}\right),$ (4.13)
where $\boldsymbol{e}_{n}$ is the $n$th canonical (column) vector in $\mathbb{R}^{M_{1}}$ and the upper bound
$\rho_{n}^{(-)}\leq\zeta_{n}^{(-)}\,\kappa^{2}\left(\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\right),$ (4.14)
with
$\zeta_{n}^{(-)}:=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left(\left|z_{j_{n}}^{(-)}\right|\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{j_{k}}^{(-)}\right|+\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{j_{k}}^{(-)}z_{j_{k}}^{(-)}\right|\right).$
Now consider the case of the sensitivities $\rho_{n}^{(+)}$ of poles $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ located outside the unit circle $\mathbb{D}$. Inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of poles $z^{(+)}_{j_{n}}$ are reconstructed as eigenvalues of the matrix pencil (4.12). That means that we will compute sensitivities $\rho_{n}^{(+)}$ of the inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of the poles $z^{(+)}_{j_{n}}$, but we will say that $\rho_{n}^{(+)}$ are also sensitivities of poles $z^{(+)}_{j_{n}}$.
To present formulas for sensitivities $\rho_{n}^{(+)}$, we have to apply some modifications to the results of the Subsection 3.1. Matrices $\mathbf{H}^{(+)}_{M_{2}}(0)$ and $\mathbf{H}^{(+)}_{M_{2}}(1)$ satisfy factorizations
$\mathbf{H}^{(+)}_{M_{2}}(0)=-\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)}),$ (4.15)
$\mathbf{H}^{(+)}_{M_{2}}(1)=-\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)}).$ (4.16)
Thus left and right eigenvectors of the GEP of the Hankel matrix pencil (4.12) corresponding to the eigenvalue $\frac{1}{z^{(+)}_{j_{n}}}$ are given by formulas
$\boldsymbol{q}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n},\quad\boldsymbol{p}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n},$ (4.17)
where $\boldsymbol{e}_{n}$ is the $n$th canonical (column) vector in $\mathbb{R}^{M_{2}}$ Taking into account (4.15) and (4.17), we get
$(\boldsymbol{p}^{(n)})^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\boldsymbol{q}^{(n)}$ $=\left(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}\right)^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}$
$=-\boldsymbol{e}_{n}^{T}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\boldsymbol{e}_{n}=-\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}.$ (4.18)
From (3.4) applying (4.15)–(4.18), we obtain the formula for the exact computation of sensitivities $\rho_{n}^{(+)}$ of the poles $z^{(+)}_{j_{n}}$ (the eigenvalues $\frac{1}{z^{(+)}_{j_{n}}}$ of the matrix pencil (4.12)) located outside the unit circle
$\rho_{n}^{(+)}=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}|}\bigg{\|}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}\bigg{\|}_{2}^{2}$
×$\times\bigg{(}\frac{1}{|z_{j_{n}}^{(+)}|}\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{2}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}+\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}\bigg{)},$ (4.19)
and the upper bound
$\rho^{(+)}_{n}\leq\frac{\frac{1}{|z_{j_{n}}^{(+)}|}\,\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big{\|}_{2}+\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big{\|}_{2}}{|\gamma_{j_{n}}^{(+)}\,(z^{(+)}_{j_{n}})^{-2}|}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right).$
Now since
$\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big{\|}_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-2}|,\ \ \ \big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big{\|}_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-3}|,$
we get for $n=1,\ldots,M_{2}$
$\rho_{n}^{(+)}\leq\zeta_{n}^{(+)}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right),$ (4.20)
with
$\zeta_{n}^{(+)}:=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z^{(+)}_{j_{n}})^{-2}|}\left(\frac{1}{|z_{j_{n}}^{(+)}|}\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-2}|+\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-3}|\right).$
Define further two vectors of unstructured sensitivities $\boldsymbol{\rho}^{(-)}:=(\rho_{1}^{(-)},\ldots,\rho^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\rho}^{(+)}:=(\rho_{1}^{(+)},\ldots,\rho^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$, and let $\boldsymbol{\zeta}^{(-)}:=(\zeta_{1}^{(-)},\ldots,\zeta^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\zeta}^{(+)}:=(\zeta_{1}^{(+)},\ldots,\zeta^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$. Then using (4.14) and (4.20), we write the following upper bounds for the $\ell_{2}$-norm of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
$\|\boldsymbol{\rho}^{(-)}\|_{2}\leq\|\boldsymbol{\zeta}^{(-)}\|_{2}\,\kappa^{2}\left(\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\right),\quad\|\boldsymbol{\rho}^{(+)}\|_{2}\leq\|\boldsymbol{\zeta}^{(+)}\|_{2}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right).$ (4.21)
Using ideas similar as in Subsection 3.1, we write also the upper bounds for the $\ell_{1}$-norm of the sensitivity vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{\dagger}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2}$ (4.22)
$\|\boldsymbol{\rho}^{(+)}\|_{1}$ $\leq\zeta^{(+)}_{\max}\,\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{2}^{2}\,\|\mathbf{V}^{\dagger}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{\mathrm{F}}^{2},$ (4.23)
where $\zeta_{\max}^{(-)}:=\max_{j=1,\ldots,M_{1}}\zeta_{j}^{(-)}$ and $\zeta_{\max}^{(+)}:=\max_{j=1,\ldots,M_{2}}\zeta_{j}^{(+)}$.
Let us now consider the case of structured perturbations. Define further matrices
$\mathbf{\mathcal{S}}^{(-)}_{M_{1},2M_{1}}:=\big{(}S_{n,m}^{(-)}\big{)}_{n,m=1}^{M_{1},2M_{1}}\in\mathbb{C}^{M_{1}\times 2M_{1}},\quad\mathbf{\mathcal{S}}^{(+)}_{M_{2},2M_{2}}:=\big{(}S_{n,m}^{(+)}\big{)}_{n,m=1}^{M_{2},2M_{2}}\in\mathbb{C}^{M_{2}\times 2M_{2}},$
and vectors
$(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{1}})^{(-)}$ $:=\left((S^{(n)}_{1})^{(-)},\ldots,(S^{(n)}_{2M_{1}})^{(-)}\right)^{T}\in\mathbb{C}^{2M_{1}}\quad\text{for}\quad n=1,\ldots,M_{1},$
$(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{2}})^{(+)}$ $:=\left((S^{(n)}_{1})^{(+)},\ldots,(S^{(n)}_{2M_{2}})^{(+)}\right)^{T}\in\mathbb{C}^{2M_{2}}\quad\text{for}\quad n=1,\ldots,M_{2},$
by their components
$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}:=$ $\frac{1}{\gamma_{j_{n}}^{(-)}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{m-k,n}$
$-z_{j_{n}}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{m-k+1,n}\bigg{)}$ (4.24)
and
$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}:=$ $-\frac{1}{\gamma_{j_{n}}^{(+)}\,z_{j_{n}}^{-2}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{m-k,n}$
$-\frac{1}{z_{j_{n}}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{m-k+1,n}\bigg{)}.$ (4.25)
Formulas (4.24) and (4.25) are obtained using similar ideas from Subsection 3.2. In (4.25), we take into account (4.18).
Now we define a structured sensitivity of the pole $z^{(-)}_{j_{n}}$ located inside the unit circle with respect to the Fourier coefficient $\hat{r}_{-m}$ by $\eta_{n,m}^{(-)}:=|S_{n,m}^{(-)}|$ for $n=1,\ldots,M_{1}$ and $m=1,\ldots,2M_{1}$. Similarly, by $\eta_{n,m}^{(+)}:=|S_{n,m}^{(+)}|$ we define a structured sensitivity of the pole $z^{(+)}_{j_{n}}$ located outside the unit circle with respect to the Fourier coefficient $\hat{r}_{m}$ for $n=1,\ldots,M_{2}$ and $m=1,\ldots,2M_{2}$. Note that Fourier coefficients $\hat{r}_{-m}$ and $\hat{r}_{m}$ are $m$th measurements in this case. We define also two vectors
$\boldsymbol{\eta}^{(-)}:=(\eta_{1}^{(-)},\ldots,\eta_{M_{1}}^{(-)})^{T}\in\mathbb{R}^{M_{1}},\quad\boldsymbol{\eta}^{(+)}:=(\eta_{1}^{(+)},\ldots,\eta_{M_{2}}^{(+)})^{T}\in\mathbb{R}^{M_{2}},$
where the entries
$\eta_{n}^{(-)}:=\|(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{1}})^{(-)}\|_{2}=\sqrt{(\eta^{(-)}_{n,1})^{2}+\ldots+(\eta^{(-)}_{n,2M_{1}})^{2}},\quad n=1,\ldots,M_{1},$ (4.26)
and
$\eta_{n}^{(+)}:=\|(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{2}})^{(+)}\|_{2}=\sqrt{(\eta^{(+)}_{n,1})^{2}+\ldots+(\eta^{(+)}_{n,2M_{2}})^{2}},\quad n=1,\ldots,M_{2},$ (4.27)
are structured sensitivities for the poles $z_{j_{n}}^{(-)}$ and $z_{j_{n}}^{(+)}$ respectively.
5 Numerical experiments
In this section, we present a few numerical experiments. The algorithm is implemented in Matlab and use IEEE standard floating point arithmetic with double precision. We use the notations $\tilde{z}_{j}$ and $\tilde{\gamma}_{j}$ for the poles and the coefficients, respectively, reconstructed by the method considered in Subsection 4.1. We define the errors for the poles $z_{j}$ and the coefficients $\gamma_{j}$ by the following formulas
[…]
Note that in the case when we deal with noisy data, we assume that a number of poles $M$, as well as numbers of poles located inside and outside the unit circle $M_{1}$ and $M_{2}$ respectively are known.
Table 1: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.1
$\sigma=10^{-3}$ $N=2$ $N=10$
$z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$ $z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$
min $2.54$e–$05$ $2.06$e–$04$ $3.49$e–$06$ $5.56$e–$05$
max $2.85$e–$04$ $3.34$e–$03$ $2.7$e–$04$ $3.15$e–$03$
average $1.34$e–$04$ $2.03$e–$03$ $1.29$e–$04$ $1.06$e–$03$
$\sigma=10^{-1}$ $N=10$ $N=30$
$z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$ $z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$
min $3.2$e–$03$ $4.09$e–$02$ $7.19$e–$03$ $4.15$e–$02$
max $2.28$e–$02$ $3.65$e–$01$ $2.72$e–$02$ $1.7$e–$01$
average $1.09$e–$02$ $1.57$e–$01$ $1.36$e–$02$ $1.05$e–$01$
We employ $2N$ (noisy) Fourier coefficients $\hat{r}_{-k}(1+n_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+n_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(+)}$. By $(n_{k})_{k=1}^{2N-1}$ we denote the random noise that is generated in Matlab by $\sigma$*randn(2N), i.e., we take multiplicative noise drawn from the standard normal distribution. We take $L=N$ and consider different values of $N$ and $\sigma$.
To compute sensitivities $\rho_{j}^{(-)}$, $\rho_{j}^{(+)}$ and ,$\eta_{j}^{(-)}$, $\eta_{j}^{(+)}$, we use formulas (4.13), (4.19) and (4.26), (4.27) respectively. Upper bounds for sensitivities $\rho_{j}^{(-)}$ and $\rho_{j}^{(+)}$ are computed via (4.14) and (4.20) respectively.
[…]
$\sigma=10^{-3}$ $N=2$ $N=10$ $\sigma=10^{-1}$ $N=10$ $N=30$
min $1.78$e–$04$ $1.17$e–$04$ min $4.03$e–$02$ $2.77$e–$02$
max $2.11$e–$03$ $2.46$e–$03$ max $3.04$e–$01$ $1.44$e–$01$
average $1.44$e–$03$ $8.48$e–$04$ average $1.11$e–$01$ $9.55$e–$02$
Example 5.1.
Let a rational function $r(z)$ in (1.1) be given by parameters $M=2$ and
$\boldsymbol{z}=(-0.1,-2.1),\quad\boldsymbol{\gamma}=(0.5,0.5)$
(see [5, 21]). When $\sigma=0$, i.e. in case of recovery from exact data, we get the following reconstruction errors with $N=2$,
[…]
For $\sigma=10^{-3}$ and $\sigma=10^{-1}$, we compute 10 iterations and present the results in Table 1. Since we are interested to investigate which pole is more sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ separately. Results for recovery of the coefficient vector are presented in Table 2.
Table 3: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.2 for $\sigma=10^{-3}$
$N=8$ $z_{1}^{(-)}=0.9$ $z_{2}^{(-)}=-0.9$ $z_{3}^{(-)}=0.9\mathrm{i}$ $z_{4}^{(-)}=-0.9\mathrm{i}$
min $7.08$e–$05$ $4.31$e–$06$ $2.79$e–$06$ $9.06$e–$06$
max $6.74$e–$04$ $3.17$e–$04$ $1.65$e–$04$ $1.45$e–$04$
average $3.18$e–$04$ $1.48$e–$04$ $7.87$e–$05$ $5.57$e–$05$
$N=8$ $z_{1}^{(+)}=1.1$ $z_{2}^{(+)}=-1.1$ $z_{3}^{(+)}=1.1\mathrm{i}$ $z_{4}^{(+)}=-1.1\mathrm{i}$
min $1.71$e–$05$ $1.13$e–$06$ $1.31$e–$05$ $8.74$e–$06$
max $2.86$e–$04$ $2.7$e–$04$ $1.89$e–$04$ $1.9$e–$04$
average $1.44$e–$04$ $1.19$e–$04$ $1.04$e–$04$ $9.43$e–$05$
Analyzing numerical results in Table 1, we conclude that for $\sigma=0.1$ our method still reconstructs location of poles quite good. On the other hand, using more (noisy) Fourier coefficients also gives us better recovery results. We notice that a pole $z_{1}^{(-)}=-0.1$ is less sensitive to perturbations than a pole $z_{1}^{(+)}=-2.1$. This numerical observation is also confirmed with sensitivities values. For a pole $z_{1}^{(-)}=-0.1$, we have $\rho_{1}^{(-)}=2\cdot 10^{-1}$ and $\eta_{1}^{(-)}=1.97$, and for a pole $z_{1}^{(+)}=-2.1$, we have $\rho_{1}^{(+)}=9.52\cdot 10^{-1}$ and $\eta_{1}^{(+)}=6.65$. Both unstructured and structured sensitivities are larger for a pole $z_{1}^{(+)}=-2.1$ located outside the unit circle. We compute also vectors of structured perturbations
$(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T}.$
That means that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ with respect to perturbation of the Fourier coefficient $\hat{r}_{2}$.
Table 4: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.2 for $\sigma=10^{-1}$
$N=8$ $z_{1}^{(-)}=0.9$ $z_{2}^{(-)}=-0.9$ $z_{3}^{(-)}=0.9\mathrm{i}$ $z_{4}^{(-)}=-0.9\mathrm{i}$
min $6.41$e–$03$ $2.65$e–$03$ $5.14$e–$04$ $1.38$e–$03$
max $2.02$e–$01$ $2.37$e–$02$ $2.89$e–$02$ $7.51$e–$03$
average $3.99$e–$02$ $1.13$e–$02$ $1.06$e–$02$ $1.82$e–$02$
$N=8$ $z_{1}^{(+)}=1.1$ $z_{2}^{(+)}=-1.1$ $z_{3}^{(+)}=1.1\mathrm{i}$ $z_{4}^{(+)}=-1.1\mathrm{i}$
min $8.17$e–$04$ $8.71$e–$04$ $1.93$e–$03$ $2.83$e–$03$
max $2.95$e–$02$ $2.61$e–$02$ $2.48$e–$02$ $2.41$e–$02$
average $1.56$e–$02$ $1.4$e–$02$ $1.24$e–$02$ $1.15$e–$02$
Example 5.2.
We consider an example when poles are located very close to the unit circle. Let a rational function $r(z)$ of type $(7,8)$ in (1.1), i.e with $M=8$, be given by vectors of poles and coefficients
$\boldsymbol{z}=(0.9,-0.9,0.9\mathrm{i},-0.9\mathrm{i},1.1,-1.1,1.1\mathrm{i},-1.1\mathrm{i}),\quad\boldsymbol{\gamma}=(1,2,3,4,5,6,7,8),$
respectively. We consider first recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we get the following reconstruction errors
$e(\boldsymbol{z})=1.45\cdot 10^{-15},\quad e(\boldsymbol{\gamma})=5.48\cdot 10^{-14}.$
[…]
$N=8$ $\sigma=10^{-3}$ $\sigma=10^{-1}$
min $4.2$e–$03$ $4.25$e–$01$
max $1.73$e–$02$ $1.22$e–$00$
average $9.7$e–$03$ $6.35$e–$01$
In case of recovery from noisy data, we compute 10 iterations and present recovery result for each pole $z_{j}$ in Tables 3 and 4. Analyzing numerical results, we conclude that a pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations. This conclusion is confirmed with sensitivities presented in Table 6. Results for recovery of the coefficient vector are presented in Table 5.
Table 6: Sensitivities for Example 5.2.
$j$     poles $z_{j}^{(-)}$      $\eta_{j}^{(-)}$      $\rho_{j}^{(-)}$      bound $\rho_{j}^{(-)}$
1 $0.9$ 0.243 7.748 13.548
2 $-0.9$ 0.121 3.874 6.774
3 $0.9\mathrm{i}$ 0.081 2.582 4.516
4 $-0.9\mathrm{i}$ 0.06 1.937 3.387
$j$     poles $z_{j}^{(+)}$      $\eta_{j}^{(+)}$      $\rho_{j}^{(+)}$      bound $\rho_{j}^{(+)}$
1 $1.1$ 0.056 3.533 5.695
2 $-1.1$ 0.047 2.944 4.745
3 $1.1\mathrm{i}$ 0.04 2.523 4.067
4 $-1.1\mathrm{i}$ 0.035 2.208 3.559
Example 5.3.
Finally, let a rational function $r(z)$ of type $(3,4)$ in (1.1), i.e with $M=4$, be given by vectors of poles and coefficients
$\boldsymbol{z}=(0.2,0.5,2,50),\quad\boldsymbol{\gamma}=(1,1,1,1),$
respectively. For $N=4$ and $\sigma=0$, we get the following reconstruction errors
$e(\boldsymbol{z})=1.42\cdot 10^{-13},\quad e(\boldsymbol{\gamma})=9.27\cdot 10^{-15}.$
Table 7: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $\sigma=10^{-7}$
$N=4$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min $1.58$e–$08$ $2.9$e–$08$ $6.07$e–$09$ $1.75$e–$03$
max $5.1$e–$07$ $2.49$e–$07$ $2.7$e–$07$ $3.21$e–$01$
average $2.48$e–$07$ $1.3$e–$07$ $1.45$e–$07$ $1.34$e–$01$
Table 8: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $\sigma=10^{-5}$
$N=4$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min $4.24$e–$06$ $1.18$e–$06$ $2.57$e–$06$ $3.16$e–$00$
max $6.45$e–$05$ $2.77$e–$05$ $2.92$e–$05$ $9.46$e+$01$
average $2.31$e–$05$ $1.13$e–$05$ $1.5$e–$05$ $2.02$e+$01$
Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$. We compute 10 iterations and present the corresponding results in Tables 7 and 8. Analyzing numerical results, we can conclude that the pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations. For $\sigma=10^{-5}$ the method does not reconstruct location of this pole. This numerical observation is also predicted by the sensitivities values presented in Table 9. Both unstructured $\rho_{2}^{(+)}=2.204\cdot 10^{3}$ and structured $\eta_{2}^{(+)}=1.577\cdot 10^{4}$ sensitivities are sufficiently larger than sensitivities of other poles. Note, that locations of poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed even for $\sigma=10^{-1}$. We present results regarding recovery of the coefficient vector in Table 10.
Table 9: Sensitivities for Example 5.3.
$j$     poles $z_{j}^{(-)}$      $\eta_{j}^{(-)}$      $\rho_{j}^{(-)}$      bound $\rho_{j}^{(-)}$
1 $0.2$ $1.807\cdot 10^{1}$ $1.766\cdot 10^{1}$ $3.937\cdot 10^{1}$
2 $0.5$ $1.518\cdot 10^{1}$ $2.249\cdot 10^{1}$ $5.625\cdot 10^{1}$
$j$     poles $z_{j}^{(+)}$      $\eta_{j}^{(+)}$      $\rho_{j}^{(+)}$      bound $\rho_{j}^{(+)}$
1 $2$ $1.973\cdot 10^{1}$ 5.43 $1.993\cdot 10^{1}$
2 $50$ $1.577\cdot 10^{4}$ $2.204\cdot 10^{3}$ $6.477\cdot 10^{3}$
We compute also vectors of structured perturbations for poles inside the unit circle
$(\boldsymbol{\mathcal{S}}^{(1)}_{4})^{(-)}=(0.555,4.999,13.333,11.111)^{T},\quad(\boldsymbol{\mathcal{S}}^{(2)}_{4})^{(-)}=(0.222,2.666,9.999,11.111)^{T},$
and outside the unit circle
$(\boldsymbol{\mathcal{S}}^{(1)}_{4})^{(+)}=(3.472\cdot 10^{-3},0.354,9.375,17.36)^{T},$
and
$(\boldsymbol{\mathcal{S}}^{(2)}_{4})^{(+)}=(54.25,2.929\cdot 10^{3},1.106\cdot 10^{4},1.085\cdot 10^{4})^{T}.$
[…]
$N=4$ $\sigma=10^{-7}$ $\sigma=10^{-5}$
min $7.38$e–$04$ $3.79$e–$02$
max $7.38$e–$03$ $6.95$e–$01$
average $4$e–$03$ $4.11$e–$01$
That means that poles $z_{1}^{(-)}=0.2$ and $z_{2}^{(-)}=0.5$ are the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-3}$, the pole $z_{1}^{(+)}=2$ – the Fourier coefficient $\hat{r}_{4}$, and the pole $z_{2}^{(+)}=50$ – the Fourier coefficient $\hat{r}_{3}$.
Table 11: $\ell_{2}$-norms and upper bounds of $\ell_{2}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3
$\|\boldsymbol{\rho}^{(-)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{2}$) $\|\boldsymbol{\rho}^{(+)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{2}$)
$N=4$ $2.86$e+$01$ $6.866$e+$01$ $2.204$e+$03$ $6.477$e+$03$
Table 12: $\ell_{1}$-norms and upper bounds of $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3
$\|\boldsymbol{\rho}^{(-)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{1}$) $\|\boldsymbol{\rho}^{(+)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{1}$)
$N=4$ $4.01$e+$01$ $5.725$e+$01$ $2.209$e+$03$ $6.802$e+$03$
Further we compute the $\ell_{1}$- and $\ell_{2}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bounds obtained in (4.21), (4.22) and (4.23). We present the results in Tables 11 and 12. Analyzing these results, we conclude that for this example upper bounds (4.21), (4.22) and (4.23) are quite sharp.
Acknowledgement
The author gratefully acknowledge support by the German Research Foundation in the framework of the RTG 2088. The author would like to thank Ion Victor Gosea (MPI Magdeburg) for several valuable discussions regarding sensitivities of the eigenvalues of a matrix pencil problem.

### v2

@@
 where $\gamma_{j},\,z_{j}\in\mathbb{C}$, $\gamma_{j}\neq 0$ and poles $z_{j}$ are pairwise distinct.
-We study reconstruction of rational functions (1.1) using as input data a finite set of their Fourier coefficients.
+We study the question of recovery of rational functions (1.1) using as input data a finite set of their Fourier coefficients $\widehat{r}_{k}$ defined by (4.1).
 We use the property that Fourier coefficients of rational functions have the exponential structure and we reconstruct this exponential structure in the frequency domain applying the ESPRIT method [19, 17].
@@
 Similar technique for the recovery of or approximation by rational functions was employed in [7, 20, 21, 22].
-We present modification of these ideas, by applying the ESPRIT method for the exponential recovery.
-Since ESPRIT is based on using the Hankel pencil, it makes the method suitable for sensitivity analysis.
-The other methods for the recovery of rational functions (in time domain) such as the AAA method [13], the RKFIT (Rational Krylov Fitting) method [3], the Vector Fitting algorithm [12] and the Loewner framework method [1], were considered in [11].
-Note that, in [6, 18, 19, 20], similar to [11], reconstruction from given noisy data is limited to numerical observations.
-In [21], for the Loewner framework method the authors study sensitivities of Loewner pencil eigenvalues (which are poles of a transfer function defined by the corresponding Loewner matrices) with respect to unstructured and structured perturbations.
-Following approach in [21], we present the sensitivity analysis for the poles computed by our method.
-According to our method, poles located inside and outside the unit circle are reconstructed separately as eigenvalues of some Hankel matrix pencils.
-Further we consider two kinds of Hankel pencil perturbations.
-For the sensitivity $\rho$ with respect to unstructured perturbations, we present the connection with condition numbers of Vandermonde matrices that are used in factorizations of the corresponding Hankel matrices.
+The paper [21] is of special interest to us.
+In [21], the author considered recovery of the rational function (1.1) based on using noisy input data and discussed reconstruction results for the poles $z_{j}$, $j=1,...,M$.
+It was pointed out that for different levels of noise, not all poles $z_{j}$ are reconstructed with the same accuracy, and sometimes even the locations of certain poles are not reconstructed correctly.
+From these numerical observations, we can conclude that some poles $z_{j}$ of the rational function (1.1) are more ”sensitive“ to perturbations in the input data than the others.
+Our goal in this paper is to present theoretical explanations of these numerical observations by applying the sensitivity analysis.
+The method developed in [21] employees the AAK theory [16] for the reconstruction of the exponential structure of the Fourier coefficients $\widehat{r}_{k}$.
+We modify these ideas, by applying the ESPRIT method in the frequency domain.
+Since ESPRIT is based on using the Hankel pencil, it makes our method suitable for the sensitivity analysis.
+In [23], the authors developed the sensitivity analysis for the eigenvalues of Loewner pencil with respect to the unstructured and structured perturbations.
+We referee the interested reader to [1] for further application of Loewner framework for model reduction of dynamical systems.
+Following approach in [23], we present the sensitivity analysis for the poles of $r$ in (1.1) computed by our method.
+According to our idea, poles located inside and outside the unit circle are reconstructed separately as eigenvalues of some Hankel matrix pencils.
+Further, we consider two kinds of Hankel pencil perturbations.
+For the sensitivity $\rho$ with respect to unstructured perturbations, we present the connection with the condition numbers of Vandermonde matrices that are used in factorizations of the corresponding Hankel matrices.
 The sensitivity $\eta$ with respect to structured perturbations of Hankel pencil is studied in [24].
-We present modification of these results for our special reconstruction problem.
-Note that the sensitivity of eigenvalues of Hankel pencil perturbations was studied also in [10], [2].
-In [10] and [2], the authors measure both structured and unstructured perturbations by the same quantity.
-Following approach in [21], we consider two different quantities for structured and unstructured sensitivities respectively.
-Further using sensitivity analysis, we explain reconstruction errors for poles from noisy Fourier coefficients.
-We refer also to the paper [4] for further application of the sensitivity analysis of Hankel pencil.
-The method offered in this paper is also very much related to our ESPIRA method for the exponential recovery [7][9].
-In ESPIRA, we use the property that the classical Fourier coefficients of exponential sums have a rational structure and to reconstruct this rational structure we apply the AAA method for rational approximation [13] in the frequency domain.
-Thus the idea of the method for rational recovery used in this paper can be considered as converse to the idea of our ESPIRA method.
+We present modification of these results for our special rational reconstruction problem (1.1).
+Note that the theory of sensitivities of eigenvalues of Hankel pencil perturbations was considered also in [11, 3].
+In [11] and [3], the authors measured both structured and unstructured perturbations by the same quantity.
+Similarly as in [23], we introduce two different quantities for structured $\eta$ and unstructured $\rho$ sensitivities, respectively, of poles of the rational function (1.1).
+By checking if $\rho$ or $\eta$ is ”large“ for some pole $z_{j}$, we measure the ill-disposedness of $z_{j}$, while consideration of $\eta$ gives us also the additional information regarding that to perturbation of which measurement $\widehat{r}_{k}$ the pole $z_{j}$ is the most sensitive.
+The method developed in this paper is also closely related to our ESPIRA method for the exponential recovery [8][10].
+ESPIRA uses the property that the classical Fourier coefficients of exponential sums have a rational structure and reconstructs this rational structure with the AAA method for rational approximation [14] in the frequency domain.
+Thus, the idea of rational reconstruction used in this paper can be considered as converse to the idea of ESPIRA.
+We want to pay a special attention also to the paper [5], where authors developed a method for exponential reconstruction using the sensitivity analysis of Hankel pencil eigenvalues.
+Finally, we want to mention the other methods for the recovery of rational functions such as the AAA method [14], the RKFIT (Rational Krylov Fitting) method [4], the Vector Fitting algorithm [13] and the Loewner framework method [1].
+In [12], the authors discussed all these algorithms for the recovery of rational functions using as input information given noisy set of samples of $r$.
+Since we are interested in the sensitivity analysis of the poles, we have to assume that we apply perturbations to the Fourier coefficients $\widehat{r}_{k}$.
+Note that we can introduce the corresponding modification of the method based on employing function value of $r$ by computing $\widehat{r}_{k}$ via the quadrature formula for some grid of points (see Remark 4.1).
+Modifying the algorithm in this way will allow us to compare its performance with known rational methods.
+Numerical experiments show us that our approach gives equally good and for some examples even better recovery results from given noisy function values than the AAA method.
 Outline.
-In Section 2, we present the main ideas of the ESPRIT method for exponential recovery.
-In Section 3 we consider sensitivity analysis for Hankel pencil eigenvalues.
-Subsection 3.1 is dedicated to unstructured perturbations and in Subsection 3.2, we consider structured perturbations.
+In Section 2, we present the main ideas of the ESPRIT method for the exponential recovery.
+In Section 3, we consider the sensitivity analysis for Hankel pencil eigenvalues: Subsection 3.1 is dedicated to unstructured perturbations and in Subsection 3.2, we consider structured perturbations.
 In Section 4, we describe our main results.
-In Subsection 4.1, we present main ideas of our method for the recovery of rational functions (1.1).
+In Subsection 4.1, we present main ideas of our new approach for the recovery of rational functions (1.1).
 Subsection 4.2 is dedicated to the sensitivity analysis of poles reconstructed with our method.
@@
 For square matrices, we often use the short notation $\mathbf{X}_{K}$ instead of $\mathbf{X}_{K,K}$.
-By $\left(\mathbf{X}_{L,K}\right)^{\dagger}$ we denote the Moore-Penrose inverse of $\mathbf{X}_{L,K}$.
+By $\mathbf{X}_{L,K}^{\dagger}$ we denote the Moore-Penrose inverse of $\mathbf{X}_{L,K}$ and by $\mathbf{X}_{K}^{-1}$ inverse of the invertible matrix $\mathbf{X}_{K}$.
 We consider $\|\mathbf{X}_{L,K}\|_{2}:=\sigma_{\max}(\mathbf{X}_{L,K})$ the 2-norm or spectral norm, and $\|\mathbf{X}_{L,K}\|_{\mathrm{F}}:=\left(\sum\limits_{i=1}^{L}\sum\limits_{j=1}^{K}|X_{ij}|^{2}\right)^{1/2}=\left(\sum\limits_{i=1}^{\min\{L,K\}}\sigma_{i}^{2}\right)^{1/2}$ the Frobenius norm of the matrix $\mathbf{X}_{L,K}$.
@@
 3 Sensitivities of Hankel pencil eigenvalues
-In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3).
-The concept of sensitivity arises from the perturbation of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [10] and further considered in [2].
+The concept of sensitivities arises from the perturbation of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [11] and further considered in [3].
 In [23], the notion of sensitivities was defined for Loewner matrix pencil.
@@
 Using similar ideas, in [24], the author considered the structured sensitivity of Hankel matrix pencil.
+In [5], the sensitivity analysis of Hankel pencil was applied to develop a new method for exponential analysis for the recovery from noisy data.
+In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3) and further in Sunbsection 4.2 we present modification of the results from this section to develop sensitivity analysis for the poles of rational functions.
 3.1 Sensitivity of Hankel pencil eigenvalues with unstructured perturbations
@@
 Formula (3.2) is obtained from the perturbed system
-$(\boldsymbol{p}^{T}+(\boldsymbol{p}^{(1)})^{T}+\ldots)$ $(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)})=(z+z^{(1)}+\ldots)(\boldsymbol{p}^{T}+(\boldsymbol{p}^{(1)})^{T}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{{\mathbf{H}_{M}(0)}}),$
-$(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)}))$ $(\boldsymbol{q}+\boldsymbol{q}^{(1)}+\ldots)=(z+z^{(1)}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)}))(\boldsymbol{q}+\boldsymbol{q}^{(1)}+\ldots),$
+$(\boldsymbol{p}^{T}+(\tilde{\boldsymbol{p}}^{(1)})^{T}+\ldots)$ $(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)})=(z+z^{(1)}+\ldots)(\boldsymbol{p}^{T}+(\tilde{\boldsymbol{p}}^{(1)})^{T}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{{\mathbf{H}_{M}(0)}}),$
+$(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)}))$ $(\boldsymbol{q}+\tilde{\boldsymbol{q}}^{(1)}+\ldots)=(z+z^{(1)}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)}))(\boldsymbol{q}+\tilde{\boldsymbol{q}}^{(1)}+\ldots),$
 by retaining only first-order terms (see [11], [3], [23] and [24] for details).
@@
 where $\varepsilon>0$. Then from (3.2) taking into account (3.3), we get for the GEP (3.1)
-$|z^{(1)}|$ $=\left|\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}_{L}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}\right|$
+$|z^{(1)}|$ $=\left|\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}\right|$
 $\leq\frac{\|\boldsymbol{p}^{T}\|_{2}\,\|\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)}\|_{2}\,\|\boldsymbol{q}\|_{2}}{|\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}|}\leq\epsilon\rho,$
@@
 Since Hankel matrices $\mathbf{H}_{M}(0)$ and $\mathbf{H}_{M}(1)$ satisfy factorizations (2.4) and (2.5) for $N=L=M$, left $\boldsymbol{p}^{(j)}$ and right $\boldsymbol{q}^{(j)}$ eigenvectors of the GEP (3.1) corresponding to the eigenvalue $z_{j}$ are given by formulas
-$\boldsymbol{q}^{(j)}=\boldsymbol{p}^{(j)}=\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j},$ (3.5)
+$\boldsymbol{q}^{(j)}=\boldsymbol{p}^{(j)}=\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j},$ (3.5)
 where $\boldsymbol{e}_{j}$ is the $j$th canonical vector in $\mathbb{R}^{M}$. Then using again the factorization (2.4) of the Hankel matrix $\mathbf{H}_{M}(0)$, we get
-$(\boldsymbol{p}^{(j)})^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}^{(j)}$ $=\left(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\right)^{T}\,\mathbf{H}_{M}(0)\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}$
+$(\boldsymbol{p}^{(j)})^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}^{(j)}$ $=\left(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\right)^{T}\,\mathbf{H}_{M}(0)\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}$
 $=\boldsymbol{e}_{j}^{T}\,\mathbf{\Gamma}_{M}\,\boldsymbol{e}_{j}=\gamma_{j}.$ (3.6)
@@
 $\rho_{j}=\frac{1}{|\gamma_{j}|}\|(\boldsymbol{p}^{(j)})^{T}\|_{2}\,(|z_{j}|\,\|\mathbf{H}_{M}(0)\|_{2}+\|\mathbf{H}_{M}(1)\|_{2})\,\|\boldsymbol{q}^{(j)}\|_{2}.$ (3.7)
-The ill-disposedness of the eigenvalue $z_{j}$ with respect to unstructured perturbations can be measured by checking whether $\rho_{j}$ is "large".
+The ill-disposedness of the eigenvalue $z_{j}$ with respect to unstructured perturbations can be measured by checking whether $\rho_{j}$ is ”large“.
 Applying formulas (2.4), (2.5) and (3.5) to (3.7), we obtain for $j=1,\ldots,M$
-$\rho_{j}$ $=\frac{1}{|\gamma_{j}|}\left\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\right\|_{2}^{2}\left(|z_{j}|\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}+\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}\right).$ (3.8)
+$\rho_{j}$ $=\frac{1}{|\gamma_{j}|}\left\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\right\|_{2}^{2}\left(|z_{j}|\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}+\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}\right).$ (3.8)
 The formula (3.8) is used to compute the exact values of sensitivities $\rho_{j}$.
@@
 $\rho_{j}\leq\frac{|z_{j}|\,\|\mathbf{\Gamma}_{M}\|_{2}+\|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}}{|\gamma_{j}|}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right),$
-where $\kappa(\mathbf{X}):=\|\mathbf{X}\|_{2}\,\|\mathbf{X}^{\dagger}\|_{2}$ denotes the condition number of the matrix $\mathbf{X}$.
+where $\kappa(\mathbf{X}):=\|\mathbf{X}\|_{2}\,\|\mathbf{X}^{-1}\|_{2}$ denotes the condition number of the matrix $\mathbf{X}$.
 For diagonal matrices $\mathbf{\Gamma}_{M}$ and $\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}$, it holds
@@
 $\zeta_{j}:=\frac{1}{|\gamma_{j}|}\left(|z_{j}|\max\limits_{k=1,\ldots,M}|\gamma_{k}|+\max\limits_{k=1,\ldots,M}|\gamma_{k}z_{k}|\right).$ (3.11)
+Note that we can proceed further with the estimation of (3.10) applying known results for upper bounds of condition numbers of the Vandermonde matrices [15, 2].
 Define further a vector of unstructured sensitivities $\boldsymbol{\rho}:=(\rho_{1},\ldots,\rho_{M})^{T}\in\mathbb{R}^{M}$ and let $\boldsymbol{\zeta}:=(\zeta_{1},\ldots,\zeta_{M})^{T}\in\mathbb{R}^{M}$.
@@
 From (3.8), taking into account (3.9) and (3.11), we obtain
-$\rho_{j}\leq\zeta_{j}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}\,\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}.$
+$\rho_{j}\leq\zeta_{j}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}\,\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}.$
 By summing up the last inequalities over $j=1,\ldots,M$, we get
-$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}$
+$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}$
 where $\zeta_{\max}:=\max_{j=1,\ldots,M}\zeta_{j}$. Finally using the property $\sum_{j=1}^{M}\|\mathbf{X}\boldsymbol{e}_{j}\|_{2}^{2}=\|\mathbf{X}\|_{\mathrm{F}}^{2}$, we obtain
-$\|\boldsymbol{\rho}\|_{1}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\big{\|}\mathbf{V}^{\dagger}_{M}(\boldsymbol{z})\big{\|}_{\mathrm{F}}^{2}$
+$\|\boldsymbol{\rho}\|_{1}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\big{\|}\mathbf{V}^{-1}_{M}(\boldsymbol{z})\big{\|}_{\mathrm{F}}^{2}$
 for the $\ell_{1}$-norm of the sensitivity vector $\boldsymbol{\rho}$.
@@
 Using (3.5), we rewrite components $S^{(j)}_{m}$ in the following way
-$S^{(j)}_{m}=$ $\frac{1}{\gamma_{j}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{m-k,j}$
-$-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{m-k+1,j}\bigg{)}.$
+$S^{(j)}_{m}=$ $\frac{1}{\gamma_{j}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{k,j}\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{m-k,j}$
+$-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{m-k+1,j}\bigg{)}.$
 Formula (3.13) can be also written in a matrix form
@@
 4 Recovery of rational functions via Hankel matrix pencil method and sensitivities of the poles
-In this section, we describe the main ideas of our new method for the rational reconstruction and present analysis of the sensitivities of the poles.
+In this section, we describe the main ideas of our new method for the recovery of a rational function of the form (1.1) and present the sensitivity analysis for its poles.
 4.1 Description of the method
@@
 which is closely related to the classical Fourier coefficient of a function $r(\mathrm{e}^{\mathrm{i}\varphi})$,
-$\widehat{r}_{k}:=\frac{1}{2\pi}\int\limits_{0}^{2\pi}r(\mathrm{e}^{\mathrm{i}\varphi})\mathrm{e}^{-\mathrm{i}k\varphi}\,\mathrm{d}\varphi=\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t}.$
+$\widehat{r}_{k}:=\frac{1}{2\pi}\int\limits_{0}^{2\pi}r(\mathrm{e}^{\mathrm{i}\varphi})\mathrm{e}^{-\mathrm{i}k\varphi}\,\mathrm{d}\varphi=\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t}.$ (4.1)
 Then (see, for example, [21])
@@
 Remark 4.1.
-Note that coefficients $\widehat{r}_{k}$ can be also computed applying the quadrature formulas on some grid of points.
-In that case, we will use the function values instead of Fourier coefficients as input data for the algorithm.
-This question was discussed in details in [19], where the author consider the random access model and the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$.
-In this paper, we are interested in the presentation of a modification of the method from [19] and in the explanation of the results of this method regarding recovery of noisy data with the use of sensitivities.
+1.
+Note that coefficients $\hat{r}_{k}$ can be also computed applying the quadrature formulas on some grid of points.
+In that case, we will use the function values of $r$ instead of its Fourier coefficients as input data for the algorithm.
+Let us chose points $t_{n}=\mathrm{e}^{\frac{2\pi\mathrm{i}n}{4N}}$, $n=0,...,4N-1$ and let $(R_{k})_{k=0}^{4N-1}$ be the DFT of $(r(t_{n}))_{n=0}^{4N-1}$.
+Then $\hat{r}_{k}$ in (4.1) can be computed as $\hat{r}_{k}=R_{k+2N}$, $k=-2N,...,2N-1$.
+This question was discussed in details in [21], where the author consider not only the points on the unit circle, but also the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$.
+On the one hand, we are interested in the presentation of a modification of the method from [21] and in the explanation of the results of this method for the recovery from noisy data with the use of the sensitivities.
+In that case, we have to employ the given noisy coefficients $\widehat{r}_{k}$ as input data.
+On the other hand, we also want to compare the performance of known methods for rational approximation with our new approach based on Hankel pencil.
+To guarantee fair comparison, we assume to have given noisy values $(r(t_{n}))_{n=0}^{4N-1}$.
+2.
+The method based on using the given coefficients $\hat{r}_{k}$, $k=-2N,...,2N$ has the overall computational cost of $\mathcal{O}(NL^{2})$, which is determined by computation of SVD of Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ and $\mathbf{H}^{(+)}_{2N-L,L+1}$.
+The complexity of the modified algorithm based on using the values $(r(t_{n}))_{n=0}^{4N-1}$ remains the same, since the additional operation computation of FFT of $(r(t_{n}))_{n=0}^{4N-1}$ requires only $\mathcal{O}(N\log N)$ flops.
 4.2 Sensitivities of poles
@@
 Since Hankel matrices in the matrix pencil (4.12) have similar structure as Hankel matrices in the matrix pencil (2.6), we apply ideas from Subsection 3.1 without any changes and obtain for the sensitivities $\rho_{n}^{(-)}$, $n=1,\ldots,M_{1}$ of the poles $z^{(-)}_{j_{1}},\ldots,z^{(-)}_{j_{M_{1}}}$ in the $\mathbb{D}$ the formula for the exact computation
-$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
+$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
 ×$\times\left(|z_{j_{n}}^{(-)}|\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}+\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{Z}_{M_{1}}\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}\right),$ (4.14)
@@
 Thus left and right eigenvectors of the GEP of the Hankel matrix pencil (4.13) corresponding to the eigenvalue $\frac{1}{z^{(+)}_{j_{n}}}$ are given by formulas
-$\boldsymbol{q}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n},\quad\boldsymbol{p}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n},$ (4.17)
+$\boldsymbol{q}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n},\quad\boldsymbol{p}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n},$ (4.18)
 where $\boldsymbol{e}_{n}$ is the $n$th canonical (column) vector in $\mathbb{R}^{M_{2}}$ Taking into account (4.16) and (4.18), we get
-$(\boldsymbol{p}^{(n)})^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\boldsymbol{q}^{(n)}$ $=\left(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}\right)^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}$
+$(\boldsymbol{p}^{(n)})^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\boldsymbol{q}^{(n)}$ $=\left(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}\right)^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}$
 $=-\boldsymbol{e}_{n}^{T}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\boldsymbol{e}_{n}=-\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}.$ (4.19)
 From (3.4) applying (4.16)–(4.19), we obtain the formula for the exact computation of sensitivities $\rho_{n}^{(+)}$ of the poles $z^{(+)}_{j_{n}}$ (the eigenvalues $\frac{1}{z^{(+)}_{j_{n}}}$ of the matrix pencil (4.13)) located outside the unit circle
-$\rho_{n}^{(+)}=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}|}\bigg{\|}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}\bigg{\|}_{2}^{2}$
+$\rho_{n}^{(+)}=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}|}\bigg{\|}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}\bigg{\|}_{2}^{2}$
 ×$\times\bigg{(}\frac{1}{|z_{j_{n}}^{(+)}|}\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{2}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}+\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}\bigg{)},$ (4.20)
@@
 Using ideas similar as in Subsection 3.1, we write also the upper bounds for the $\ell_{1}$-norm of the sensitivity vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
-$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{\dagger}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2}$ (4.22)
-$\|\boldsymbol{\rho}^{(+)}\|_{1}$ $\leq\zeta^{(+)}_{\max}\,\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{2}^{2}\,\|\mathbf{V}^{\dagger}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{\mathrm{F}}^{2},$ (4.23)
+$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{-1}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2}$ (4.23)
+$\|\boldsymbol{\rho}^{(+)}\|_{1}$ $\leq\zeta^{(+)}_{\max}\,\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{2}^{2}\,\|\mathbf{V}^{-1}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{\mathrm{F}}^{2},$ (4.24)
 where $\zeta_{\max}^{(-)}:=\max_{j=1,\ldots,M_{1}}\zeta_{j}^{(-)}$ and $\zeta_{\max}^{(+)}:=\max_{j=1,\ldots,M_{2}}\zeta_{j}^{(+)}$.
+Naturally, arises the question how sharp are the obtained inequalities (4.224.24).
+In Example 5.3, we present the numerical computations for the exact values of the $\ell_{2}$- and $\ell_{1}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bound.
+Based on the results presented in Table 10, we can conclude that estimates (4.224.24) are quite sharp.
 Let us now consider the case of structured perturbations. Define further matrices
@@
 by their components
-$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}:=$ $\frac{1}{\gamma_{j_{n}}^{(-)}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{m-k,n}$
-$-z_{j_{n}}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{m-k+1,n}\bigg{)}$ (4.24)
+$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}:=$ $\frac{1}{\gamma_{j_{n}}^{(-)}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{m-k,n}$
+$-z_{j_{n}}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{m-k+1,n}\bigg{)}$ (4.25)
 and
-$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}:=$ $-\frac{1}{\gamma_{j_{n}}^{(+)}\,z_{j_{n}}^{-2}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{m-k,n}$
-$-\frac{1}{z_{j_{n}}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{m-k+1,n}\bigg{)}.$ (4.25)
+$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}:=$ $-\frac{1}{\gamma_{j_{n}}^{(+)}\,z_{j_{n}}^{-2}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{m-k,n}$
+$-\frac{1}{z_{j_{n}}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{m-k+1,n}\bigg{)}.$ (4.26)
 Formulas (4.25) and (4.26) are obtained using similar ideas from Subsection 3.2. In (4.26), we take into account (4.19).
@@
 are structured sensitivities for the poles $z_{j_{n}}^{(-)}$ and $z_{j_{n}}^{(+)}$ respectively.
-5 Numerical experiments
-In this section, we present a few numerical experiments.
-The algorithm is implemented in Matlab and use IEEE standard floating point arithmetic with double precision.
-We use the notations $\tilde{z}_{j}$ and $\tilde{\gamma}_{j}$ for the poles and the coefficients, respectively, reconstructed by the method considered in Subsection 4.1.
-We define the errors for the poles $z_{j}$ and the coefficients $\gamma_{j}$ by the following formulas
-[…]
-Note that in the case when we deal with noisy data, we assume that a number of poles $M$, as well as numbers of poles located inside and outside the unit circle $M_{1}$ and $M_{2}$ respectively are known.
 Table 1: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.1
@@
 average $1.09$e–$02$ $1.57$e–$01$ $1.36$e–$02$ $1.05$e–$01$
-We employ $2N$ (noisy) Fourier coefficients $\hat{r}_{-k}(1+n_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+n_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(+)}$.
-By $(n_{k})_{k=1}^{2N-1}$ we denote the random noise that is generated in Matlab by $\sigma$*randn(2N), i.e., we take multiplicative noise drawn from the standard normal distribution.
+5 Numerical experiments
+In this section, we present a few numerical experiments.
+The algorithms are implemented in Matlab and use IEEE standard floating point arithmetic with double precision.
+By $\tilde{z}_{j}$ and $\tilde{\gamma}_{j}$ we denote reconstructed poles and coefficients, respectively, and by the following formulas
+[…]
+we define the corresponding reconstruction errors.
+Note that in the case when we deal with noisy data, we assume that a number of poles $M$, as well as numbers of poles located inside and outside the unit circle $M_{1}$ and $M_{2}$, respectively, are known.
+In Examples 5.1-5.3, we demonstrate recovery results and analyze sensitivities of poles with the method developed in Section 4.
+We employ $2N$ (noisy) Fourier coefficients $\hat{r}_{-k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(+)}$.
+By $(\varepsilon_{k})_{k=1}^{2N}$ we denote the random noise that is generated in Matlab by $\sigma$*randn(2N), i.e., we take multiplicative noise drawn from the standard normal distribution.
 We take $L=N$ and consider different values of $N$ and $\sigma$.
-To compute sensitivities $\rho_{j}^{(-)}$, $\rho_{j}^{(+)}$ and ,$\eta_{j}^{(-)}$, $\eta_{j}^{(+)}$, we use formulas (4.13), (4.19) and (4.26), (4.27) respectively.
-Upper bounds for sensitivities $\rho_{j}^{(-)}$ and $\rho_{j}^{(+)}$ are computed via (4.14) and (4.20) respectively.
+To compute sensitivities $\rho_{j}^{(-)}$, $\rho_{j}^{(+)}$ and ,$\eta_{j}^{(-)}$, $\eta_{j}^{(+)}$, we use formulas (4.14), (4.20) and (4.27), (4.28), respectively.
+Upper bounds for sensitivities $\rho_{j}^{(-)}$ and $\rho_{j}^{(+)}$ are computed via (4.15) and (4.21), respectively.
 […]
@@
 average $1.44$e–$03$ $8.48$e–$04$ average $1.11$e–$01$ $9.55$e–$02$
-Example 5.1.
-Let a rational function $r(z)$ in (1.1) be given by parameters $M=2$ and
-$\boldsymbol{z}=(-0.1,-2.1),\quad\boldsymbol{\gamma}=(0.5,0.5)$
-(see [5, 21]). When $\sigma=0$, i.e. in case of recovery from exact data, we get the following reconstruction errors with $N=2$,
-[…]
-For $\sigma=10^{-3}$ and $\sigma=10^{-1}$, we compute 10 iterations and present the results in Table 1.
-Since we are interested to investigate which pole is more sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ separately.
-Results for recovery of the coefficient vector are presented in Table 2.
 Table 3: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.2 for $\sigma=10^{-3}$
@@
 average $1.44$e–$04$ $1.19$e–$04$ $1.04$e–$04$ $9.43$e–$05$
-Analyzing numerical results in Table 1, we conclude that for $\sigma=0.1$ our method still reconstructs location of poles quite good.
-On the other hand, using more (noisy) Fourier coefficients also gives us better recovery results.
-We notice that a pole $z_{1}^{(-)}=-0.1$ is less sensitive to perturbations than a pole $z_{1}^{(+)}=-2.1$.
-This numerical observation is also confirmed with sensitivities values.
-For a pole $z_{1}^{(-)}=-0.1$, we have $\rho_{1}^{(-)}=2\cdot 10^{-1}$ and $\eta_{1}^{(-)}=1.97$, and for a pole $z_{1}^{(+)}=-2.1$, we have $\rho_{1}^{(+)}=9.52\cdot 10^{-1}$ and $\eta_{1}^{(+)}=6.65$.
-Both unstructured and structured sensitivities are larger for a pole $z_{1}^{(+)}=-2.1$ located outside the unit circle.
-We compute also vectors of structured perturbations
-$(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T}.$
-That means that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ with respect to perturbation of the Fourier coefficient $\hat{r}_{2}$.
 Table 4: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.2 for $\sigma=10^{-1}$
@@
 average $1.56$e–$02$ $1.4$e–$02$ $1.24$e–$02$ $1.15$e–$02$
-Example 5.2.
-We consider an example when poles are located very close to the unit circle. Let a rational function $r(z)$ of type $(7,8)$ in (1.1), i.e with $M=8$, be given by vectors of poles and coefficients
-$\boldsymbol{z}=(0.9,-0.9,0.9\mathrm{i},-0.9\mathrm{i},1.1,-1.1,1.1\mathrm{i},-1.1\mathrm{i}),\quad\boldsymbol{\gamma}=(1,2,3,4,5,6,7,8),$
-respectively. We consider first recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we get the following reconstruction errors
-$e(\boldsymbol{z})=1.45\cdot 10^{-15},\quad e(\boldsymbol{\gamma})=5.48\cdot 10^{-14}.$
 […]
@@
 average $9.7$e–$03$ $6.35$e–$01$
-In case of recovery from noisy data, we compute 10 iterations and present recovery result for each pole $z_{j}$ in Tables 3 and 4.
-Analyzing numerical results, we conclude that a pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations.
-This conclusion is confirmed with sensitivities presented in Table 6.
-Results for recovery of the coefficient vector are presented in Table 5.
 Table 6: Sensitivities for Example 5.2.
@@
 Example 5.1.
-Finally, let a rational function $r(z)$ of type $(3,4)$ in (1.1), i.e with $M=4$, be given by vectors of poles and coefficients
-$\boldsymbol{z}=(0.2,0.5,2,50),\quad\boldsymbol{\gamma}=(1,1,1,1),$
-respectively. For $N=4$ and $\sigma=0$, we get the following reconstruction errors
-$e(\boldsymbol{z})=1.42\cdot 10^{-13},\quad e(\boldsymbol{\gamma})=9.27\cdot 10^{-15}.$
-Table 7: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $\sigma=10^{-7}$
-$N=4$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
+Let a rational function $r(z)$ in (1.1) be given by parameters $M=2$ and
+$\boldsymbol{z}=(-0.1,-2.1),\quad\boldsymbol{\gamma}=(0.5,0.5)$
+(see [6, 23]). When $\sigma=0$, i.e. in case of recovery from exact data, we get the following reconstruction errors with $N=2$,
+[…]
+For $\sigma=10^{-3}$ and $\sigma=10^{-1}$, we compute 10 iterations and present the results in Table 1.
+Since we are interested to investigate which pole is more sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ separately.
+Results for the recovery of the coefficient vector are presented in Table 2.
+Table 7: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $N=4$
+$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
 min $1.58$e–$08$ $2.9$e–$08$ $6.07$e–$09$ $1.75$e–$03$
@@
 average $2.48$e–$07$ $1.3$e–$07$ $1.45$e–$07$ $1.34$e–$01$
-Table 8: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $\sigma=10^{-5}$
-$N=4$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
+$\sigma=10^{-5}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
 min $4.24$e–$06$ $1.18$e–$06$ $2.57$e–$06$ $3.16$e–$00$
@@
 average $2.31$e–$05$ $1.13$e–$05$ $1.5$e–$05$ $2.02$e+$01$
-Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$.
-We compute 10 iterations and present the corresponding results in Tables 7 and 8.
-Analyzing numerical results, we can conclude that the pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations.
-For $\sigma=10^{-5}$ the method does not reconstruct location of this pole.
-This numerical observation is also predicted by the sensitivities values presented in Table 9.
-Both unstructured $\rho_{2}^{(+)}=2.204\cdot 10^{3}$ and structured $\eta_{2}^{(+)}=1.577\cdot 10^{4}$ sensitivities are sufficiently larger than sensitivities of other poles.
-Note, that locations of poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed even for $\sigma=10^{-1}$.
-We present results regarding recovery of the coefficient vector in Table 10.
+Analyzing numerical results in Table 1, we conclude that for $\sigma=0.1$ our method still reconstructs the locations of the poles quite good.
+On the other hand, using more (noisy) Fourier coefficients also gives us better recovery results.
+We notice that a pole $z_{1}^{(-)}=-0.1$ is less sensitive to perturbations than a pole $z_{1}^{(+)}=-2.1$.
+This numerical observation is also confirmed with sensitivities values.
+For a pole $z_{1}^{(-)}=-0.1$, we have $\rho_{1}^{(-)}=2\cdot 10^{-1}$ and $\eta_{1}^{(-)}=1.97$, and for a pole $z_{1}^{(+)}=-2.1$, we have $\rho_{1}^{(+)}=9.52\cdot 10^{-1}$ and $\eta_{1}^{(+)}=6.65$.
+Both unstructured and structured sensitivities are larger for a pole $z_{1}^{(+)}=-2.1$ located outside the unit circle.
+We compute also vectors of structured perturbations
+$(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T},$
+which means that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive with respect to the perturbation of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ with respect to the perturbation of the Fourier coefficient $\hat{r}_{2}$.
 Table 8: Sensitivities for Example 5.3.
@@
 2 $50$ $1.577\cdot 10^{4}$ $2.204\cdot 10^{3}$ $6.477\cdot 10^{3}$
+Example 5.2.
+We consider an example when poles are located very close to the unit circle. Let a rational function $r(z)$ of type $(7,8)$ in (1.1), i.e with $M=8$, be given by vectors of poles and coefficients
+$\boldsymbol{z}=(0.9,-0.9,0.9\mathrm{i},-0.9\mathrm{i},1.1,-1.1,1.1\mathrm{i},-1.1\mathrm{i}),\quad\boldsymbol{\gamma}=(1,2,3,4,5,6,7,8),$
+respectively. First, we consider recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we get the following reconstruction errors
+$e(\boldsymbol{z})=1.45\cdot 10^{-15},\quad e(\boldsymbol{\gamma})=5.48\cdot 10^{-14}.$
+[…]
+$N=4$ $\sigma=10^{-7}$ $\sigma=10^{-5}$
+min $7.38$e–$04$ $3.79$e–$02$
+max $7.38$e–$03$ $6.95$e–$01$
+average $4$e–$03$ $4.11$e–$01$
+In case of recovery from noisy data, we compute 10 iterations and present recovery result for each pole $z_{j}$ in Tables 3 and 4.
+Analyzing numerical results, we conclude that the pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations.
+This conclusion is confirmed with the sensitivities presented in Table 6.
+Results for the recovery of the coefficient vector are presented in Table 5.
+Table 10: $\ell_{p}$-norm and upper bound of $\ell_{p}$-norm of $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3 for $p=1,2$
+$\|\boldsymbol{\rho}^{(-)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{2}$) $\|\boldsymbol{\rho}^{(+)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{2}$)
+$N=4$ $2.86$e+$01$ $6.866$e+$01$ $2.204$e+$03$ $6.477$e+$03$
+$\|\boldsymbol{\rho}^{(-)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{1}$) $\|\boldsymbol{\rho}^{(+)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{1}$)
+$N=4$ $4.01$e+$01$ $5.725$e+$01$ $2.209$e+$03$ $6.802$e+$03$
+Example 5.3.
+Finally, let a rational function $r(z)$ of type $(3,4)$ in (1.1), i.e with $M=4$, be given by vectors of poles and coefficients
+$\boldsymbol{z}=(0.2,0.5,2,50),\quad\boldsymbol{\gamma}=(1,1,1,1),$
+respectively. For $N=4$ and $\sigma=0$, we get the following reconstruction errors
+$e(\boldsymbol{z})=1.42\cdot 10^{-13},\quad e(\boldsymbol{\gamma})=9.27\cdot 10^{-15}.$
+Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$.
+We compute 10 iterations and present the corresponding results in Table 7.
+Analyzing numerical results, we can conclude that the pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations.
+For $\sigma=10^{-5}$ the method does not reconstruct location of this pole correctly.
+This numerical observation is also predicted by the sensitivities values presented in Table 8.
+Both unstructured $\rho_{2}^{(+)}=2.204\cdot 10^{3}$ and structured $\eta_{2}^{(+)}=1.577\cdot 10^{4}$ sensitivities are sufficiently larger than sensitivities of other poles.
+Note, that the locations of the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed even for $\sigma=10^{-1}$.
+We present results regarding recovery of the coefficient vector in Table 9.
 We compute also vectors of structured perturbations for poles inside the unit circle
@@
 $(\boldsymbol{\mathcal{S}}^{(2)}_{4})^{(+)}=(54.25,2.929\cdot 10^{3},1.106\cdot 10^{4},1.085\cdot 10^{4})^{T}.$
+That means that poles $z_{1}^{(-)}=0.2$ and $z_{2}^{(-)}=0.5$ are the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-3}$, the pole $z_{1}^{(+)}=2$ – the Fourier coefficient $\hat{r}_{4}$, and the pole $z_{2}^{(+)}=50$ – the Fourier coefficient $\hat{r}_{3}$.
+In Subsection 4.2, we derived the upper bounds (4.224.24) for the $\ell_{2}$- and the $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and we want to understand how sharp these estimates are.
+In order to do it, we compute the exact values for the corresponding norms and the upper bounds obtained in (4.224.24) and present the corresponding results in Table 10.
+Comparing these results we can conclude that the estimates (4.224.24) are quite sharp.
+Table 11: Reconstruction results for $r$ from Example 5.1
+Method via Hankel pencil AAA method
 […]
-$N=4$ $\sigma=10^{-7}$ $\sigma=10^{-5}$
-min $7.38$e–$04$ $3.79$e–$02$
-max $7.38$e–$03$ $6.95$e–$01$
-average $4$e–$03$ $4.11$e–$01$
-That means that poles $z_{1}^{(-)}=0.2$ and $z_{2}^{(-)}=0.5$ are the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-3}$, the pole $z_{1}^{(+)}=2$ – the Fourier coefficient $\hat{r}_{4}$, and the pole $z_{2}^{(+)}=50$ – the Fourier coefficient $\hat{r}_{3}$.
-Table 11: $\ell_{2}$-norms and upper bounds of $\ell_{2}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3
-$\|\boldsymbol{\rho}^{(-)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{2}$) $\|\boldsymbol{\rho}^{(+)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{2}$)
-$N=4$ $2.86$e+$01$ $6.866$e+$01$ $2.204$e+$03$ $6.477$e+$03$
-Table 12: $\ell_{1}$-norms and upper bounds of $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3
-$\|\boldsymbol{\rho}^{(-)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{1}$) $\|\boldsymbol{\rho}^{(+)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{1}$)
-$N=4$ $4.01$e+$01$ $5.725$e+$01$ $2.209$e+$03$ $6.802$e+$03$
-Further we compute the $\ell_{1}$- and $\ell_{2}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bounds obtained in (4.21), (4.22) and (4.23).
-We present the results in Tables 11 and 12.
-Analyzing these results, we conclude that for this example upper bounds (4.21), (4.22) and (4.23) are quite sharp.
+min $1.52$e–$08$ $1.23$e–$08$ min $3.54$e–$08$ $4.55$e–$08$
+max $1.19$e–$07$ $5.51$e–$08$ max $6.62$e–$07$ $4.84$e–$07$
+average $7.51$e–$08$ $2.89$e–$08$ average $3.83$e–$07$ $2.82$e–$07$
+Method via Hankel pencil AAA method
+[…]
+min $4.34$e–$03$ $1.06$e–$03$ min $2.37$e–$02$ $2.27$e–$02$
+max $1.23$e–$02$ $4.24$e–$03$ max $2.05$e+$00$ $7.56$e–$02$
+average $7.64$e–$03$ $2.71$e–$03$ average $4.52$e–$01$ $4.55$e–$02$
+Table 12: Reconstruction results for $r$ from Example 5.2
+Method via Hankel pencil AAA method
+[…]
+min $1.96$e–$05$ $4.79$e–$06$ min $7.03$e–$05$ $7.18$e–$05$
+max $1.37$e–$04$ $6.22$e–$05$ max $4.49$e–$04$ $3.43$e–$04$
+average $5.89$e–$05$ $2.34$e–$05$ average $2.98$e–$04$ $2.25$e–$04$
+Method via Hankel pencil AAA method
+[…]
+min $2.2$e–$03$ $2.42$e–$03$ min $1.54$e–$02$ $9.73$e–$03$
+max $1.37$e–$02$ $4.86$e–$02$ max $6.69$e–$02$ $3.54$e–$02$
+average $6.93$e–$03$ $1.06$e–$02$ average $4.42$e–$02$ $5.75$e–$02$
+In Example 5.4 we present the comparison results of the performance of our approach with the AAA method for rational approximation.
+We employ $4N$ (noisy) function values $r(t_{k})(1+\varepsilon_{k})$ with $t_{k}=\mathrm{e}^{\frac{2\pi\mathrm{i}k}{4N}}$ for $k=0,\ldots,4N-1$, as input information for both methods.
+By $(\varepsilon_{k})_{k=0}^{4N-1}$ as above we denote the random noise that is generated in Matlab by $\sigma$*randn(4N), i.e., we take multiplicative noise drawn from the standard normal distribution.
+The corresponding modification of our method is discussed in Remark 4.1.
+We referee the reader to [14] regarding description of the AAA method.
+Note only that AAA takes the form of iteration with respect to a type of a rational approximant that is represented at each step in the barycentric form.
+In case of reconstruction of rational function of type $(M-1,M)$ using exact data, AAA terminates after $M+1$ iteration steps and returns $r$ as
+$r(t)=\sum\limits_{s=1}^{M+1}\frac{r(t_{k_{s}})w_{s}}{t-t_{k_{s}}}\bigg{/}\sum\limits_{s=1}^{M+1}\frac{w_{s}}{t-t_{k_{s}}},$
+where points $t_{k_{s}}$ are chosen by some greedy algorithm and weights $w_{s}$ are computed by solving a linear least-squares problem over the subset of sample points that have not been selected as support points, $\{t_{0},\ldots,t_{4N-1}\}\setminus\{t_{k_{1}},\ldots,t_{k_{M+1}}\}$ (see also [8, 9, 10] for details).
+The AAA method has the overall computational cost of $\mathcal{O}(NM^{3})$.
+Table 13: Reconstruction results for $r$ from Example 5.3
+Method via Hankel pencil AAA method
+[…]
+min $4.35$e–$03$ $7.99$e–$05$ min $1.26$e–$01$ $4.89$e–$03$
+max $3.00$e–$02$ $5.66$e–$04$ max $2.03$e+$00$ $7.86$e–$02$
+average $1.42$e–$02$ $2.66$e–$04$ average $1.48$e+$00$ $4.44$e–$02$
+Method via Hankel pencil AAA method
+[…]
+min $7.89$e–$04$ $4.48$e–$06$ min $2.6$e–$02$ $1.01$e–$03$
+max $1.28$e–$03$ $5.58$e–$05$ max $2.6$e–$01$ $1.01$e–$02$
+average $2.98$e–$03$ $2.4$e–$05$ average $1.25$e–$01$ $4.84$e–$03$
+Example 5.4.
+We consider again recovery of rational functions $r$ from Examples 5.1-5.3.
+We reconstruct $r$ from $4N$ samples $r(t_{k})(1+\varepsilon_{k})$, $k=0,\ldots,4N-1$ by using our method based on Hankel pencil and AAA.
+In Tables 11- 13 we present the results for $N=15$, $L=20$ and different values of $\sigma$.
+We see that for the rational function from Example 5.2 both methods work very well.
+Analysing results of Tables 11 and 13, we conclude that for $r$ from Examples 5.1 and 5.3 we get a slightly better recovery from noisy input data with our method.
+The reason for this is that the method based on the Hankel pencil works better for the recovery of poles with certain locations.
+For Example 5.3 the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed equally good by both methods, while the Hankel pencil based approach shows better performance for the recovery of the pole $z_{2}^{(+)}=50$ than the AAA method (see Tables 14).
+Table 14: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 with Hankel pencil and the AAA methods
+Method via Hankel pencil
+$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
+min $2.82$e–$08$ $6.98$e–$08$ $2.08$e–$07$ $4.35$e–$03$
+max $5.09$e–$07$ $3.92$e–$07$ $1.14$e–$06$ $3$e–$02$
+average $2.51$e–$07$ $1.79$e–$07$ $2.08$e–$07$ $1.72$e–$02$
+AAA method
+$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
+min $1.15$e–$07$ $6.42$e–$08$ $6.65$e–$07$ $1.26$e–$01$
+max $7.4$e–$07$ $5.24$e–$07$ $2.55$e–$06$ $2.03$e+$00$
+average $4.63$e–$07$ $2.34$e–$07$ $1.19$e–$06$ $1.14$e+$00$
 Acknowledgement
-The author gratefully acknowledge support by the German Research Foundation in the framework of the RTG 2088.
 The author would like to thank Ion Victor Gosea (MPI Magdeburg) for several valuable discussions regarding sensitivities of the eigenvalues of a matrix pencil problem.
+The author would also like to thank the anonymous referees for their valuable comments that helped to improve the manuscript.

### v3

@@
 Abstract
-In the paper, we develop a new method for the recovery of rational functions.
-Our idea is based on the property that Fourier coefficients of rational functions have the exponential structure and reconstruction of this exponential structure with the ESPRIT method in the frequency domain.
-Further we present sensitivity analysis for poles of rational functions reconstructed with our method in case of unstructured and structured perturbations.
+In this paper, we introduce a new approach for the recovery of rational functions.
+The concept we propose is based on using the exponential structure of the Fourier coefficients of rational functions and the reconstruction of this exponential structure in the frequency domain.
+We choose ESPRIT as a method for the exponential recovery.
+The matrix pencil structure of this approach is the reason for its selection, as it makes our method suitable for the sensitivity analysis.
+According to our method, poles located inside and outside the unit circle are reconstructed independently as eigenvalues of some special Hankel matrix pencils.
+Furthermore, we derived formulas for sensitivities of poles of rational functions in case of unstructured and structured perturbations.
 Finally, we consider several numerical experiments and, using sensitivities, explain the recovery errors for poles.
@@
 1 Introduction
-For $M\in\mathbb{N}$, we define a rational function $r(z)$ of type $(M-1,M)$ by means of its partial fraction decomposition
+For $M\in\mathbb{N}$, we define a rational function $r(z)$ of type (M1,M)(M-1,M) by means of its partial fraction decomposition
 $r(z):=\sum\limits_{j=1}^{M}\frac{\gamma_{j}}{z-z_{j}},$ (1.1)
-where $\gamma_{j},\,z_{j}\in\mathbb{C}$, $\gamma_{j}\neq 0$ and poles $z_{j}$ are pairwise distinct.
+where $\gamma_{j},\,z_{j}\in\mathbb{C}$, coefficients $\gamma_{j}\neq 0$ and poles $z_{j}$ are pairwise distinct.
 We study the question of recovery of rational functions (1.1) using as input data a finite set of their Fourier coefficients $\widehat{r}_{k}$ defined by (4.1).
-We use the property that Fourier coefficients of rational functions have the exponential structure and we reconstruct this exponential structure in the frequency domain applying the ESPRIT method [19, 17].
+We use the property that Fourier coefficients of rational functions have the exponential structure, and we reconstruct this exponential structure in the frequency domain applying the ESPRIT method [20, 18].
 The idea of recovery of rational functions by estimation of signal parameters in the frequency domain is not new.
@@
 The paper [22] is of special interest to us.
-In [21], the author considered recovery of the rational function (1.1) based on using noisy input data and discussed reconstruction results for the poles $z_{j}$, $j=1,...,M$.
+In [22], the author considered recovery of rational functions (1.1) based on using noisy input data and discussed reconstruction results for the poles $z_{j}$, j=1,,Mj=1,...,M.
 It was pointed out that for different levels of noise, not all poles $z_{j}$ are reconstructed with the same accuracy, and sometimes even the locations of certain poles are not reconstructed correctly.
-From these numerical observations, we can conclude that some poles $z_{j}$ of the rational function (1.1) are more ”sensitive“ to perturbations in the input data than the others.
+From these numerical observations, we can conclude that some poles $z_{j}$ of $r$ as in (1.1) are more ”sensitive“ to perturbations in the input data than the others.
 Our goal in this paper is to present theoretical explanations of these numerical observations by applying the sensitivity analysis.
@@
 In [24], the authors developed the sensitivity analysis for the eigenvalues of Loewner pencil with respect to the unstructured and structured perturbations.
-We referee the interested reader to [1] for further application of Loewner framework for model reduction of dynamical systems.
-Following approach in [23], we present the sensitivity analysis for the poles of $r$ in (1.1) computed by our method.
+We refer the interested reader to [1] for further application of Loewner framework for model reduction of dynamical systems.
+Following the approach in [24], we present the sensitivity analysis for the poles of $r$ in (1.1) computed by our method.
 According to our idea, poles located inside and outside the unit circle are reconstructed separately as eigenvalues of some Hankel matrix pencils.
@@
 In [12] and [3], the authors measured both structured and unstructured perturbations by the same quantity.
-Similarly as in [23], we introduce two different quantities for structured $\eta$ and unstructured $\rho$ sensitivities, respectively, of poles of the rational function (1.1).
-By checking if $\rho$ or $\eta$ is ”large“ for some pole $z_{j}$, we measure the ill-disposedness of $z_{j}$, while consideration of $\eta$ gives us also the additional information regarding that to perturbation of which measurement $\widehat{r}_{k}$ the pole $z_{j}$ is the most sensitive.
-The method developed in this paper is also closely related to our ESPIRA method for the exponential recovery [8][10].
-ESPIRA uses the property that the classical Fourier coefficients of exponential sums have a rational structure and reconstructs this rational structure with the AAA method for rational approximation [14] in the frequency domain.
-Thus, the idea of rational reconstruction used in this paper can be considered as converse to the idea of ESPIRA.
-We want to pay a special attention also to the paper [5], where authors developed a method for exponential reconstruction using the sensitivity analysis of Hankel pencil eigenvalues.
+Similarly as in [24], we introduce two different quantities, sensitivities $\eta$ and $\rho$ with respect to structured und unstructured perturbations, respectively, for poles of the rational function (1.1).
+The ill-disposedness of $z_{j}$ is measured by determining whether $\rho$ and $\eta$ are ”large“ for this specific pole $z_{j}$.
+Additionally, the sensitivity $\eta$ reveals the perturbations of which Fourier coefficient $\widehat{r}_{k}$ the pole $z_{j}$ is most sensitive to.
+The method developed in this paper is also closely related to our ESPIRA method for exponential recovery [9][11] (see also [8] for the multivariate case).
+ESPIRA uses the property that the classical Fourier coefficients of exponential sums have a rational structure and reconstructs this rational structure with the AAA method for rational approximation [15].
+Thus, the rational reconstruction concept employed in this paper can be viewed as the opposite of the ESPIRA concept.
+We would also like to draw particular attention to the paper [5], in which the authors used the sensitivity analysis of Hankel pencil eigenvalues to developed a method for exponential reconstruction.
 Finally, we want to mention the other methods for the recovery of rational functions such as the AAA method [15], the RKFIT (Rational Krylov Fitting) method [4], the Vector Fitting algorithm [14] and the Loewner framework method [1].
@@
 Since we are interested in the sensitivity analysis of the poles, we have to assume that we apply perturbations to the Fourier coefficients $\widehat{r}_{k}$.
-Note that we can introduce the corresponding modification of the method based on employing function value of $r$ by computing $\widehat{r}_{k}$ via the quadrature formula for some grid of points (see Remark 4.1).
-Modifying the algorithm in this way will allow us to compare its performance with known rational methods.
-Numerical experiments show us that our approach gives equally good and for some examples even better recovery results from given noisy function values than the AAA method.
+Note that in this paper, we also introduce the corresponding modification of our method, which is based on employing function values of $r$ by computing $\widehat{r}_{k}$ via a quadrature formula for some grid of points (see Remark 4.1).
+Modifying our algorithm in this way, allows us to compare its performance with known rational methods.
+Numerical experiments show us that our approach gives equally good and, for some examples, even better recovery results from given noisy function values than the AAA method.
 Outline.
@@
 Notation.
-Let us agree that throughout the paper, we will use the matrix notation $\mathbf{X}_{L,K}:=(X_{ij})_{i,j=1}^{L,K}$ for complex matrices of size $L\times K$ and the submatrix notation $\mathbf{X}_{L,K}(m:n,k:\ell)$, to denote a submatrix of $\mathbf{X}_{L,K}$ with rows indexed $m$ to $n$ and columns indexed $k$ to $\ell$, where first row and first column has index 1 (even though the row and column indices for the definition of the matrix may start with 0).
+As usual, $\mathbb{N}$, $\mathbb{Z}$, $\mathbb{R}$ and $\mathbb{C}$ denote the natural, integer, real and complex numbers, respectively.
+Let us agree that throughout the paper, we use the matrix notation $\mathbf{X}_{L,K}:=(X_{i,j})_{i,j=1}^{L,K}$ for matrices of the size $L\times K$ and the submatrix notation $\mathbf{X}_{L,K}(m:n,k:\ell)$, to denote a submatrix of $\mathbf{X}_{L,K}$ with rows indexed from $m$ to $n$ and columns indexed from $k$ to $\ell$, where first row and first column has index 1 (even though the row and column indices for the definition of the matrix may start with 0).
 For square matrices, we often use the short notation $\mathbf{X}_{K}$ instead of $\mathbf{X}_{K,K}$.
 By $\mathbf{X}_{L,K}^{\dagger}$ we denote the Moore-Penrose inverse of $\mathbf{X}_{L,K}$ and by $\mathbf{X}_{K}^{-1}$ inverse of the invertible matrix $\mathbf{X}_{K}$.
-We consider $\|\mathbf{X}_{L,K}\|_{2}:=\sigma_{\max}(\mathbf{X}_{L,K})$ the 2-norm or spectral norm, and $\|\mathbf{X}_{L,K}\|_{\mathrm{F}}:=\left(\sum\limits_{i=1}^{L}\sum\limits_{j=1}^{K}|X_{ij}|^{2}\right)^{1/2}=\left(\sum\limits_{i=1}^{\min\{L,K\}}\sigma_{i}^{2}\right)^{1/2}$ the Frobenius norm of the matrix $\mathbf{X}_{L,K}$.
-The $\ell_{1}$- and $\ell_{2}$-norm of the vector $\boldsymbol{x}=(x_{1},\ldots,x_{K})$ are given as usually by $\|\boldsymbol{x}\|_{1}=\sum\limits_{j=1}^{K}|x_{j}|$ and $\|\boldsymbol{x}\|_{2}=\left(\sum\limits_{j=1}^{K}|x_{j}|^{2}\right)^{1/2}$ respectively.
+We consider $\|\mathbf{X}_{L,K}\|_{2}:=\sigma_{\max}(\mathbf{X}_{L,K})$ the 2-norm or spectral norm, and $\|\mathbf{X}_{L,K}\|_{\mathrm{F}}:=\left(\sum_{i=1}^{L}\sum_{j=1}^{K}|X_{i,j}|^{2}\right)^{1/2}=\left(\sum_{i=1}^{\min\{L,K\}}\sigma_{i}^{2}\right)^{1/2}$ the Frobenius norm of the matrix $\mathbf{X}_{L,K}$.
+By $\boldsymbol{x}=(x_{1},\ldots,x_{K})^{T}$ we denote a column vector of the size $K$.
+The $\ell_{1}$- and $\ell_{2}$-norms of a vector $\boldsymbol{x}$ are given by $\|\boldsymbol{x}\|_{1}=\sum_{j=1}^{K}|x_{j}|$ and $\|\boldsymbol{x}\|_{2}=\left(\sum_{j=1}^{K}|x_{j}|^{2}\right)^{1/2}$, respectively.
 2 ESPRIT method for the recovery of exponential sums
-As was described in Introduction, we will apply the ESPRIT method [19, 17] to reconstruct the exponential structure of the Fourier coefficients of rational functions (1.1).
-In this section, we explain the main ideas of this method.
-Let $M\in\mathbb{N}$ and $f$ be an exponential sum defined by
-$f(t):=\sum\limits_{j=1}^{M}\gamma_{j}z_{j}^{t},\ \gamma_{j},z_{j}\in\mathbb{C},\ \gamma_{j}\neq 0,$ (2.1)
-where nodes $z_{j}$ are pairwise distinct.
-Let $L,N\in\mathbb{N}$ be given such that $M\leq L\leq N$, $L$ is the upper bound for $M$.
-We want to reconstruct an exponential sum (2.1), i.e.
-to reconstruct parameters $M$, $\gamma_{j}$ and $z_{j}$ for $j=1,\ldots,M$, using as input data samples $\boldsymbol{f}:=\big{(}f(0),f(1),\ldots,f(2N-1)\big{)}^{T}$.
+As previously mentioned in Introduction, we employ the ESPRIT method [20, 18] to reconstruct the exponential structure of the Fourier coefficients of rational functions (1.1).
+In this section, we describe the key concepts of this approach.
+Let $f$ be an exponential sum of order $M\in\mathbb{N}$ defined by
+$f(t):=\sum\limits_{j=1}^{M}\gamma_{j}z_{j}^{t},\ \gamma_{j},z_{j}\in\mathbb{C},$ (2.1)
+where nodes $z_{j}$ are pairwise distinct and coefficients $\gamma_{j}\neq 0$.
+Let $L,N\in\mathbb{N}$ be given such that $M\leq L\leq N$, and $L$ is the upper bound for $M$.
+Our goal is to reconstruct the exponential sum (2.1), i.e.
+to reconstruct parameters $M$, $\gamma_{j}$ and $z_{j}$ for $j=1,\ldots,M$, using as input data a vector of samples $\boldsymbol{f}:=\big(f(0),f(1),\ldots,f(2N-1)\big)^{T}$ of the size $2N$.
 For $n,m\in\mathbb{N}$ and a vector $\boldsymbol{a}:=(a_{1},\ldots,a_{m})^{T}$, we define a Vandermonde matrix $\mathbf{V}_{n,m}:=\mathbf{V}_{n,m}(\boldsymbol{a})=(a_{j}^{k})_{k=0,j=1}^{n-1,m}$.
@@
 Since Vandermonde matrices in (2.2) have full column ranks, the matrix $\mathbf{H}_{2N-L,L+1}$ has rank $M$.
-Therefore in case of exact given data, the order $M$ of the exponential sum (2.1) is reconstructed as the numerical rank of the matrix $\mathbf{H}_{2N-L,L+1}$.
+Thus, in case of exact given data, the order $M$ of the exponential sum (2.1) is reconstructed as the numerical rank of the matrix $\mathbf{H}_{2N-L,L+1}$.
 Consider further a Hankel matrix pencil problem
@@
 $z\,{\mathbf{H}}_{2N-L,L}(0)-{\mathbf{H}}_{2N-L,L}(1)=\mathbf{V}_{2N-L,M}(\boldsymbol{z})\,\mathrm{diag}\left(z-z_{j}\right)_{j=1}^{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{L,M}(\boldsymbol{z}).$
-From the last property, we conclude that nodes $z_{j}$, $j=1,\ldots,M$ in (2.1) are eigenvalues of the matrix pencil (2.3).
-Let us denote by $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ the vector of eigenvalues of (2.3).
+From the last formula, we conclude that nodes $z_{j}$, $j=1,\ldots,M$ in (2.1) are eigenvalues of the matrix pencil (2.3).
 In case when $M$ is known, we can choose $N=L=M$ and compute $z_{j}$ as eigenvalues of
 $z\mathbf{H}_{M}(0)-\mathbf{H}_{M}(1).$ (2.6)
-Now we discuss how we solve the matrix pencil (2.3). Consider the SVD of ${\mathbf{H}}_{2N-L,L+1}$,
+Next, we discuss how we solve the matrix pencil (2.3). Consider SVD of ${\mathbf{H}}_{2N-L,L+1}$,
 ${\mathbf{H}}_{2N-L,L+1}={\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,L+1}{\mathbf{W}}_{L+1},$ (2.7)
-where ${\mathbf{U}}_{2N-L}$ and ${\mathbf{W}}_{L+1}$ are unitary square matrices and, in the case of unperturbed data,
+where ${\mathbf{U}}_{2N-L}$ and ${\mathbf{W}}_{L+1}$ are unitary matrices and, in case of unperturbed data,
 ${\mathbf{D}}_{2N-L,L+1}=\left(\begin{array}[]{cc}\mathrm{diag}\left((\sigma_{j})_{j=1}^{M}\right)&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{0}}\end{array}\right)\in{\mathbb{R}}^{(2N-L)\times(L+1)},$
-where $\sigma_{1}\geq\ldots\geq\sigma_{M}$ denote the positive singular values of ${\mathbf{H}}_{2N-L,L+1}$. Thus (2.7) can be simplified to
+where $\sigma_{1}\geq\ldots\geq\sigma_{M}>0$ are singular values of ${\mathbf{H}}_{2N-L,L+1}$. Consequently, (2.7) can be simplified to
 ${\mathbf{H}}_{2N-L,L+1}={\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1},$ (2.8)
-where the (last) $L+1-M$ zero-columns in ${\mathbf{D}}_{2N-L,L+1}$ and the last $L+1-M$ rows in ${\mathbf{W}}_{L+1}$ are removed.
-Similarly as before, the SVDs of ${\mathbf{H}}_{2N-L,L}(0)$ and ${\mathbf{H}}_{2N-L,L}(1)$ can be directly derived from (2.8) and the matrix pencil in (2.3) now reads
+where the last L+1ML+1-M zero-columns in ${\mathbf{D}}_{2N-L,L+1}$ and the last L+1ML+1-M zero-rows in ${\mathbf{W}}_{L+1}$ are removed.
+Similarly as before, SVDs of ${\mathbf{H}}_{2N-L,L}(0)$ and ${\mathbf{H}}_{2N-L,L}(1)$ can be directly derived from (2.8) and the matrix pencil (2.3) now reads
 $z\left({\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1}\right)(1:2N-L,1:L)$
@@
 $z\,{\mathbf{W}}_{M,L+1}(1:M,1:L)-{\mathbf{W}}_{M,L+1}(1:M,2:L+1).$ (2.9)
-Thus, eigenvalues of the matrix pencil (2.3) can be computed as eigenvalues of the matrix pencil (2.9).
-Then we define ${\mathbf{W}}_{M,L}(0)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,1\!:\!L)$, ${\mathbf{W}}_{M,L}(1)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,2\!:\!L+1)$, and determine the vector $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ as vector of eigenvalues of $\left({\mathbf{W}}_{M,L}(0)^{T}\right)^{\dagger}{\mathbf{W}}_{M,L}(1)^{T}$.
-Having nodes $z_{j}$, we reconstruct coefficients $\gamma_{j}$, $j=1,\ldots,M$ as a least squares solution of the system of linear equations
+Therefore, eigenvalues of the matrix pencil (2.3) can be computed as eigenvalues of the matrix pencil (2.9).
+Then we define ${\mathbf{W}}_{M,L}(0)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,1\!:\!L)$ and ${\mathbf{W}}_{M,L}(1)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,2\!:\!L+1)$, and determine the vector $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ as a vector of eigenvalues of $\left({\mathbf{W}}_{M,L}(0)^{T}\right)^{\dagger}{\mathbf{W}}_{M,L}(1)^{T}$.
+Having nodes $z_{j}$, we compute coefficients $\gamma_{j}$, $j=1,\ldots,M$ as a least squares solution of the system of linear equations
 $\sum\limits_{j=1}^{M}\gamma_{j}z_{j}^{k}=f(k),\quad k=0,\ldots,2N-1,$
@@
 3 Sensitivities of Hankel pencil eigenvalues
-The concept of sensitivities arises from the perturbation of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [11] and further considered in [3].
+The concept of sensitivities arises from the perturbations of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [12] and further considered in [3].
 In [24], the notion of sensitivities was defined for Loewner matrix pencil.
@@
 Using similar ideas, in [25], the author considered the structured sensitivity of Hankel matrix pencil.
-In [5], the sensitivity analysis of Hankel pencil was applied to develop a new method for exponential analysis for the recovery from noisy data.
-In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3) and further in Sunbsection 4.2 we present modification of the results from this section to develop sensitivity analysis for the poles of rational functions.
-3.1 Sensitivity of Hankel pencil eigenvalues with unstructured perturbations
-Following approach offered in [23] for Loewner pencil, we will show that in case of unstructured perturbations of Hankel pencil, the sensitivity of eigenvalues depends on the numerical conditions of Vandermonde matrices from factorizations (2.4) and (2.5) (for $N=L=M$).
-We consider the generalized eigenvalue problem (GEP) for the Hankel matrix pencil (2.6),
+In [5], the sensitivity analysis of Hankel pencil was applied to develop a new method for the exponential analysis.
+In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3) and in Subsection 4.2, we present modification of these results to develop sensitivity analysis for the poles of rational functions (1.1).
+3.1 Sensitivities of Hankel pencil eigenvalues with unstructured perturbations
+Following the approach for Loewner pencil from [24], we show that in case of unstructured perturbations of Hankel pencil, the sensitivities of eigenvalues depend on the numerical conditions of Vandermonde matrices from factorizations (2.4) and (2.5) (for $N=L=M$).
+We consider the generalized eigenvalue problem (GEP) for the Hankel matrix pencil (2.6)
 $\boldsymbol{p}^{T}\mathbf{H}_{M}(1)=z\boldsymbol{p}^{T}\mathbf{H}_{M}(0),\quad\mathbf{H}_{M}(1)\boldsymbol{q}=z\mathbf{H}_{M}(0)\boldsymbol{q},$ (3.1)
-where $\boldsymbol{p}=(p_{1},\ldots,p_{M})^{T}$ and $\boldsymbol{q}=(q_{1},\ldots,q_{M})^{T}$ are right and left eigenvectors of the GEP (3.1) respectively. Then under the perturbation
-$\tilde{\mathbf{H}}_{M}(0):=\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)},\ \ \tilde{\mathbf{H}}_{M}(1):=\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)},$
+where $\boldsymbol{p}=(p_{1},\ldots,p_{M})^{T}$ and $\boldsymbol{q}=(q_{1},\ldots,q_{M})^{T}$ are left and right eigenvectors, respectively, of the GEP (3.1). Then under the perturbations
+$\tilde{\mathbf{H}}_{M}(0):=\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)},\ \ \tilde{\mathbf{H}}_{M}(1):=\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)},$ (3.2)
 the first order approximation of the eigenvalue perturbation $z^{(1)}$ is given by
 $z^{(1)}=\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}.$ (3.3)
-Formula (3.2) is obtained from the perturbed system
+Formula (3.3) is derived from the perturbed system
 $(\boldsymbol{p}^{T}+(\tilde{\boldsymbol{p}}^{(1)})^{T}+\ldots)$ $(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)})=(z+z^{(1)}+\ldots)(\boldsymbol{p}^{T}+(\tilde{\boldsymbol{p}}^{(1)})^{T}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{{\mathbf{H}_{M}(0)}}),$
@@
 by retaining only first-order terms (see [12], [3], [24] and [25] for details).
-Assume that the norm of perturbation matrices is bounded as follows
+Assume that the $\ell_{2}$-norms of perturbation matrices $\Delta_{\mathbf{H}_{M}(0)}$ and $\Delta_{\mathbf{H}_{M}(1)}$ in (3.2) are bounded as follows
 $\|\Delta_{\mathbf{H}_{M}(0)}\|_{2}\leq\epsilon\|\mathbf{H}_{M}(0)\|_{2},\ \ \|\Delta_{\mathbf{H}_{M}(1)}\|_{2}\leq\epsilon\|\mathbf{H}_{M}(1)\|_{2},$ (3.4)
-where $\varepsilon>0$. Then from (3.2) taking into account (3.3), we get for the GEP (3.1)
+for some $\varepsilon>0$. Then from (3.3) taking into account (3.4), we obtain
 $|z^{(1)}|$ $=\left|\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}\right|$
 $\leq\frac{\|\boldsymbol{p}^{T}\|_{2}\,\|\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)}\|_{2}\,\|\boldsymbol{q}\|_{2}}{|\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}|}\leq\epsilon\rho,$
-where $\rho$ is the sensitivity of the eigenvalue $z$ of the GEP (3.1) defined by
+where $\rho$ is the sensitivity with respect to unstructured perturbations of the eigenvalue $z$ of the GEP (3.1) defined by
 $\rho:=\frac{\|\boldsymbol{p}^{T}\|_{2}\,(|z|\,\|\mathbf{H}_{M}(0)\|_{2}+\|\mathbf{H}_{M}(1)\|_{2})\,\|\boldsymbol{q}\|_{2}}{|\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}|}.$ (3.5)
-We derive further formulas for the exact computation and upper bounds of sensitivities $\rho_{j}$ of eigenvalues $z_{j}$, $j=1,\ldots,M$ of the GEP (3.1).
-Since Hankel matrices $\mathbf{H}_{M}(0)$ and $\mathbf{H}_{M}(1)$ satisfy factorizations (2.4) and (2.5) for $N=L=M$, left $\boldsymbol{p}^{(j)}$ and right $\boldsymbol{q}^{(j)}$ eigenvectors of the GEP (3.1) corresponding to the eigenvalue $z_{j}$ are given by formulas
+Next, we derive formulas for the exact computation of sensitivities $\rho_{j}$ of the eigenvalues $z_{j}$, $j=1,\ldots,M$ of the GEP (3.1) as well as their upper bounds.
+Since the Hankel matrices $\mathbf{H}_{M}(0)$ and $\mathbf{H}_{M}(1)$ satisfy factorizations (2.4) and (2.5) for $N=L=M$, left $\boldsymbol{p}^{(j)}$ and right $\boldsymbol{q}^{(j)}$ eigenvectors corresponding to the eigenvalue $z_{j}$ can be determined by formulas
 $\boldsymbol{q}^{(j)}=\boldsymbol{p}^{(j)}=\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j},$ (3.6)
-where $\boldsymbol{e}_{j}$ is the $j$th canonical vector in $\mathbb{R}^{M}$. Then using again the factorization (2.4) of the Hankel matrix $\mathbf{H}_{M}(0)$, we get
+where $\boldsymbol{e}_{j}$ is the $j$th canonical column vector in $\mathbb{R}^{M}$. Then using again the factorization (2.4) of the Hankel matrix $\mathbf{H}_{M}(0)$, we get
 $(\boldsymbol{p}^{(j)})^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}^{(j)}$ $=\left(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\right)^{T}\,\mathbf{H}_{M}(0)\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}$
 $=\boldsymbol{e}_{j}^{T}\,\mathbf{\Gamma}_{M}\,\boldsymbol{e}_{j}=\gamma_{j}.$ (3.7)
-From (3.4) taking into account (3.6), we obtain that the sensitivity $\rho_{j}$ of the eigenvalue $z_{j}$ is given by
+From (3.5) taking into account (3.7), we obtain that the sensitivity $\rho_{j}$ is given by
 $\rho_{j}=\frac{1}{|\gamma_{j}|}\|(\boldsymbol{p}^{(j)})^{T}\|_{2}\,(|z_{j}|\,\|\mathbf{H}_{M}(0)\|_{2}+\|\mathbf{H}_{M}(1)\|_{2})\,\|\boldsymbol{q}^{(j)}\|_{2}.$ (3.8)
 The ill-disposedness of the eigenvalue $z_{j}$ with respect to unstructured perturbations can be measured by checking whether $\rho_{j}$ is ”large“.
-Applying formulas (2.4), (2.5) and (3.5) to (3.7), we obtain for $j=1,\ldots,M$
+Applying (2.4), (2.5) and (3.6) to (3.8), we obtain the following final formula for the computation of sensitivities $\rho_{j}$
 $\rho_{j}$ $=\frac{1}{|\gamma_{j}|}\left\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\right\|_{2}^{2}\left(|z_{j}|\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}+\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}\right).$ (3.9)
-The formula (3.8) is used to compute the exact values of sensitivities $\rho_{j}$.
-We proceed with estimation of (3.7). Since from (2.4) and (2.5), we have
-$\|\mathbf{H}_{M}(0)\|_{2}$ $\leq\|\mathbf{V}_{M}(\boldsymbol{z})\|_{2}\,\|\mathbf{\Gamma}_{M}\|_{2}\,\|\mathbf{V}^{T}_{M}(\boldsymbol{z})\|_{2},$
+We proceed with the estimation of (3.8). Since (2.4) and (2.5) implies
+$\|\mathbf{H}_{M}(0)\|_{2}$ $\leq\|\mathbf{V}_{M}(\boldsymbol{z})\|_{2}\,\|\mathbf{\Gamma}_{M}\|_{2}\,\|\mathbf{V}^{T}_{M}(\boldsymbol{z})\|_{2},$ (3.10)
 $\|\mathbf{H}_{M}(1)\|_{2}$ $\leq\|\mathbf{V}_{M}(\boldsymbol{z})\|_{2}\,\|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}\,\|\mathbf{V}^{T}_{M}(\boldsymbol{z})\|_{2},$ (3.11)
-and taking into account the fact $\|\mathbf{X}\,\boldsymbol{e}_{j}\|_{2}\leq\|\mathbf{X}\|_{2}$, we obtain
+and taking into account the inequality $\|\mathbf{X}\,\boldsymbol{e}_{j}\|_{2}\leq\|\mathbf{X}\|_{2}$, we conclude that
 $\rho_{j}\leq\frac{|z_{j}|\,\|\mathbf{\Gamma}_{M}\|_{2}+\|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}}{|\gamma_{j}|}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right),$
@@
 $\|\mathbf{\Gamma}_{M}\|_{2}=\max\limits_{k=1,\ldots,M}|\gamma_{k}|,\ \ \ \|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}=\max\limits_{k=1,\ldots,M}|z_{k}\gamma_{k}|.$
-Thus
+Thus, we derive the following upper bounds for sensitivities $\rho_{j}$,
 $\rho_{j}\leq\zeta_{j}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right),$ (3.12)
@@
 $\zeta_{j}:=\frac{1}{|\gamma_{j}|}\left(|z_{j}|\max\limits_{k=1,\ldots,M}|\gamma_{k}|+\max\limits_{k=1,\ldots,M}|\gamma_{k}z_{k}|\right).$ (3.13)
-Note that we can proceed further with the estimation of (3.10) applying known results for upper bounds of condition numbers of the Vandermonde matrices [15, 2].
-Define further a vector of unstructured sensitivities $\boldsymbol{\rho}:=(\rho_{1},\ldots,\rho_{M})^{T}\in\mathbb{R}^{M}$ and let $\boldsymbol{\zeta}:=(\zeta_{1},\ldots,\zeta_{M})^{T}\in\mathbb{R}^{M}$.
-Then from (3.10) for the $\ell_{2}$-norm of the vector $\boldsymbol{\rho}$ we get
+Note that we can estimate (3.12) further by applying known estimates for condition numbers of Vandermonde matrices [16, 2].
+Defining $\boldsymbol{\zeta}:=(\zeta_{1},\ldots,\zeta_{M})^{T}\in\mathbb{R}^{M}$ and the vector of sensitivities $\boldsymbol{\rho}:=(\rho_{1},\ldots,\rho_{M})^{T}\in\mathbb{R}^{M}$, and applying (3.12), we get for the $\ell_{2}$-norm of the vector $\boldsymbol{\rho}$,
 $\|\boldsymbol{\rho}\|_{2}\leq\|\boldsymbol{\zeta}\|_{2}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right).$
-From (3.8), taking into account (3.9) and (3.11), we obtain
+From (3.9), taking into account (3.10), (3.11) and (3.13), we obtain
 $\rho_{j}\leq\zeta_{j}\bigg\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\bigg\|_{2}^{2}\,\big\|\mathbf{V}_{M}(\boldsymbol{z})\big\|_{2}^{2}.$
-By summing up the last inequalities over $j=1,\ldots,M$, we get
-$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}$
-where $\zeta_{\max}:=\max_{j=1,\ldots,M}\zeta_{j}$. Finally using the property $\sum_{j=1}^{M}\|\mathbf{X}\boldsymbol{e}_{j}\|_{2}^{2}=\|\mathbf{X}\|_{\mathrm{F}}^{2}$, we obtain
+By summing up the last inequalities over $j=1,\ldots,M$, we arrive at
+$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big\|\mathbf{V}_{M}(\boldsymbol{z})\big\|_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\,\boldsymbol{e}_{j}\bigg\|_{2}^{2},$
+where $\zeta_{\max}:=\max_{j=1,\ldots,M}\zeta_{j}$.
+Finally, using the property $\sum_{j=1}^{M}\|\mathbf{X}\boldsymbol{e}_{j}\|_{2}^{2}=\|\mathbf{X}\|_{\mathrm{F}}^{2}$, we derive the following upper bound
 $\|\boldsymbol{\rho}\|_{1}\leq\zeta_{\max}\big\|\mathbf{V}_{M}(\boldsymbol{z})\big\|_{2}^{2}\,\big\|\mathbf{V}^{-1}_{M}(\boldsymbol{z})\big\|_{\mathrm{F}}^{2}$
 for the $\ell_{1}$-norm of the sensitivity vector $\boldsymbol{\rho}$.
-3.2 Sensitivity Hankel pencil eigenvalues with structured perturbations
-In this section we consider structured perturbations of the GEP (3.1) for the matrix pencil (2.6), i.e.
-the perturbation matrices have a particular structure, they are again Hankel matrices.
-We give a short description of the results obtained in [24].
-Let $\boldsymbol{\delta}_{2M}:=(\delta_{0},\delta_{1},\ldots,\delta_{2M-1})^{T}$ be a vector of perturbations and we define
-$\Delta_{\mathbf{H}_{M}(0)}:=(\delta_{k+\ell})_{k,\ell=0}^{M-1,M-1},\ \Delta_{\mathbf{H}_{M}(1)}:=(\delta_{k+\ell+1})_{k,\ell=0}^{M-1,M-1}.$
+3.2 Sensitivities of Hankel pencil eigenvalues with structured perturbations
+In this subsection, we consider structured perturbations of the GEP (3.1) for the matrix pencil (2.6).
+Now the perturbation matrices $\Delta_{\mathbf{H}_{M}(0)}$ and $\Delta_{\mathbf{H}_{M}(1)}$ in (3.2) have a particular structure, they are again Hankel matrices.
+We give a short description of the results from [25].
+Let $\boldsymbol{\delta}_{2M}:=(\delta_{0},\delta_{1},\ldots,\delta_{2M-1})^{T}$ be a vector of perturbations and let
+$\Delta_{\mathbf{H}_{M}(0)}=(\delta_{k+\ell})_{k,\ell=0}^{M-1,M-1},\ \Delta_{\mathbf{H}_{M}(1)}=(\delta_{k+\ell+1})_{k,\ell=0}^{M-1,M-1}.$
 For the first order approximation of the eigenvalue perturbation $z_{j}^{(1)}$, $j=1,\ldots,M$ from (3.3) using (3.7), we obtain
@@
 $\left((\boldsymbol{p}^{(j)})^{T}\,\mathbf{Q}_{M,2M}^{(j)}(1)\right)_{m}$ $=\sum\limits_{k=1}^{M}p^{(j)}_{k}\,Q_{k,m}^{(j)}(1)=\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}p^{(j)}_{k}\,q^{(j)}_{m-k}$,
-taking into account (3.12), we define vectors $\boldsymbol{\mathcal{S}}^{(j)}_{2M}:=\left(S^{(j)}_{1},\ldots,S^{(j)}_{2M}\right)^{T}\in\mathbb{C}^{2M}$, $j=1,\ldots,M$ by their components
+and taking into account (3.14), we define vectors $\boldsymbol{\mathcal{S}}^{(j)}_{2M}:=\left(S^{(j)}_{1},\ldots,S^{(j)}_{2M}\right)^{T}\in\mathbb{C}^{2M}$, $j=1,\ldots,M$ by their components
 $S^{(j)}_{m}:=\frac{1}{\gamma_{j}}\left(\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}p^{(j)}_{k}\,q^{(j)}_{m-k}-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}p^{(j)}_{k}\,q^{(j)}_{m-k+1}\right)$,
-$m=1,\ldots,2M$. Then the first order approximation of the eigenvalue perturbation $z_{j}^{(1)}$ in (3.12) can be given by
-$z_{j}^{(1)}=(\boldsymbol{\mathcal{S}}^{(j)}_{2M})^{T}\,\boldsymbol{\delta}_{2M},\quad j=1,\ldots,M.$ (3.13)
-The component $S^{(j)}_{m}$ shows the influence of the perturbation of the $m$th measurement (in this case it is the function value $f(m-1)$) on the $j$th eigenvalue $z_{j}$ of the Hankel matrix pencil (2.6).
-Using (3.5), we rewrite components $S^{(j)}_{m}$ in the following way
-$S^{(j)}_{m}=$ $\frac{1}{\gamma_{j}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{k,j}\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg{)}_{m-k,j}$
+for $m=1,\ldots,2M$. Then $z_{j}^{(1)}$ as in (3.14) can be presented by
+$z_{j}^{(1)}=(\boldsymbol{\mathcal{S}}^{(j)}_{2M})^{T}\,\boldsymbol{\delta}_{2M}.$ (3.15)
+The component $S^{(j)}_{m}$ shows the influence of the perturbations of the $m$th measurement (in this case it is the function value f(m1)f(m-1)) on the $j$th eigenvalue $z_{j}$ of the Hankel matrix pencil (2.6).
+Using (3.6), we rewrite components $S^{(j)}_{m}$ as follows
+$S^{(j)}_{m}$= $\frac{1}{\gamma_{j}}\bigg(\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\bigg({\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}}\bigg)_{k,j}\,\bigg(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg)_{m-k,j}$
 $-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}\bigg(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg)_{k,j}\,\bigg(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{-1}\bigg)_{m-k+1,j}\bigg).$
@@
 $\boldsymbol{z}^{(1)}=\mathbf{\mathcal{S}}_{M,2M}\,\boldsymbol{\delta}_{2M},$
-where $\boldsymbol{z}^{(1)}:=(z_{1}^{(1)},\ldots,z_{M}^{(1)})^{T}$ and $\mathbf{\mathcal{S}}_{M,2M}:=(S_{j,m})_{j,m=1}^{M,2M}\in\mathbb{C}^{M\times 2M}$ be a matrix with the entries $S_{j,m}:=S^{(j)}_{m}$ that quantify the perturbation of the eigenvalues of matrix pencil (2.6).
-Finally, we define a structured sensitivity of the eigenvalue $z_{j}$ of matrix pencil (2.6) with respect to the $m$th measurement by
-$\eta_{j,m}:=|S_{j,m}|$
+where $\boldsymbol{z}^{(1)}:=(z_{1}^{(1)},\ldots,z_{M}^{(1)})^{T}$ and $\mathbf{\mathcal{S}}_{M,2M}:=(S_{j,m})_{j,m=1}^{M,2M}\in\mathbb{C}^{M\times 2M}$ be a matrix with the entries $S_{j,m}:=S^{(j)}_{m}$, which quantify the perturbation of the eigenvalues of the matrix pencil (2.6).
+Finally, we define the sensitivity $\eta_{j,m}$ of the eigenvalue $z_{j}$ of the matrix pencil (2.6) with respect to the $m$th measurement by
+$\eta_{j,m}:=|S_{j,m}|,$
 for $j=1,\ldots,M$ and $m=1,\ldots,2M$.
-Then the eigenvalue sensitivity $\eta_{j}$ for the eigenvalue $z_{j}$ with respect to structured perturbations of Hankel matrix pencil (2.6) is defined as the $\ell_{2}$-norm of the vector $\boldsymbol{\mathcal{S}}^{(j)}_{2M}\in\mathbb{C}^{2M}$.
-Thus we further introduce a vector $\boldsymbol{\eta}:=(\eta_{1},\ldots,\eta_{M})^{T}\in\mathbb{R}^{M}$ with its entries
+By $\boldsymbol{\eta}_{2M}^{(j)}:=(\eta_{j,1},\ldots,\eta_{j,2M})^{T}\in\mathbb{R}^{2M}$ we denote the corresponding vector of sensitivities for the eigenvalue $z_{j}$.
+Then the sensitivity $\eta_{j}$ with respect to structured perturbations for the eigenvalue $z_{j}$ of the Hankel matrix pencil (2.6) is defined as the $\ell_{2}$-norm of the vector $\boldsymbol{\mathcal{S}}^{(j)}_{2M}\in\mathbb{C}^{2M}$.
+Thus, we further define a vector $\boldsymbol{\eta}:=(\eta_{1},\ldots,\eta_{M})^{T}\in\mathbb{R}^{M}$ by its components
 $\eta_{j}:=\|\boldsymbol{\mathcal{S}}^{(j)}_{2M}\|_{2}=\sqrt{\eta_{j,1}^{2}+\ldots+\eta^{2}_{j,2M}},\quad j=1,\ldots,M.$
-The ill-disposedness of the eigenvalue $z_{j}$ with respect to structured perturbations can be measured by checking whether $\eta_{j}$ is "large".
+The ill-disposedness of the eigenvalue $z_{j}$ with respect to structured perturbations can be measured by checking whether $\eta_{j}$ is ”large“.
 The entry $\eta_{j}$ determines also the standard deviation of the eigenvalue perturbation $z_{j}^{(1)}$ (see [25] for details).
 4 Recovery of rational functions via Hankel matrix pencil method and sensitivities of the poles
-In this section, we describe the main ideas of our new method for the recovery of a rational function of the form (1.1) and present the sensitivity analysis for its poles.
+In this section, we describe the main ideas of our new method for the recovery of rational functions (1.1) and present the sensitivity analysis for the poles.
 4.1 Description of the method
-Let $r(t)$ be a rational function defined by (1.1) and by $\mathbb{D}$ we denote the unit circle. Following approach in [21], for $k\in\mathbb{Z}$ we consider the integral
+Let $r(t)$ be a rational function defined by (1.1) and by $\mathbb{D}:=\{z\in\mathbb{C}:\,|z|=1\}$ we denote the unit circle in the complex plane.
+Following approach in [22], for $k\in\mathbb{Z}\setminus\{0\}$ we consider the integral
 $\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t},$
-which is closely related to the classical Fourier coefficient of a function $r(\mathrm{e}^{\mathrm{i}\varphi})$,
+which is closely related to the classical Fourier coefficient $\widehat{r}_{k}$ of the function $r(\mathrm{e}^{\mathrm{i}\varphi})$,
 $\widehat{r}_{k}:=\frac{1}{2\pi}\int\limits_{0}^{2\pi}r(\mathrm{e}^{\mathrm{i}\varphi})\mathrm{e}^{-\mathrm{i}k\varphi}\,\mathrm{d}\varphi=\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t}.$ (4.1)
 Then (see, for example, [22])
-$\widehat{r}_{k}=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{-(k+1)},&k\leq-1,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{-(k+1)},&k\geq 1.\end{cases}$ (4.2)
-From (4.2), we conclude that Fourier coefficients $\widehat{r}_{k}$ with negative indeces $k=-1,-2,\ldots$ contain information about poles inside the unite circle, and Fourier coefficients $\widehat{r}_{k}$ with positive indises $k=1,2,\ldots$ contain information about poles outside the unite circle.
+$\widehat{r}_{k}=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{-(k+1)},&k=-1,-2,\ldots,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{-(k+1)},&k=1,2,\ldots.\end{cases}$ (4.2)
+From (4.2), we conclude that Fourier coefficients $\widehat{r}_{k}$ with negative indices $k=-1,-2,\ldots$ contain information about poles located inside the unite circle, and Fourier coefficients $\widehat{r}_{k}$ with positive indices $k=1,2,\ldots$ contain information about poles located outside the unite circle.
 Taking into account (4.2), we define an exponential sum $f(t)$ by
-$f(t)=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{t},&t\geq 0,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{t},&t\leq-2,\end{cases}$ (4.3)
+$f(t)=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{t},&t\geq 0,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{t},&t<0,\end{cases}$ (4.3)
 and from (4.2) we obtain the following interpolation property
 $f(-(k+1))=\widehat{r}_{k},\quad k\in\mathbb{Z}\setminus\{0\}.$ (4.4)
-Therefore the rational reconstruction problem (1.1) can be reformulated as the exponential interpolation problem (4.4).
-Our further goal is to reconstruct an exponential sum (4.3), using samples $f(m)$, $m\in\mathbb{Z}\setminus\{-1\}$ as input data.
-From (4.3), we get that values $f(m)$ with $m=0,1,2,\ldots$ contain information about nodes $z_{j}$ inside $\mathbb{D}$, and values $f(m)$ with $m=-2,-3,\ldots$ contain information about nodes $z_{j}$ outside $\mathbb{D}$.
-Let $M_{1}$ and $M_{2}$ be numbers of nodes $z_{j}$ of the exponential sum (4.3) located inside $\mathbb{D}$ and outside $\mathbb{D}$ respectively.
-We assume that a number of poles $M$ as well as $M_{1}$ and $M_{2}$ are also unknown.
-Let $L,N\in\mathbb{N}$ and $M_{1},M_{2}\leq L\leq N$, where $L$ is the upper bound for $M$.
-We use values $f(k)$, $k=0,\ldots,2N-1$ ($\widehat{r}_{-k}$, $k=1,\ldots,2N$) to reconstruct nodes $z_{j}$ inside $\mathbb{D}$, and values $f(-k)$, $k=2,\ldots,2N+1$ ($\widehat{r}_{k}$, $k=1,\ldots,2N$) to reconstruct nodes $z_{j}$ outside $\mathbb{D}$.
-Further we rewrite an exponential sum (4.3) in the following way
-$f(t)=\begin{cases}\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{j_{n}}(z^{(-)}_{j_{n}})^{t},&t\geq 0,\\ -\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{j_{n}}(z^{(+)}_{j_{n}})^{t}&t\leq-2,\end{cases}$ (4.5)
+Thus, the rational reconstruction problem (1.1) can be reformulated as the exponential interpolation problem (4.4).
+Our further goal is to reconstruct the exponential sum (4.3), using samples $f(m)$, $m\in\mathbb{Z}\setminus\{-1\}$ as inputs.
+From (4.3) and (4.4), we conclude that values $f(m)$ with $m=0,1,\ldots$ contain information about nodes $z_{j}$ located inside $\mathbb{D}$, and values $f(m)$ with $m=-2,-3,\ldots$ contain information about nodes $z_{j}$ located outside $\mathbb{D}$.
+Let $M_{1}$ and $M_{2}$ be the numbers of nodes $z_{j}$ of $f$ as in (4.3) located inside $\mathbb{D}$ and outside $\mathbb{D}$, respectively. Thus, we can rewrite the exponential sum (4.3) as follows
+$f(t)=\begin{cases}\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{n}(z^{(-)}_{n})^{t},&t\geq 0,\\ -\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{n}(z^{(+)}_{n})^{t}&t<0,\end{cases}$ (4.5)
 i.e.
-by $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ we denote poles located inside $\mathbb{D}$ and let $\gamma^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ be the corresponding coefficients and by $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$ we denote poles located outside $\mathbb{D}$ and let $\gamma^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$ be the corresponding coefficients.
-To reconstruct exponential sums (4.5), we use the ESPRIT method described in Section 2.
-We present some modification of this method for our special parameter estimation problem (4.5).
-Note that the ESPRIT is not the only method to apply in the frequency domain.
-We can use the other methods for exponential recovery, for example, MPM (matrix-pencil method) [17] or approximate Prony method [18].
-In [21], the author uses the AAK theory [16] in the frequency domain to reconstruct (1.1) and a matrix-valued version of (1.1).
-Let $\boldsymbol{z}^{(-)}\in\mathbb{C}^{M_{1}}$ and $\boldsymbol{z}^{(+)}\in\mathbb{C}^{M_{2}}$ be two vectors defined by
-$\boldsymbol{z}^{(-)}:=(z^{(-)}_{j_{1}},\ldots,z^{(-)}_{j_{M_{1}}}),\ \ \boldsymbol{z}^{(+)}:=\left(\frac{1}{z_{j_{1}}^{(+)}},\ldots,\frac{1}{z^{(+)}_{j_{M_{2}}}}\right).$
-Let us define also diagonal matrices
-$\mathbf{\Gamma}^{(-)}_{M_{1}}:=\mathrm{diag}(\gamma^{(-)}_{j_{n}})_{n=1}^{M_{1}},\ \mathbf{\Gamma}^{(+)}_{M_{2}}:=\mathrm{diag}(\gamma^{(+)}_{j_{n}})_{n=1}^{M_{2}},\ \mathbf{Z}_{M_{1}}^{(+)}:=\mathrm{diag}\left(z^{(-)}_{j_{n}}\right)_{n=1}^{M_{1}},\ \mathbf{Z}_{M_{2}}^{(+)}:=\mathrm{diag}\left(\frac{1}{z^{(+)}_{j_{n}}}\right)_{n=1}^{M_{2}}.$
-According to the ESPRIT method described in Section 2, to reconstruct nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$, located inside the unit circle, we use the Hankel matrix
+by $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$, and $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$, we denote nodes located inside $\mathbb{D}$ and outside $\mathbb{D}$, respectively, and let $\gamma^{(-)}_{n}$, $n=1,\ldots,M_{1}$, and $\gamma^{(+)}_{n}$, $n=1,\ldots,M_{2}$, be the corresponding coefficients.
+Next, we present some modification of the ESPRIT method to solve our special parameter estimation problem (4.5).
+Note that ESPRIT is not the only method to apply in the frequency domain.
+We can use also the other methods for the exponential recovery, for example, MPM (matrix-pencil method) [18], approximate Prony method [19], or AAK theory [17] (as in [22]).
+By $\boldsymbol{z}^{(-)}\in\mathbb{C}^{M_{1}}$ and $\boldsymbol{z}^{(+)}\in\mathbb{C}^{M_{2}}$ we denote two vectors
+$\boldsymbol{z}^{(-)}:=(z^{(-)}_{1},\ldots,z^{(-)}_{M_{1}})^{T},\ \ \boldsymbol{z}^{(+)}:=\left(\frac{1}{z_{1}^{(+)}},\ldots,\frac{1}{z^{(+)}_{M_{2}}}\right)^{T},$
+and let us also define the following diagonal matrices
+$\mathbf{\Gamma}^{(-)}_{M_{1}}:=\mathrm{diag}(\gamma^{(-)}_{n})_{n=1}^{M_{1}},\ \mathbf{\Gamma}^{(+)}_{M_{2}}:=\mathrm{diag}(\gamma^{(+)}_{n})_{n=1}^{M_{2}},\ \mathbf{Z}_{M_{1}}^{(+)}:=\mathrm{diag}\left(z^{(-)}_{n}\right)_{n=1}^{M_{1}},\ \mathbf{Z}_{M_{2}}^{(+)}:=\mathrm{diag}\left(\frac{1}{z^{(+)}_{n}}\right)_{n=1}^{M_{2}}.$
+Let further $L,N\in\mathbb{N}$ be given such that $M_{1},M_{2}\leq L\leq N$, and $L$ is the upper bound for $M$.
+We use values $f(k)$, $k=0,\ldots,2N-1$ ($\widehat{r}_{-k}$, $k=1,\ldots,2N$) for recovering nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$ located inside $\mathbb{D}$, and values f(k)f(-k), $k=2,\ldots,2N+1$ ($\widehat{r}_{k}$, $k=1,\ldots,2N$) for recovering nodes $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$ located outside $\mathbb{D}$.
+According to the ESPRIT method described in Section 2, to reconstruct nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$, and $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$, we employ the Hankel matrices
 $\mathbf{H}^{(-)}_{2N-L,L+1}:=\left(f(\ell+k)\right)_{k,\ell=0}^{2N-L-1,L}=\left(\widehat{r}_{-(k+\ell)}\right)_{k=0,\ell=1}^{2N-L-1,L+1},$
-and to reconstruct nodes $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$, located outside the unit circle, we use the Hankel matrix
-$\mathbf{H}^{(+)}_{2N-L,L+1}:=(f(-(\ell+k+2)))_{k,\ell=0}^{2N-L-1,L}=\left(\widehat{r}_{k+\ell}\right)_{k=0,\ell=1}^{2N-L-1,L+1}.$
-Consider factorizations of these matrices
+and
+$\mathbf{H}^{(+)}_{2N-L,L+1}:=(f(-(\ell+k+2)))_{k,\ell=0}^{2N-L-1,L}=\left(\widehat{r}_{k+\ell}\right)_{k=0,\ell=1}^{2N-L-1,L+1},$
+respectively. We consider factorizations of these matrices
 $\mathbf{H}^{(-)}_{2N-L,L+1}$ $=\mathbf{V}_{2N-L,M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}^{(-)}_{M_{1}}\,\mathbf{V}^{T}_{L+1,M_{1}}(\boldsymbol{z}^{(-)}),$ (4.6)
 $\mathbf{H}^{(+)}_{2N-L,L+1}$ $=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L+1,M_{2}}(\boldsymbol{z}^{(+)}).$ (4.7)
-Since all Vandermonde matrices in (4.6) and (4.7) have full column ranks, the matrix $\mathbf{H}_{2N-L,L+1}^{(-)}$ has rank $M_{1}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ has rank $M_{2}$.
-Therefore in case of exact given data, the numbers $M_{1}$ and $M_{2}$ are reconstructed as numerical ranks of matrices $\mathbf{H}_{2N-L,L+1}^{(-)}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ respectively.
-First we consider the reconstruction of nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$, located inside the $\mathbb{D}$.
-Since in this case, the Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ have structure similar to the structure of the Hankel matrices $\mathbf{H}_{2N-L,L+1}$ in the Section 2, following ideas of the Section 2, we get that nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ are reconstructed as eigenvalues of the Hankel matrix pencil
+Since all Vandermonde matrices in (4.6) and (4.7) have full column ranks, $\mathbf{H}_{2N-L,L+1}^{(-)}$ has rank $M_{1}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ has rank $M_{2}$.
+Therefore, in case of exact given data, the numbers $M_{1}$ and $M_{2}$ are reconstructed as numerical ranks of matrices $\mathbf{H}_{2N-L,L+1}^{(-)}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$, respectively.
+Let us start with the case of nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$, located inside $\mathbb{D}$.
+Since the Hankel matrix $\mathbf{H}^{(-)}_{2N-L,L+1}$ have the structure similar to those of the Hankel matrix $\mathbf{H}_{2N-L,L+1}$ as in (2.2), following ideas of Section 2, we conclude that nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$ can be computed as eigenvalues of the Hankel matrix pencil
 $z\mathbf{H}^{(-)}_{2N-L,L}(0)-\mathbf{H}^{(-)}_{2N-L,L}(1),$ (4.8)
-where for $s=0,1$
+where for $s=0,1$,
 $\mathbf{H}^{(-)}_{2N-L,L}(s):=\mathbf{H}^{(-)}_{2N-L,L+1}(1:2N-L,1+s:L+s)=(f(\ell+k+s))_{k,\ell=0}^{2N-L-1,L-1}.$
-Consider now the case of nodes $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$, located outside the $\mathbb{D}$. For $s=0,1$, define matrices
+Consider now the case of nodes $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$, located outside $\mathbb{D}$. For $s=0,1$, define matrices
 $\mathbf{H}^{(+)}_{2N-L,L}(s):=\mathbf{H}^{(+)}_{2N-L,L+1}(1:2N-L,1+s:L+s)=(f(-(\ell+k+s+2)))_{k,\ell=0}^{2N-L-1,L-1},$
-and consider their factorizations
+and write their factorizations
 $\mathbf{H}^{(+)}_{2N-L,L}(0)$ $=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}),$ (4.9)
 $\mathbf{H}^{(+)}_{2N-L,L}(1)$ $=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{3}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}).$ (4.10)
-Since according to (4.9) and (4.10) we have
-$z\,{\mathbf{H}}^{(+)}_{2N-L,L}(0)-{\mathbf{H}}^{(+)}_{2N-L,L}(1)=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathrm{diag}\left(z-\frac{1}{z^{(+)}_{j_{n}}}\right)_{n=1}^{M_{2}}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}),$
-then inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of nodes $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ are reconstructed as eigenvalues of the Hankel matrix pencil
+According to (4.9) and (4.10) we have
+$z\,{\mathbf{H}}^{(+)}_{2N-L,L}(0)-{\mathbf{H}}^{(+)}_{2N-L,L}(1)=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathrm{diag}\left(z-\frac{1}{z^{(+)}_{n}}\right)_{n=1}^{M_{2}}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}),$
+Thus, the inverse $\frac{1}{z^{(+)}_{n}}$ of nodes $z^{(+)}_{n}$ for $n=1,\ldots,M_{2}$ are computed as eigenvalues of the Hankel matrix pencil
 $z\mathbf{H}^{(+)}_{2N-L,L}(0)-\mathbf{H}^{(+)}_{2N-L,L}(1).$ (4.11)
-If we assume that $M_{1}$ and $M_{2}$ are known, we can reconstruct nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ as eigenvalues of the Hankel matrix pencil ($N=L=M_{1}$ in (4.8))
+If we assume that $M_{1}$ and $M_{2}$ are known, we can reconstruct nodes $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$ as eigenvalues of the Hankel matrix pencil ($N=L=M_{1}$ in (4.8))
 $z\mathbf{H}^{(-)}_{M_{1}}(0)-\mathbf{H}^{(-)}_{M_{1}}(1),$ (4.12)
-and inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of nodes $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ are reconstructed as eigenvalues of the Hankel matrix pencil ($N=L=M_{2}$ in (4.11))
+and we can reconstruct the inverse $\frac{1}{z^{(+)}_{n}}$ of nodes $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$, as eigenvalues of the Hankel matrix pencil ($N=L=M_{2}$ in (4.11))
 $z\mathbf{H}^{(+)}_{M_{2}}(0)-\mathbf{H}^{(+)}_{M_{2}}(1).$ (4.13)
-Having nodes $z^{(-)}_{j_{n}}$ and $z^{(+)}_{j_{n}}$, coefficients $\gamma^{(-)}_{j_{n}}$ and $\gamma^{(+)}_{j_{n}}$ can be computed as least squares solutions of the overdetermined systems of linear equations
-$\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{j_{n}}(z^{(-)}_{j_{n}})^{-(k+1)}=\widehat{r}_{-k},\quad k=1,\ldots,2N,$
+Having nodes $z^{(-)}_{n}$ and $z^{(+)}_{n}$, coefficients $\gamma^{(-)}_{n}$ and $\gamma^{(+)}_{n}$ can be computed as least squares solutions of the overdetermined systems of linear equations
+$\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{n}(z^{(-)}_{n})^{-(k+1)}=\widehat{r}_{-k},\quad k=1,\ldots,2N,$
 and
-$-\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{j_{n}}(z^{(+)}_{j_{n}})^{-(k+1)}=\widehat{r}_{k},\quad k=1,\ldots,2N,$
+$\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{n}(z^{(+)}_{n})^{-(k+1)}=\widehat{r}_{k},\quad k=1,\ldots,2N,$
 respectively.
@@
 1.
-Note that coefficients $\hat{r}_{k}$ can be also computed applying the quadrature formulas on some grid of points.
-In that case, we will use the function values of $r$ instead of its Fourier coefficients as input data for the algorithm.
-Let us chose points $t_{n}=\mathrm{e}^{\frac{2\pi\mathrm{i}n}{4N}}$, $n=0,...,4N-1$ and let $(R_{k})_{k=0}^{4N-1}$ be the DFT of $(r(t_{n}))_{n=0}^{4N-1}$.
-Then $\hat{r}_{k}$ in (4.1) can be computed as $\hat{r}_{k}=R_{k+2N}$, $k=-2N,...,2N-1$.
-This question was discussed in details in [21], where the author consider not only the points on the unit circle, but also the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$.
+Note that Fourier coefficients $\hat{r}_{k}$ can also be computed applying quadrature formulas on some grid of points.
+In that case, we use the function values of $r$ instead of its Fourier coefficients as inputs for the algorithm.
+Let us chose points $t_{n}=\mathrm{e}^{\frac{2\pi\mathrm{i}n}{4N}}$, n=0,,4N1n=0,...,4N-1 and let $(R_{k})_{k=0}^{4N-1}$ be the DFT of $(r(t_{n}))_{n=0}^{4N-1}$.
+Then $\hat{r}_{k}$ in (4.1) can be determined by $\hat{r}_{k}=R_{k+2N}$ for k=2N,,2N1k=-2N,...,2N-1.
+This question was discussed in details in [22], where the author considered not only the points on the unit circle, but also the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$.
 On the one hand, we are interested in the presentation of a modification of the method from [22] and in the explanation of the results of this method for the recovery from noisy data with the use of the sensitivities.
@@
 2.
-The method based on using the given coefficients $\hat{r}_{k}$, $k=-2N,...,2N$ has the overall computational cost of $\mathcal{O}(NL^{2})$, which is determined by computation of SVD of Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ and $\mathbf{H}^{(+)}_{2N-L,L+1}$.
-The complexity of the modified algorithm based on using the values $(r(t_{n}))_{n=0}^{4N-1}$ remains the same, since the additional operation computation of FFT of $(r(t_{n}))_{n=0}^{4N-1}$ requires only $\mathcal{O}(N\log N)$ flops.
+The method based on using the given coefficients $\hat{r}_{k}$, k=2N,,2Nk=-2N,...,2N has the overall computational cost of $\mathcal{O}(NL^{2})$ flops, which is determined by the complexity of the computation of SVDs of Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ and $\mathbf{H}^{(+)}_{2N-L,L+1}$.
+The complexity of the modified algorithm based on employing the values $(r(t_{n}))_{n=0}^{4N-1}$ is $\mathcal{O}(N(L^{2}+\log N))$, since the additional step of the computation of FFT of $(r(t_{n}))_{n=0}^{4N-1}$ requires $\mathcal{O}(N\log N)$ operations.
 4.2 Sensitivities of poles
-In this subsection, we consider sensitivities of poles $z_{j}$ of the reconstruction problem (1.1) that is solved with the method described in Subsection 4.1.
-First we consider the case of unstructured perturbations.
-Poles $z_{j_{n}}^{(-)}$, $n=1,\ldots,M_{1}$ located inside the unit circle $\mathbb{D}$, are reconstructed as eigenvalues of the matrix pencil (4.12).
-Since Hankel matrices in the matrix pencil (4.12) have similar structure as Hankel matrices in the matrix pencil (2.6), we apply ideas from Subsection 3.1 without any changes and obtain for the sensitivities $\rho_{n}^{(-)}$, $n=1,\ldots,M_{1}$ of the poles $z^{(-)}_{j_{1}},\ldots,z^{(-)}_{j_{M_{1}}}$ in the $\mathbb{D}$ the formula for the exact computation
-$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
-×$\times\left(|z_{j_{n}}^{(-)}|\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}+\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{Z}_{M_{1}}\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}\right),$ (4.14)
-where $\boldsymbol{e}_{n}$ is the $n$th canonical (column) vector in $\mathbb{R}^{M_{1}}$ and the upper bound
+In this subsection, we present sensitivity analysis for poles $z_{j}$ of the rational function (1.1), reconstructed with the method described in Subsection 4.1.
+First, we consider the case of unstructured perturbations.
+Poles $z_{n}^{(-)}$, $n=1,\ldots,M_{1}$ located inside the unit circle $\mathbb{D}$ are reconstructed as eigenvalues of the matrix pencil (4.12).
+Since Hankel matrices in (4.12) have structures similar to those of Hankel matrices in (2.6), we apply the ideas from Subsection 3.1 without any changes and obtain the formula for the exact computation of sensitivities $\rho_{n}^{(-)}$ with respect to unstructured perturbations of the poles $z^{(-)}_{n}$, $n=1,\ldots,M_{1}$,
+$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{n}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
+$\times\left(|z_{n}^{(-)}|\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}+\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{Z}_{M_{1}}\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}\right),$ (4.14)
+where $\boldsymbol{e}_{n}$ is the $n$th canonical column vector in $\mathbb{R}^{M_{1}}$. Correspondingly, for the upper bound of $\rho_{n}^{(-)}$ we have
 $\rho_{n}^{(-)}\leq\zeta_{n}^{(-)}\,\kappa^{2}\left(\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\right),$ (4.15)
 with
-$\zeta_{n}^{(-)}:=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left(\left|z_{j_{n}}^{(-)}\right|\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{j_{k}}^{(-)}\right|+\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{j_{k}}^{(-)}z_{j_{k}}^{(-)}\right|\right).$
-Now consider the case of the sensitivities $\rho_{n}^{(+)}$ of poles $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ located outside the unit circle $\mathbb{D}$.
-Inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of poles $z^{(+)}_{j_{n}}$ are reconstructed as eigenvalues of the matrix pencil (4.13).
-That means that we will compute sensitivities $\rho_{n}^{(+)}$ of the inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of the poles $z^{(+)}_{j_{n}}$, but we will say that $\rho_{n}^{(+)}$ are also sensitivities of poles $z^{(+)}_{j_{n}}$.
-To present formulas for sensitivities $\rho_{n}^{(+)}$, we have to apply some modifications to the results of the Subsection 3.1.
+$\zeta_{n}^{(-)}:=\frac{1}{\left|\gamma_{n}^{(-)}\right|}\left(\left|z_{n}^{(-)}\right|\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{k}^{(-)}\right|+\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{k}^{(-)}z_{k}^{(-)}\right|\right).$
+Now consider the case of sensitivities $\rho_{n}^{(+)}$ with respect to unstructured perturbations for the poles $z^{(+)}_{n}$, $n=1,\ldots,M_{2}$ located outside $\mathbb{D}$.
+As was already mentioned, the inverse $\frac{1}{z^{(+)}_{n}}$ of the poles $z^{(+)}_{n}$ are reconstructed as eigenvalues of the matrix pencil (4.13).
+Thus, we compute sensitivities $\rho_{n}^{(+)}$ of the inverse $\frac{1}{z^{(+)}_{n}}$ of the poles $z^{(+)}_{n}$, but we say that $\rho_{n}^{(+)}$ are also sensitivities of $z^{(+)}_{n}$.
+To present formulas for sensitivities $\rho_{n}^{(+)}$, we have to apply some modifications to the results from Subsection 3.1.
 Matrices $\mathbf{H}^{(+)}_{M_{2}}(0)$ and $\mathbf{H}^{(+)}_{M_{2}}(1)$ satisfy factorizations
@@
 $\mathbf{H}^{(+)}_{M_{2}}(1)=-\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)}).$ (4.17)
-Thus left and right eigenvectors of the GEP of the Hankel matrix pencil (4.13) corresponding to the eigenvalue $\frac{1}{z^{(+)}_{j_{n}}}$ are given by formulas
+Thus, right and left eigenvectors of the GEP of the Hankel matrix pencil (4.13) corresponding to the eigenvalue $\frac{1}{z^{(+)}_{n}}$ are given by formulas
 $\boldsymbol{q}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n},\quad\boldsymbol{p}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n},$ (4.18)
-where $\boldsymbol{e}_{n}$ is the $n$th canonical (column) vector in $\mathbb{R}^{M_{2}}$ Taking into account (4.16) and (4.18), we get
+respectively, where $\boldsymbol{e}_{n}$ is the $n$th canonical column vector in $\mathbb{R}^{M_{2}}$ . Taking into account (4.16) and (4.18), we get
 $(\boldsymbol{p}^{(n)})^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\boldsymbol{q}^{(n)}$ $=\left(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}\right)^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}$
-$=-\boldsymbol{e}_{n}^{T}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\boldsymbol{e}_{n}=-\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}.$ (4.19)
-From (3.4) applying (4.16)–(4.19), we obtain the formula for the exact computation of sensitivities $\rho_{n}^{(+)}$ of the poles $z^{(+)}_{j_{n}}$ (the eigenvalues $\frac{1}{z^{(+)}_{j_{n}}}$ of the matrix pencil (4.13)) located outside the unit circle
-$\rho_{n}^{(+)}=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}|}\bigg{\|}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}\bigg{\|}_{2}^{2}$
-×$\times\bigg{(}\frac{1}{|z_{j_{n}}^{(+)}|}\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{2}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}+\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}\bigg{)},$ (4.20)
-and the upper bound
-$\rho^{(+)}_{n}\leq\frac{\frac{1}{|z_{j_{n}}^{(+)}|}\,\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big{\|}_{2}+\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big{\|}_{2}}{|\gamma_{j_{n}}^{(+)}\,(z^{(+)}_{j_{n}})^{-2}|}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right).$
-Now since
-$\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big{\|}_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-2}|,\ \ \ \big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big{\|}_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-3}|,$
-we get for $n=1,\ldots,M_{2}$
+$=-\boldsymbol{e}_{n}^{T}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\boldsymbol{e}_{n}=-\gamma_{n}^{(+)}\,(z_{n}^{(+)})^{-2}.$ (4.19)
+[…]
+$\rho_{n}^{(+)}=\frac{1}{|\gamma_{n}^{(+)}\,(z_{n}^{(+)})^{-2}|}\bigg\|\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\,\boldsymbol{e}_{n}\bigg\|_{2}^{2}$
+$\times\bigg(\frac{1}{|z_{n}^{(+)}|}\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{2}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}+\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}\bigg),$ (4.20)
+and the corresponding upper bound
+$\rho^{(+)}_{n}\leq\frac{\frac{1}{|z_{n}^{(+)}|}\,\big\|\big(\mathbf{Z}^{(+)}_{M_{2}}\big)^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big\|_{2}+\big\|\big(\mathbf{Z}^{(+)}_{M_{2}}\big)^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big\|_{2}}{|\gamma_{n}^{(+)}\,(z^{(+)}_{n})^{-2}|}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right).$
+Since
+$\big\|\big(\mathbf{Z}^{(+)}_{M_{2}}\big)^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big\|_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{k}\,(z^{(+)}_{k})^{-2}|,\ \ \ \big\|\big(\mathbf{Z}^{(+)}_{M_{2}}\big)^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big\|_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{k}\,(z^{(+)}_{k})^{-3}|,$
+the last inequality implies
 $\rho_{n}^{(+)}\leq\zeta_{n}^{(+)}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right),$ (4.21)
-with
-$\zeta_{n}^{(+)}:=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z^{(+)}_{j_{n}})^{-2}|}\left(\frac{1}{|z_{j_{n}}^{(+)}|}\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-2}|+\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-3}|\right).$
-Define further two vectors of unstructured sensitivities $\boldsymbol{\rho}^{(-)}:=(\rho_{1}^{(-)},\ldots,\rho^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\rho}^{(+)}:=(\rho_{1}^{(+)},\ldots,\rho^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$, and let $\boldsymbol{\zeta}^{(-)}:=(\zeta_{1}^{(-)},\ldots,\zeta^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\zeta}^{(+)}:=(\zeta_{1}^{(+)},\ldots,\zeta^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$.
-Then using (4.15) and (4.21), we write the following upper bounds for the $\ell_{2}$-norm of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
+where
+$\zeta_{n}^{(+)}:=\frac{1}{|\gamma_{n}^{(+)}\,(z^{(+)}_{n})^{-2}|}\left(\frac{1}{|z_{n}^{(+)}|}\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{k}\,(z^{(+)}_{k})^{-2}|+\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{k}\,(z^{(+)}_{k})^{-3}|\right).$
+Define further two vectors of sensitivities $\boldsymbol{\rho}^{(-)}:=(\rho_{1}^{(-)},\ldots,\rho^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\rho}^{(+)}:=(\rho_{1}^{(+)},\ldots,\rho^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$, and let $\boldsymbol{\zeta}^{(-)}:=(\zeta_{1}^{(-)},\ldots,\zeta^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\zeta}^{(+)}:=(\zeta_{1}^{(+)},\ldots,\zeta^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$.
+Then using (4.15) and (4.21), we write the following upper bounds for the $\ell_{2}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
 $\|\boldsymbol{\rho}^{(-)}\|_{2}\leq\|\boldsymbol{\zeta}^{(-)}\|_{2}\,\kappa^{2}\left(\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\right),\quad\|\boldsymbol{\rho}^{(+)}\|_{2}\leq\|\boldsymbol{\zeta}^{(+)}\|_{2}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right).$ (4.22)
-Using ideas similar as in Subsection 3.1, we write also the upper bounds for the $\ell_{1}$-norm of the sensitivity vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
-$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{-1}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2}$ (4.23)
+Applying ideas similar to those in Subsection 3.1, we write also the upper bounds for the $\ell_{1}$-norms of the sensitivity vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
+$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{-1}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2},$ (4.23)
 $\|\boldsymbol{\rho}^{(+)}\|_{1}$ $\leq\zeta^{(+)}_{\max}\,\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{2}^{2}\,\|\mathbf{V}^{-1}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{\mathrm{F}}^{2},$ (4.24)
 where $\zeta_{\max}^{(-)}:=\max_{j=1,\ldots,M_{1}}\zeta_{j}^{(-)}$ and $\zeta_{\max}^{(+)}:=\max_{j=1,\ldots,M_{2}}\zeta_{j}^{(+)}$.
-Naturally, arises the question how sharp are the obtained inequalities (4.224.24).
-In Example 5.3, we present the numerical computations for the exact values of the $\ell_{2}$- and $\ell_{1}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bound.
-Based on the results presented in Table 10, we can conclude that estimates (4.224.24) are quite sharp.
+Naturally, arises the question of how sharp are the inequalities (4.22)–(4.24).
+This question is discussed in Example 5.3.
 Let us now consider the case of structured perturbations. Define further matrices
@@
 by their components
-$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}:=$ $\frac{1}{\gamma_{j_{n}}^{(-)}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{m-k,n}$
-$-z_{j_{n}}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg{)}_{m-k+1,n}\bigg{)}$ (4.25)
+$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}$:= $\frac{1}{\gamma_{n}^{(-)}}\bigg(\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg(\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg)_{k,n}\,\bigg(\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg)_{m-k,n}$
+$-z_{n}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg(\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg)_{k,n}\,\bigg(\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{-1}\bigg)_{m-k+1,n}\bigg)$ (4.25)
 and
-$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}:=$ $-\frac{1}{\gamma_{j_{n}}^{(+)}\,z_{j_{n}}^{-2}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{m-k,n}$
-$-\frac{1}{z_{j_{n}}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg{)}_{m-k+1,n}\bigg{)}.$ (4.26)
-Formulas (4.25) and (4.26) are obtained using similar ideas from Subsection 3.2. In (4.26), we take into account (4.19).
-Now we define a structured sensitivity of the pole $z^{(-)}_{j_{n}}$ located inside the unit circle with respect to the Fourier coefficient $\hat{r}_{-m}$ by $\eta_{n,m}^{(-)}:=|S_{n,m}^{(-)}|$ for $n=1,\ldots,M_{1}$ and $m=1,\ldots,2M_{1}$.
-Similarly, by $\eta_{n,m}^{(+)}:=|S_{n,m}^{(+)}|$ we define a structured sensitivity of the pole $z^{(+)}_{j_{n}}$ located outside the unit circle with respect to the Fourier coefficient $\hat{r}_{m}$ for $n=1,\ldots,M_{2}$ and $m=1,\ldots,2M_{2}$.
-Note that Fourier coefficients $\hat{r}_{-m}$ and $\hat{r}_{m}$ are $m$th measurements in this case.
-We define also two vectors
+$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}$:= $-\frac{1}{\gamma_{n}^{(+)}\,z_{n}^{-2}}\bigg(\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg)_{k,n}\,\bigg(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg)_{m-k,n}$
+$-\frac{1}{z_{n}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg)_{k,n}\,\bigg(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{-1}\bigg)_{m-k+1,n}\bigg).$ (4.26)
+Formulas (4.25) and (4.26) are obtained using ideas similar to those presented in Subsection 3.2. In (4.26) we take into account (4.19).
+Now we define the sensitivity $\eta_{n,m}^{(-)}$ of the pole $z^{(-)}_{n}$ located inside the unit circle with respect to the Fourier coefficient $\hat{r}_{-m}$ by $\eta_{n,m}^{(-)}:=|S_{n,m}^{(-)}|$ for $n=1,\ldots,M_{1}$ and $m=1,\ldots,2M_{1}$.
+Similarly, by $\eta_{n,m}^{(+)}:=|S_{n,m}^{(+)}|$ we define the sensitivity of the pole $z^{(+)}_{n}$ located outside the unit circle with respect to the Fourier coefficient $\hat{r}_{m}$ for $n=1,\ldots,M_{2}$ and $m=1,\ldots,2M_{2}$.
+By
+$(\boldsymbol{\eta}^{(n)}_{2M_{1}})^{(-)}:=(\eta_{n,1}^{(-)},\ldots,\eta_{n,2M_{1}}^{(-)})^{T}\text{ and }(\boldsymbol{\eta}^{(n)}_{2M_{2}})^{(+)}:=(\eta_{n,1}^{(+)},\ldots,\eta_{n,2M_{2}}^{(+)})^{T}$ (4.27)
+we denote the corresponding sensitivity vectors. Note that Fourier coefficients $\hat{r}_{-m}$ and $\hat{r}_{m}$ are $m$th measurements in this case. We consider also two vectors
 $\boldsymbol{\eta}^{(-)}:=(\eta_{1}^{(-)},\ldots,\eta_{M_{1}}^{(-)})^{T}\in\mathbb{R}^{M_{1}},\quad\boldsymbol{\eta}^{(+)}:=(\eta_{1}^{(+)},\ldots,\eta_{M_{2}}^{(+)})^{T}\in\mathbb{R}^{M_{2}},$
-where the entries
+where their components
 $\eta_{n}^{(-)}$ $:=\|(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{1}})^{(-)}\|_{2}=\sqrt{(\eta^{(-)}_{n,1})^{2}+\ldots+(\eta^{(-)}_{n,2M_{1}})^{2}},\quad n=1,\ldots,M_{1},$ (4.28)
-and
 $\eta_{n}^{(+)}$ $:=\|(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{2}})^{(+)}\|_{2}=\sqrt{(\eta^{(+)}_{n,1})^{2}+\ldots+(\eta^{(+)}_{n,2M_{2}})^{2}},\quad n=1,\ldots,M_{2},$ (4.29)
-are structured sensitivities for the poles $z_{j_{n}}^{(-)}$ and $z_{j_{n}}^{(+)}$ respectively.
-Table 1: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.1
+are sensitivities of the poles $z_{n}^{(-)}$ and $z_{n}^{(+)}$, respectively, with respect to structured perturbations.
+Table 1: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ for Example 5.1
 $\sigma=10^{-3}$ $N=2$ $N=10$
 $z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$ $z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$
-min $2.54$e–$05$ $2.06$e–$04$ $3.49$e–$06$ $5.56$e–$05$
-max $2.85$e–$04$ $3.34$e–$03$ $2.7$e–$04$ $3.15$e–$03$
-average $1.34$e–$04$ $2.03$e–$03$ $1.29$e–$04$ $1.06$e–$03$
+min 2.542.54e–$05$ 2.062.06e–$04$ 3.493.49e–$06$ 5.565.56e–$05$
+max 2.852.85e–$04$ 3.343.34e–$03$ 2.72.7e–$04$ 3.153.15e–$03$
+average 1.341.34e–$04$ 2.032.03e–$03$ 1.291.29e–$04$ 1.061.06e–$03$
 $\sigma=10^{-1}$ $N=10$ $N=30$
 $z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$ $z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$
-min $3.2$e–$03$ $4.09$e–$02$ $7.19$e–$03$ $4.15$e–$02$
-max $2.28$e–$02$ $3.65$e–$01$ $2.72$e–$02$ $1.7$e–$01$
-average $1.09$e–$02$ $1.57$e–$01$ $1.36$e–$02$ $1.05$e–$01$
+min 3.23.2e–$03$ 4.094.09e–$02$ 7.197.19e–$03$ 4.154.15e–$02$
+max 2.282.28e–$02$ 3.653.65e–$01$ 2.722.72e–$02$ 1.71.7e–$01$
+average 1.091.09e–$02$ 1.571.57e–$01$ 1.361.36e–$02$ 1.051.05e–$01$
 5 Numerical experiments
-In this section, we present a few numerical experiments.
+In this section, we present several numerical experiments.
 The algorithms are implemented in Matlab and use IEEE standard floating point arithmetic with double precision.
@@
 Note that in the case when we deal with noisy data, we assume that a number of poles $M$, as well as numbers of poles located inside and outside the unit circle $M_{1}$ and $M_{2}$, respectively, are known.
-In Examples 5.1-5.3, we demonstrate recovery results and analyze sensitivities of poles with the method developed in Section 4.
-We employ $2N$ (noisy) Fourier coefficients $\hat{r}_{-k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(+)}$.
-By $(\varepsilon_{k})_{k=1}^{2N}$ we denote the random noise that is generated in Matlab by $\sigma$*randn(2N), i.e., we take multiplicative noise drawn from the standard normal distribution.
-We take $L=N$ and consider different values of $N$ and $\sigma$.
-To compute sensitivities $\rho_{j}^{(-)}$, $\rho_{j}^{(+)}$ and ,$\eta_{j}^{(-)}$, $\eta_{j}^{(+)}$, we use formulas (4.14), (4.20) and (4.27), (4.28), respectively.
-Upper bounds for sensitivities $\rho_{j}^{(-)}$ and $\rho_{j}^{(+)}$ are computed via (4.15) and (4.21), respectively.
 […]
 $\sigma=10^{-3}$ $N=2$ $N=10$ $\sigma=10^{-1}$ $N=10$ $N=30$
-min $1.78$e–$04$ $1.17$e–$04$ min $4.03$e–$02$ $2.77$e–$02$
-max $2.11$e–$03$ $2.46$e–$03$ max $3.04$e–$01$ $1.44$e–$01$
-average $1.44$e–$03$ $8.48$e–$04$ average $1.11$e–$01$ $9.55$e–$02$
-Table 3: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.2 for $\sigma=10^{-3}$
+min 1.781.78e–$04$ 1.171.17e–$04$ min 4.034.03e–$02$ 2.772.77e–$02$
+max 2.112.11e–$03$ 2.462.46e–$03$ max 3.043.04e–$01$ 1.441.44e–$01$
+average 1.441.44e–$03$ 8.488.48e–$04$ average 1.111.11e–$01$ 9.559.55e–$02$
+In Examples 5.1-5.3, we demonstrate recovery results for three different rational functions and analyze them employing sensitivity analysis as was discussed in Section 4.
+We utilize $2N$ noisy Fourier coefficients $\hat{r}_{-k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{n}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+\varepsilon_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{n}^{(+)}$.
+By $(\varepsilon_{k})_{k=1}^{2N}$ we denote the random noise that is generated in Matlab by $\sigma*$randn(2N), i.e.
+we consider a multiplicative noise drawn from the standard normal distribution.
+We set $L=N$ and perform computations for different values of $N$ and $\sigma$.
+For $\sigma>0$ we perform 10 itterations.
+To compute sensitivities $\rho_{n}^{(-)}$, $\rho_{n}^{(+)}$ and $\eta_{n}^{(-)}$, $\eta_{n}^{(+)}$, we use formulas (4.14), (4.20) and (4.28), (4.29), respectively.
+Upper bounds for sensitivities $\rho_{n}^{(-)}$ and $\rho_{n}^{(+)}$ are computed via (4.15) and (4.21), respectively.
+Table 3: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ for Example 5.2 for $\sigma=10^{-3}$
 $N=8$ $z_{1}^{(-)}=0.9$ $z_{2}^{(-)}=-0.9$ $z_{3}^{(-)}=0.9\mathrm{i}$ $z_{4}^{(-)}=-0.9\mathrm{i}$
-min $7.08$e–$05$ $4.31$e–$06$ $2.79$e–$06$ $9.06$e–$06$
-max $6.74$e–$04$ $3.17$e–$04$ $1.65$e–$04$ $1.45$e–$04$
-average $3.18$e–$04$ $1.48$e–$04$ $7.87$e–$05$ $5.57$e–$05$
+min 7.087.08e–$05$ 4.314.31e–$06$ 2.792.79e–$06$ 9.069.06e–$06$
+max 6.746.74e–$04$ 3.173.17e–$04$ 1.651.65e–$04$ 1.451.45e–$04$
+average 3.183.18e–$04$ 1.481.48e–$04$ 7.877.87e–$05$ 5.575.57e–$05$
 $N=8$ $z_{1}^{(+)}=1.1$ $z_{2}^{(+)}=-1.1$ $z_{3}^{(+)}=1.1\mathrm{i}$ $z_{4}^{(+)}=-1.1\mathrm{i}$
-min $1.71$e–$05$ $1.13$e–$06$ $1.31$e–$05$ $8.74$e–$06$
-max $2.86$e–$04$ $2.7$e–$04$ $1.89$e–$04$ $1.9$e–$04$
-average $1.44$e–$04$ $1.19$e–$04$ $1.04$e–$04$ $9.43$e–$05$
-Table 4: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.2 for $\sigma=10^{-1}$
+min 1.711.71e–$05$ 1.131.13e–$06$ 1.311.31e–$05$ 8.748.74e–$06$
+max 2.862.86e–$04$ 2.72.7e–$04$ 1.891.89e–$04$ 1.91.9e–$04$
+average 1.441.44e–$04$ 1.191.19e–$04$ 1.041.04e–$04$ 9.439.43e–$05$
+Table 4: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ for Example 5.2 for $\sigma=10^{-1}$
 $N=8$ $z_{1}^{(-)}=0.9$ $z_{2}^{(-)}=-0.9$ $z_{3}^{(-)}=0.9\mathrm{i}$ $z_{4}^{(-)}=-0.9\mathrm{i}$
-min $6.41$e–$03$ $2.65$e–$03$ $5.14$e–$04$ $1.38$e–$03$
-max $2.02$e–$01$ $2.37$e–$02$ $2.89$e–$02$ $7.51$e–$03$
-average $3.99$e–$02$ $1.13$e–$02$ $1.06$e–$02$ $1.82$e–$02$
+min 6.416.41e–$03$ 2.652.65e–$03$ 5.145.14e–$04$ 1.381.38e–$03$
+max 2.022.02e–$01$ 2.372.37e–$02$ 2.892.89e–$02$ 7.517.51e–$03$
+average 3.993.99e–$02$ 1.131.13e–$02$ 1.061.06e–$02$ 1.821.82e–$02$
 $N=8$ $z_{1}^{(+)}=1.1$ $z_{2}^{(+)}=-1.1$ $z_{3}^{(+)}=1.1\mathrm{i}$ $z_{4}^{(+)}=-1.1\mathrm{i}$
-min $8.17$e–$04$ $8.71$e–$04$ $1.93$e–$03$ $2.83$e–$03$
-max $2.95$e–$02$ $2.61$e–$02$ $2.48$e–$02$ $2.41$e–$02$
-average $1.56$e–$02$ $1.4$e–$02$ $1.24$e–$02$ $1.15$e–$02$
+min 8.178.17e–$04$ 8.718.71e–$04$ 1.931.93e–$03$ 2.832.83e–$03$
+max 2.952.95e–$02$ 2.612.61e–$02$ 2.482.48e–$02$ 2.412.41e–$02$
+average 1.561.56e–$02$ 1.41.4e–$02$ 1.241.24e–$02$ 1.151.15e–$02$
 […]
 $N=8$ $\sigma=10^{-3}$ $\sigma=10^{-1}$
-min $4.2$e–$03$ $4.25$e–$01$
-max $1.73$e–$02$ $1.22$e–$00$
-average $9.7$e–$03$ $6.35$e–$01$
+min 4.24.2e–$03$ 4.254.25e–$01$
+max 1.731.73e–$02$ 1.221.22e–$00$
+average 9.79.7e–$03$ 6.356.35e–$01$
+Example 5.1.
+Let a rational function $r(z)$ as in (1.1) be given by parameters $M=2$ and
+$\boldsymbol{z}=(-0.1,-2.1)^{T},\quad\boldsymbol{\gamma}=(0.5,0.5)^{T}$
+(see [6, 24]). When $\sigma=0$, i.e. in case of recovery from exact data, we get the following reconstruction errors for $N=2$
+[…]
+Since we are interested to investigate which pole is the most sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ independently and present the results in Table 1.
+Results for the recovery of the coefficient vector are shown in Table 2.
 Table 6: Sensitivities for Example 5.2.
@@
 1 $0.9$ 0.243 7.748 13.548
-2 $-0.9$ 0.121 3.874 6.774
-3 $0.9\mathrm{i}$ 0.081 2.582 4.516
-4 $-0.9\mathrm{i}$ 0.06 1.937 3.387
+2 0.9-0.9 0.121 3.874 6.774
+3 $9\mathrm{i} 0.0$81 2.582 4.516
+4 $9\mathrm{i} 0.0$6 1.937 3.387
 $j$ poles $z_{j}^{(+)}$ $\eta_{j}^{(+)}$ $\rho_{j}^{(+)}$ bound $\rho_{j}^{(+)}$
 1 $1.1$ 0.056 3.533 5.695
-2 $-1.1$ 0.047 2.944 4.745
+2 1.1-1.1 0.047 2.944 4.745
 3 $1.1\mathrm{i}$ 0.04 2.523 4.067
-4 $-1.1\mathrm{i}$ 0.035 2.208 3.559
-Example 5.1.
-Let a rational function $r(z)$ in (1.1) be given by parameters $M=2$ and
-$\boldsymbol{z}=(-0.1,-2.1),\quad\boldsymbol{\gamma}=(0.5,0.5)$
-(see [6, 23]). When $\sigma=0$, i.e. in case of recovery from exact data, we get the following reconstruction errors with $N=2$,
-[…]
-For $\sigma=10^{-3}$ and $\sigma=10^{-1}$, we compute 10 iterations and present the results in Table 1.
-Since we are interested to investigate which pole is more sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ separately.
-Results for the recovery of the coefficient vector are presented in Table 2.
-Table 7: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $N=4$
-$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
-min $1.58$e–$08$ $2.9$e–$08$ $6.07$e–$09$ $1.75$e–$03$
-max $5.1$e–$07$ $2.49$e–$07$ $2.7$e–$07$ $3.21$e–$01$
-average $2.48$e–$07$ $1.3$e–$07$ $1.45$e–$07$ $1.34$e–$01$
-$\sigma=10^{-5}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
-min $4.24$e–$06$ $1.18$e–$06$ $2.57$e–$06$ $3.16$e–$00$
-max $6.45$e–$05$ $2.77$e–$05$ $2.92$e–$05$ $9.46$e+$01$
-average $2.31$e–$05$ $1.13$e–$05$ $1.5$e–$05$ $2.02$e+$01$
-Analyzing numerical results in Table 1, we conclude that for $\sigma=0.1$ our method still reconstructs the locations of the poles quite good.
-On the other hand, using more (noisy) Fourier coefficients also gives us better recovery results.
+4 $1.1\mathrm{i}$ 0.035 2.208 3.559
+Analyzing numerical results from Table 1, we conclude that for $\sigma=10^{-1}$ our method still reconstructs the locations of the poles quite good.
+On the other hand, employing more noisy Fourier coefficients also gives us better recovery results.
 We notice that a pole $z_{1}^{(-)}=-0.1$ is less sensitive to perturbations than a pole $z_{1}^{(+)}=-2.1$.
-This numerical observation is also confirmed with sensitivities values.
+This numerical observation is also confirmed with the sensitivities values.
 For a pole $z_{1}^{(-)}=-0.1$, we have $\rho_{1}^{(-)}=2\cdot 10^{-1}$ and $\eta_{1}^{(-)}=1.97$, and for a pole $z_{1}^{(+)}=-2.1$, we have $\rho_{1}^{(+)}=9.52\cdot 10^{-1}$ and $\eta_{1}^{(+)}=6.65$.
 Both unstructured and structured sensitivities are larger for a pole $z_{1}^{(+)}=-2.1$ located outside the unit circle.
-We compute also vectors of structured perturbations
-$(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T},$
-which means that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive with respect to the perturbation of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ with respect to the perturbation of the Fourier coefficient $\hat{r}_{2}$.
+We compute also vectors (4.27)
+$(\boldsymbol{\eta}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\eta}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T},$
+which provide us with the additional information that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive to the perturbations of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ is the most sensitive to the perturbations of the Fourier coefficient $\hat{r}_{2}$.
+Table 7: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ for Example 5.3 for $N=4$
+$\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
+min 1.581.58e–$08$ 2.92.9e–$08$ 6.076.07e–$09$ 1.751.75e–$03$
+max 5.15.1e–$07$ 2.492.49e–$07$ 2.72.7e–$07$ 3.213.21e–$01$
+average 2.482.48e–$07$ 1.31.3e–$07$ 1.451.45e–$07$ 1.341.34e–$01$
+$\sigma=10^{-5}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
+min 4.244.24e–$06$ 1.181.18e–$06$ 2.572.57e–$06$ 3.163.16e–$00$
+max 6.456.45e–$05$ 2.772.77e–$05$ 2.922.92e–$05$ 9.469.46e+$01$
+average 2.312.31e–$05$ 1.131.13e–$05$ 1.51.5e–$05$ 2.022.02e+$01$
+Example 5.2.
+We consider the case when poles are located very close to the unit circle. Let a rational function $r(z)$ as in (1.1) of type $(7,8)$, i.e with $M=8$, be given by vectors of poles and coefficients
+$\boldsymbol{z}=(0.9,-0.9,0.9\mathrm{i},-0.9\mathrm{i},1.1,-1.1,1.1\mathrm{i},-1.1\mathrm{i})^{T},\quad\boldsymbol{\gamma}=(1,2,3,4,5,6,7,8)^{T},$
+respectively. First, we consider recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we obtain the following reconstruction errors
+$e(\boldsymbol{z})=1.45\cdot 10^{-15},\quad e(\boldsymbol{\gamma})=5.48\cdot 10^{-14}.$
+To reconstruct from noisy data, we consider $\sigma=10^{-3}$ and $\sigma=10^{-1}$.
+The recovery results for each pole $z_{j}$ are presented in Tables 3 and 4, as well as for the coefficient vector in Table 5.
+Based on these numerical results, we conclude that a pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations, which is corroborated by the sensitivities shown in Table 6.
 Table 8: Sensitivities for Example 5.3.
@@
 $j$ poles $z_{j}^{(+)}$ $\eta_{j}^{(+)}$ $\rho_{j}^{(+)}$ bound $\rho_{j}^{(+)}$
-1 $2$ $1.973\cdot 10^{1}$ 5.43 $1.993\cdot 10^{1}$
+1 22 $1.973\cdot 10^{1}$ 5.43 $1.993\cdot 10^{1}$
 2 $50$ $1.577\cdot 10^{4}$ $2.204\cdot 10^{3}$ $6.477\cdot 10^{3}$
 Example 5.3.
-We consider an example when poles are located very close to the unit circle. Let a rational function $r(z)$ of type $(7,8)$ in (1.1), i.e with $M=8$, be given by vectors of poles and coefficients
-$\boldsymbol{z}=(0.9,-0.9,0.9\mathrm{i},-0.9\mathrm{i},1.1,-1.1,1.1\mathrm{i},-1.1\mathrm{i}),\quad\boldsymbol{\gamma}=(1,2,3,4,5,6,7,8),$
-respectively. First, we consider recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we get the following reconstruction errors
-$e(\boldsymbol{z})=1.45\cdot 10^{-15},\quad e(\boldsymbol{\gamma})=5.48\cdot 10^{-14}.$
-[…]
-$N=4$ $\sigma=10^{-7}$ $\sigma=10^{-5}$
-min $7.38$e–$04$ $3.79$e–$02$
-max $7.38$e–$03$ $6.95$e–$01$
-average $4$e–$03$ $4.11$e–$01$
-In case of recovery from noisy data, we compute 10 iterations and present recovery result for each pole $z_{j}$ in Tables 3 and 4.
-Analyzing numerical results, we conclude that the pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations.
-This conclusion is confirmed with the sensitivities presented in Table 6.
-Results for the recovery of the coefficient vector are presented in Table 5.
-Table 10: $\ell_{p}$-norm and upper bound of $\ell_{p}$-norm of $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3 for $p=1,2$
-$\|\boldsymbol{\rho}^{(-)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{2}$) $\|\boldsymbol{\rho}^{(+)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{2}$)
-$N=4$ $2.86$e+$01$ $6.866$e+$01$ $2.204$e+$03$ $6.477$e+$03$
-$\|\boldsymbol{\rho}^{(-)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{1}$) $\|\boldsymbol{\rho}^{(+)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{1}$)
-$N=4$ $4.01$e+$01$ $5.725$e+$01$ $2.209$e+$03$ $6.802$e+$03$
-Example 5.3.
-Finally, let a rational function $r(z)$ of type $(3,4)$ in (1.1), i.e with $M=4$, be given by vectors of poles and coefficients
-$\boldsymbol{z}=(0.2,0.5,2,50),\quad\boldsymbol{\gamma}=(1,1,1,1),$
+Finally, let a rational function $r(z)$ as in (1.1) of type $(3,4)$, i.e with $M=4$, be given by vectors of poles and coefficients
+$\boldsymbol{z}=(0.2,0.5,2,50)^{T},\quad\boldsymbol{\gamma}=(1,1,1,1)^{T},$
 respectively. For $N=4$ and $\sigma=0$, we get the following reconstruction errors
 $e(\boldsymbol{z})=1.42\cdot 10^{-13},\quad e(\boldsymbol{\gamma})=9.27\cdot 10^{-15}.$
-Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$.
-We compute 10 iterations and present the corresponding results in Table 7.
-Analyzing numerical results, we can conclude that the pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations.
+Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$, and present the corresponding results in Table 7.
+Analyzing these numerical results, we can conclude that a pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations.
 For $\sigma=10^{-5}$ the method does not reconstruct location of this pole correctly.
-This numerical observation is also predicted by the sensitivities values presented in Table 8.
+This numerical observation is also predicted by the sensitivities values shown in Table 8.
 Both unstructured $\rho_{2}^{(+)}=2.204\cdot 10^{3}$ and structured $\eta_{2}^{(+)}=1.577\cdot 10^{4}$ sensitivities are sufficiently larger than sensitivities of other poles.
@@
 We present results regarding recovery of the coefficient vector in Table 9.
-We compute also vectors of structured perturbations for poles inside the unit circle
-$(\boldsymbol{\mathcal{S}}^{(1)}_{4})^{(-)}=(0.555,4.999,13.333,11.111)^{T},\quad(\boldsymbol{\mathcal{S}}^{(2)}_{4})^{(-)}=(0.222,2.666,9.999,11.111)^{T},$
+[…]
+$N=4$ $\sigma=10^{-7}$ $\sigma=10^{-5}$
+min 7.387.38e–$04$ 3.793.79e–$02$
+max 7.387.38e–$03$ 6.956.95e–$01$
+average 44e–$03$ 4.114.11e–$01$
+We compute also vectors (4.27) for poles located inside the unit circle
+$(\boldsymbol{\eta}^{(1)}_{4})^{(-)}=(0.555,4.999,13.333,11.111)^{T},\quad(\boldsymbol{\eta}^{(2)}_{4})^{(-)}=(0.222,2.666,9.999,11.111)^{T},$
 and outside the unit circle
-$(\boldsymbol{\mathcal{S}}^{(1)}_{4})^{(+)}=(3.472\cdot 10^{-3},0.354,9.375,17.36)^{T},$
-and
-$(\boldsymbol{\mathcal{S}}^{(2)}_{4})^{(+)}=(54.25,2.929\cdot 10^{3},1.106\cdot 10^{4},1.085\cdot 10^{4})^{T}.$
-That means that poles $z_{1}^{(-)}=0.2$ and $z_{2}^{(-)}=0.5$ are the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-3}$, the pole $z_{1}^{(+)}=2$ – the Fourier coefficient $\hat{r}_{4}$, and the pole $z_{2}^{(+)}=50$ – the Fourier coefficient $\hat{r}_{3}$.
-In Subsection 4.2, we derived the upper bounds (4.224.24) for the $\ell_{2}$- and the $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and we want to understand how sharp these estimates are.
-In order to do it, we compute the exact values for the corresponding norms and the upper bounds obtained in (4.224.24) and present the corresponding results in Table 10.
-Comparing these results we can conclude that the estimates (4.224.24) are quite sharp.
-Table 11: Reconstruction results for $r$ from Example 5.1
+$(\boldsymbol{\eta}^{(1)}_{4})^{(+)}$ $=(3.472\cdot 10^{-3},0.354,9.375,17.36)^{T},$
+$(\boldsymbol{\eta}^{(2)}_{4})^{(+)}$ $=(54.25,2.929\cdot 10^{3},1.106\cdot 10^{4},1.085\cdot 10^{4})^{T}.$
+This implies that the poles $z_{1}^{(-)}=0.2$ and $z_{2}^{(-)}=0.5$ are the most sensitive to perturbations of the Fourier coefficients $\hat{r}_{-3}$ and $\hat{r}_{-4}$, respectively, while the poles $z_{1}^{(+)}=2$ and $z_{2}^{(+)}=50$ are the most sensitive to the perturbations of the Fourier coefficients $\hat{r}_{4}$ and $\hat{r}_{3}$, respectively.
+Table 10: $\ell_{p}$-norms of $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bounds for Example 5.3 for $p=1,2$
+$\|\boldsymbol{\rho}^{(-)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{2}$) $\|\boldsymbol{\rho}^{(+)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{2}$)
+$N=4$ 2.862.86e+$01$ 6.8666.866e+$01$ 2.2042.204e+$03$ 6.4776.477e+$03$
+$\|\boldsymbol{\rho}^{(-)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{1}$) $\|\boldsymbol{\rho}^{(+)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{1}$)
+$N=4$ 4.014.01e+$01$ 5.7255.725e+$01$ 2.2092.209e+$03$ 6.8026.802e+$03$
+In Subsection 4.2, we derived the upper bounds (4.22)–(4.24) for the $\ell_{2}$- and the $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$.
+In order to understand how sharp these estimates are, we compute the exact values for the corresponding norms and their upper bounds as in (4.22)–(4.24) and present the obtained numerical results in Table 10.
+Comparing these results, we can conclude that the estimates (4.22)–(4.24) are quite sharp.
+Table 11: Reconstruction results for Example 5.1
 Method via Hankel pencil AAA method
 […]
-min $1.52$e–$08$ $1.23$e–$08$ min $3.54$e–$08$ $4.55$e–$08$
-max $1.19$e–$07$ $5.51$e–$08$ max $6.62$e–$07$ $4.84$e–$07$
-average $7.51$e–$08$ $2.89$e–$08$ average $3.83$e–$07$ $2.82$e–$07$
+min 1.521.52e–$08$ 1.231.23e–$08$ min 3.543.54e–$08$ 4.554.55e–$08$
+max 1.191.19e–$07$ 5.515.51e–$08$ max 6.626.62e–$07$ 4.844.84e–$07$
+average 7.517.51e–$08$ 2.892.89e–$08$ average 3.833.83e–$07$ 2.822.82e–$07$
 Method via Hankel pencil AAA method
 […]
-min $4.34$e–$03$ $1.06$e–$03$ min $2.37$e–$02$ $2.27$e–$02$
-max $1.23$e–$02$ $4.24$e–$03$ max $2.05$e+$00$ $7.56$e–$02$
-average $7.64$e–$03$ $2.71$e–$03$ average $4.52$e–$01$ $4.55$e–$02$
-Table 12: Reconstruction results for $r$ from Example 5.2
+min 4.344.34e–$03$ 1.061.06e–$03$ min 2.372.37e–$02$ 2.272.27e–$02$
+max 1.231.23e–$02$ 4.244.24e–$03$ max 2.052.05e+$00$ 7.567.56e–$02$
+average 7.647.64e–$03$ 2.712.71e–$03$ average 4.524.52e–$01$ 4.554.55e–$02$
+Table 12: Reconstruction results for Example 5.2
 Method via Hankel pencil AAA method
 […]
-min $1.96$e–$05$ $4.79$e–$06$ min $7.03$e–$05$ $7.18$e–$05$
-max $1.37$e–$04$ $6.22$e–$05$ max $4.49$e–$04$ $3.43$e–$04$
-average $5.89$e–$05$ $2.34$e–$05$ average $2.98$e–$04$ $2.25$e–$04$
+min 1.961.96e–$05$ 4.794.79e–$06$ min 7.037.03e–$05$ 7.187.18e–$05$
+max 1.371.37e–$04$ 6.226.22e–$05$ max 4.494.49e–$04$ 3.433.43e–$04$
+average 5.895.89e–$05$ 2.342.34e–$05$ average 2.982.98e–$04$ 2.252.25e–$04$
 Method via Hankel pencil AAA method
 […]
-min $2.2$e–$03$ $2.42$e–$03$ min $1.54$e–$02$ $9.73$e–$03$
-max $1.37$e–$02$ $4.86$e–$02$ max $6.69$e–$02$ $3.54$e–$02$
-average $6.93$e–$03$ $1.06$e–$02$ average $4.42$e–$02$ $5.75$e–$02$
-In Example 5.4 we present the comparison results of the performance of our approach with the AAA method for rational approximation.
-We employ $4N$ (noisy) function values $r(t_{k})(1+\varepsilon_{k})$ with $t_{k}=\mathrm{e}^{\frac{2\pi\mathrm{i}k}{4N}}$ for $k=0,\ldots,4N-1$, as input information for both methods.
-By $(\varepsilon_{k})_{k=0}^{4N-1}$ as above we denote the random noise that is generated in Matlab by $\sigma$*randn(4N), i.e., we take multiplicative noise drawn from the standard normal distribution.
+min 2.22.2e–$03$ 2.422.42e–$03$ min 1.541.54e–$02$ 9.739.73e–$03$
+max 1.371.37e–$02$ 4.864.86e–$02$ max 6.696.69e–$02$ 3.543.54e–$02$
+average 6.936.93e–$03$ 1.061.06e–$02$ average 4.424.42e–$02$ 5.755.75e–$02$
+In Example 5.4, we discuss the comparison of the performance of our approach with the AAA method for rational approximation.
+We employ $4N$ noisy function values $r(t_{k})(1+\varepsilon_{k})$ with $t_{k}=\mathrm{e}^{\frac{2\pi\mathrm{i}k}{4N}}$ for $k=0,\ldots,4N-1$, as input information for both methods.
+By $(\varepsilon_{k})_{k=0}^{4N-1}$ as above we denote the random noise that is generated in Matlab by $\sigma*$randn(4N).
 The corresponding modification of our method is discussed in Remark 4.1.
-We referee the reader to [14] regarding description of the AAA method.
+We refer the interested reader to [15] regarding description of the AAA method.
 Note only that AAA takes the form of iteration with respect to a type of a rational approximant that is represented at each step in the barycentric form.
-In case of reconstruction of rational function of type $(M-1,M)$ using exact data, AAA terminates after $M+1$ iteration steps and returns $r$ as
+In case of reconstruction of rational function of type (M1,M)(M-1,M) using exact data, AAA terminates after $M+1$ iteration steps and returns $r$ as
 $r(t)=\sum\limits_{s=1}^{M+1}\frac{r(t_{k_{s}})w_{s}}{t-t_{k_{s}}}\bigg/\sum\limits_{s=1}^{M+1}\frac{w_{s}}{t-t_{k_{s}}},$
 where points $t_{k_{s}}$ are chosen by some greedy algorithm and weights $w_{s}$ are computed by solving a linear least-squares problem over the subset of sample points that have not been selected as support points, $\{t_{0},\ldots,t_{4N-1}\}\setminus\{t_{k_{1}},\ldots,t_{k_{M+1}}\}$ (see also [9, 10, 11] for details).
-The AAA method has the overall computational cost of $\mathcal{O}(NM^{3})$.
-Table 13: Reconstruction results for $r$ from Example 5.3
+The AAA method has the overall computational cost of $\mathcal{O}(NM^{3})$ operations.
+Table 13: Reconstruction results for Example 5.3
 Method via Hankel pencil AAA method
 […]
-min $4.35$e–$03$ $7.99$e–$05$ min $1.26$e–$01$ $4.89$e–$03$
-max $3.00$e–$02$ $5.66$e–$04$ max $2.03$e+$00$ $7.86$e–$02$
-average $1.42$e–$02$ $2.66$e–$04$ average $1.48$e+$00$ $4.44$e–$02$
+min 4.354.35e–$03$ 7.997.99e–$05$ min 1.261.26e–$01$ 4.894.89e–$03$
+max 3.003.00e–$02$ 5.665.66e–$04$ max 2.032.03e+$00$ 7.867.86e–$02$
+average 1.421.42e–$02$ 2.662.66e–$04$ average 1.481.48e+$00$ 4.444.44e–$02$
 Method via Hankel pencil AAA method
 […]
-min $7.89$e–$04$ $4.48$e–$06$ min $2.6$e–$02$ $1.01$e–$03$
-max $1.28$e–$03$ $5.58$e–$05$ max $2.6$e–$01$ $1.01$e–$02$
-average $2.98$e–$03$ $2.4$e–$05$ average $1.25$e–$01$ $4.84$e–$03$
+min 7.897.89e–$04$ 4.484.48e–$06$ min 2.62.6e–$02$ 1.011.01e–$03$
+max 1.281.28e–$03$ 5.585.58e–$05$ max 2.62.6e–$01$ 1.011.01e–$02$
+average 2.982.98e–$03$ 2.42.4e–$05$ average 1.251.25e–$01$ 4.844.84e–$03$
 Example 5.4.
-We consider again recovery of rational functions $r$ from Examples 5.1-5.3.
-We reconstruct $r$ from $4N$ samples $r(t_{k})(1+\varepsilon_{k})$, $k=0,\ldots,4N-1$ by using our method based on Hankel pencil and AAA.
-In Tables 11- 13 we present the results for $N=15$, $L=20$ and different values of $\sigma$.
-We see that for the rational function from Example 5.2 both methods work very well.
-Analysing results of Tables 11 and 13, we conclude that for $r$ from Examples 5.1 and 5.3 we get a slightly better recovery from noisy input data with our method.
-The reason for this is that the method based on the Hankel pencil works better for the recovery of poles with certain locations.
-For Example 5.3 the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed equally good by both methods, while the Hankel pencil based approach shows better performance for the recovery of the pole $z_{2}^{(+)}=50$ than the AAA method (see Tables 14).
-Table 14: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 with Hankel pencil and the AAA methods
+We consider again recovery of rational functions $r$ from Examples 5.15.3.
+In Tables 1113 we present the results for $N=15$, $L=20$ (for the modified method $L\leq 2N$) and different values of $\sigma$.
+Analyzing results of Tables 1113, we conclude that for $r$ from Example 5.2 both methods work very well, while for $r$ from Examples 5.1 and 5.3 we get a slightly better recovery applying our method.
+This is because the Hankel pencil-based approach is more effective for the recovery of poles with certain locations.
+For Example 5.3 the poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed equally good by both methods, while the Hankel pencil method shows better performance for the recovery of the pole $z_{2}^{(+)}=50$ than the AAA method (see Table 14).
+Table 14: Reconstruction errors $|z_{j}-\tilde{z}_{j}|$ for poles $z_{j}$ from Example 5.3 with the Hankel pencil and the AAA methods
 Method via Hankel pencil
 $\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
-min $2.82$e–$08$ $6.98$e–$08$ $2.08$e–$07$ $4.35$e–$03$
-max $5.09$e–$07$ $3.92$e–$07$ $1.14$e–$06$ $3$e–$02$
-average $2.51$e–$07$ $1.79$e–$07$ $2.08$e–$07$ $1.72$e–$02$
+min 2.822.82e–$08$ 6.986.98e–$08$ 2.082.08e–$07$ 4.354.35e–$03$
+max 5.095.09e–$07$ 3.923.92e–$07$ 1.141.14e–$06$ 33e–$02$
+average 2.512.51e–$07$ 1.791.79e–$07$ 2.082.08e–$07$ 1.721.72e–$02$
 AAA method
 $\sigma=10^{-7}$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
-min $1.15$e–$07$ $6.42$e–$08$ $6.65$e–$07$ $1.26$e–$01$
-max $7.4$e–$07$ $5.24$e–$07$ $2.55$e–$06$ $2.03$e+$00$
-average $4.63$e–$07$ $2.34$e–$07$ $1.19$e–$06$ $1.14$e+$00$
+min 1.151.15e–$07$ 6.426.42e–$08$ 6.656.65e–$07$ 1.261.26e–$01$
+max 7.47.4e–$07$ 5.245.24e–$07$ 2.552.55e–$06$ 2.032.03e+$00$
+average 4.634.63e–$07$ 2.342.34e–$07$ 1.191.19e–$06$ 1.141.14e+$00$
 Acknowledgement
-The author would like to thank Ion Victor Gosea (MPI Magdeburg) for several valuable discussions regarding sensitivities of the eigenvalues of a matrix pencil problem.
+The author would like to thank Ion Victor Gosea (MPI Magdeburg) for several valuable discussions regarding the sensitivity analysis.
 The author would also like to thank the anonymous referees for their valuable comments that helped to improve the manuscript.